A Complete Bibliography of Publications in the Journal of Mathematics and Physics (MIT)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

−1 [Ced57]. 0 [Ced57]. 1 [Ced57]. 2n [Oon50, Sta36a, Tel53]. 2n + 1 [Rut23]. 2 [Ada43]. ^{3} [Ada43]. ^{5} [Ada43]. _{I} [Gar35, Gar37]. _{II} [Gar35, Gar37]. A_{mn}(h, k) [Ste54a]. {e^{inθ}} [Kal35a]. e^{X} [Bar64a]. γ = 3 [KF56]. H(m, a, x) = exp(−ιx)F (m + 1 − ιa, 2m + 2; 2ιx) [LH42]. H_{n}(hw) [IR52]. H_{v}(x) [MP46]. ∫_{0}^{x} J_{0}(t) dt [Sch55a]. J_{0}(x)Y_{0}(kt) − J_{0}(kx)Y_{0}(x) = 0 [LH43]. j_{0}(z, θ) [Fet57b]. J_{2}, J_{1}′, J_{2}′, N_{2}, N_{1}′, N_{2}′ [Dwi46]. J_{i_{0}(x)} = ∫_{x}^{∞} J_{0}(t)/t dt [Smi43]. J_{n}(1000) [Cor45]. J_{n}(x) [CM45]. J_{ν}(x) [Bon61]. L^{p} [HS36]. L_{v}(x) [MP46]. Λ_{0} [Fem58]. M [Kap34]. N [Cra55, Hos36b, Lip22, Amo60, Cro65, Gol58, Moo29a]. ∇^{2}u + λu = 0 [KS52]. O(4) [Sha68]. P [Hit28]. P (X) = A [Ing34]. P_{ν}^{μ}(cos θ) [Bon61]. ∂J_{ν}(z)/∂ν [LR60]. P ′_{nι}(x_{0}) [SCKH52]. R [Bur37]. R_{6} [Kap36]. R_{m+1} [Kap34]. → [Gar35, Gar37]. τ [Luk58]. θ [Bon61]. V_{3} [Kap36]. V_{m} [Kap34, Moo29c]. V_{n} [Moo29c]. x [Bon61, GD49, Sch55a]. x^{n}sechx [Chu36]. Y_{1}(mx)[xJ_{1}(x) − BJ_{0}(x)] − J_{1}(mx)[xY_{1}(x) − BY_{0}(x)] = 0 [LZ55]. z [Fet57b]. z = w^{−1} coth w [HS48]. z = w^{−1} tanh w [HS48].

-Dimension [Amo60]. -Matrices [Bur37]. -Method [Luk58]. -Regular [Cro65]. -Space [Hos36b, Lip22]. -Step [Cra55]. -Terminal [Oon50]. -tuply [Kap34]. -Way [Hit28].

1 [Hov26].

4-Poles [Dar39].

8000^{o} [Moo37].

= [Ada43].

Abel [Bar64b, EHY62]. Aberrated [Kei63]. Abridged [Sal53a, Sav38]. Abscissae [MA65]. Absolute [KTY22]. Absolutely [Pit37]. Absorbing [Fab25]. Absorption [Eis35]. According [YL56]. Accounting [TM40]. Accuracy [HW47, Lot52, Str62a, Wes68]. Accurate [Rou67]. Achieve [Fla54]. Acoustic [Hor50a]. Acoustical [HF47]. Acoustics [Pau59]. Adaptation [MS45]. Adaption [SS61]. Addition [ASC63, BM62a]. Additional [Sal61]. Adiabatic [LC34]. Adjacent [DL30]. Adjoining [Hov34]. Adjoint [Osb66, Ric26a, Coh64, Coh65]. Adjustment [Hov26, Su59]. Admittances [BP32]. Admitting [Zel23b]. Advancing [Sal44b]. Aeolotropic [Ōku52]. Aerodynamic [Tur49]. Aerofoil [Est35]. Aerofoil-Joukovsky [Est35]. Affine [Spe44a, Sta36a]. Aid [Wag51a, Wag52]. Aided [KC62]. Air [BB23]. Airfoil [Nag51]. Airfoils [Lin49, MG54]. Airship [Hov26]. Airy [Don67]. Algebra [Par49]. Algebraic [Cro29a, Fra31, Hit24, Hit38, Hit39, nL43, LU51, MF31a]. Algebraical [Sav41a, Sav41c]. Algorithm [FS55, NG47]. Alkali [Lin24]. Almost [Ava35, Cam36, Cam38, Fra26b, Fra26c, Koz63, Lan59, Muc29, Roe48, Win42]. Along [Ang59, Sal45a, Sal47a, Smi57, Kap34]. Alphabets [Fle58]. Alternative [Sal47a]. Ammonia [Bea30]. Amplitude [BL64, FR57, PS56, Red49a, Rob57a]. Amplitude-Modulated [PS56]. Amplitudes [Gri61a, Gri61b]. Analog [Fra58, Fri56, Gol46, Su59]. Analogue [BP61, BH64a, Dou59]. Analysis [Bol61, Bow56, Bow66, Cro64, Hit30, HB53, Hos36b, Kre50, KC62, Lif42a, MM53, MA65, PS56, Pen28, Ric22a, Wie26a, Wie26b, Wie29a, Wie29b]. Analytic [BY50, Bla40c, Cam38, Lev37, MS46, Sai28, Spe41b, Sta36a, Sta36b, Wie25b, You52, WYA64]. Analytical [Ike31a, Jes35, Lan38]. Analytics [Tay23]. Analytiques [BM42, Ber42]. Analyzer [Sha41a]. Analyzers [MM53]. Analyzing [Wee34b]. Anecdote [Rub64]. Angled [KP67]. Anisotropic [Kuš55, LP63, Str62b, Syn56]. Anistotropic [Sat66]. Announcement [Ano43]. Annular [EH66, HE64, Hod57, Hig42]. Antenna [BHPBT65, Kue60, Pap54, PB64, Yeh63]. Antithetic [Mor57]. Any [Dou36a, Ric22b]. Aperiodic [BW26, Han57b]. Apparent [BN62]. Appell [KD48]. Applicable [Elr58, Ste68]. Application [AA65, CW52, Cou41, Cro29a, Cro36, Cro40, FW64, GC36, Gla64, Her64, Hit25b, Hor50a, Kei62b, Kin65, KS52, Kre50, Lel67, Lit29, Mah27, Pap54, PVW66, Wie27]. Applications [Bic48, Dar39, ELN57, Fri35, Hei35, Ima49, IH49, Kli48, Law34, Phi22a, SB42, SK52, Pet65]. Approach [Cob38, Wan68c]. Approximate [Bel51, Bla40c, Cov58, Cro28, Cro29b, DS58, Gri61b, KM62, LO51, Leu51, MS50, Nag67, Red49a, Sch57]. Approximates [Rey67]. Approximation [Ald48, Cro30, Cro40, Cro59, Fra25b, HC41, Hsu68, MW52, Osb66, Sar56, SR52, Ste55]. Approximations [Bur64, Cer61, Lin41, MBC56, Zaj64]. Arbitrary [Doy58, Est35, Gui60, Mac41, Ros58]. Arc [Moo25, Smi57]. Arches [HB53]. Arcs [Sal45a, Sal47a]. Area [Cro29a, Fra35b, Rut22]. Argument [Bur66c, Gil68, Mon46]. Arguments [CM45]. Arising [Cro40, Hun60, KP53, Mez58, Zem61]. Arithmetic [Kog35]. Arithmetical [Fej34, Mah27, Wie27]. Arithmétiques [Kog35]. Army [Hov26]. Array [Mah27, Wie27]. Aspects [Che66, Zel22b]. Assigned [Dou36a]. Associated [Bic48, CT50, Cod51, Cod65, ES67, For56, Luk52a, Tsu61]. Asymmetric [Wei65]. Asymptotic [Abr49a, Ban39, Ben67, Bon61, Cra59, Di 54, Ell67, Jen60, JK58, Kan65, Kan68, Lel67, Mor54, Mow67, Mur67, Nay67, Obe55, Par54, RS66a, sNR62, Sch59, Seg60, Sle65, Smi43, Ste57, Str68, Tay39, Was44]. Atmosphere [Fab25]. Atom [GC36, Val24b]. Atomic [Lin24]. Attached [BW61, Tho48]. Attenuation [Mez58, Val65]. Attraction [Blo52]. Auto [Kre50]. Auto-Correlation [Kre50]. Automorphisms [TA63]. Autonomous [Lou61]. Auxiliary [Her64, Kap46, Kap48]. Average [Sut37]. Averages [ML59b]. Ax [Ada43]. Axi [Sim56]. Axi-Symmetrical [Sim56]. Axial [Bla40a, Bla40c, RW68]. Axially [Ber50, Cha61, Ran65, TH50]. Axisymmetric [Ber68, EH66, Par57, Ros66b].

Bôcher [MS61]. Backflow [Abr49b]. Backward [Sal62]. Band [Bar64b]. Band-Limited [Bar64b]. Bandpass [Mor65a]. Barriers [Gup60, Lew63]. Based [Fox55, Gri61b, Lin54]. Basic [Gol55]. Basis [MM53, Phi33, Red49b]. Bateman [AA65]. Beam [Hov25, ML59a, RT46, Was64]. Beams [BD60, HH25, Hil43, Hov23, Mar56]. Behavior [Ben67, BA68, CD62, KB63, LF67, Lud51, Zwi58a, Zwi58b, Zwi58c, Gri61a, Gri61b]. Behind [LR48, Ran50]. Below [Bus67, KTY22]. Bend [Eic53]. Bending [BD60, HE64, HS68, Hod57, Hov23, KW60, Rei37, Rei44b, Rei63a, Sim56, Wan68d]. Bends [Hov27, Hov28a, Hov29]. Bent [RT46]. Bessel [Coh65, Abb49, Abr48, Abr50a, Abr54, Bog51, Bon61, Cla37, Coc67, Coh64, CM45, Cor45, Don67, Dwi46, Fet57a, For58, Gol57, Hen57, Hor50b, ISC47, Jae48, Jon68, Kar49, Kir57, Kli48, Luk50, Luk52a, ME63, ME64, Mon46, Mul55, Nay66, Ng67, Obe58, Par54, Pea60, Pin57, Son65, Spi57, Str65]. Best [MS50, MB50]. Between [Bus26, Doy58, Kue51, KF56, Mul67, Pea37, Por60, Syn36, Mow67]. Bi [DG48, Tor53]. Bi-harmonic [DG48]. Bi-Parametric [Tor53]. Bicubic [dB62]. Biharmonic [BP63, Duf48, Hit35, PW54, Ran64]. Biology [Hal35]. Birkhoff [Kei63]. Block [Sal53a]. Blunted [Ell67]. Bodies [Cov58]. Body [Bag53, Cha61, LRT48, Man55]. Boiling [Pea37]. Bond [Pay64]. Bound [Cro59, Lyn66, You52]. Boundaries [BH63b, Dou36a, GS50, Hun61, MS66]. Boundary [Ald48, Bow56, BP63, BH64a, BH64b, Che61, Cob44a, Cob44b, Coh64, Coh65, Cra59, DG48, DR52, Dun67, Fis35, For54, For56, Gea64, GR55, HW67, Hsu68, Joh63, Kan62, Kue51, Mar63a, Pea68a, Pea68b, Ran64, RS66a, Roe48, Seg60, Ste57, Tho40, Tin65, Tor52, Tor53, sT47, Was53, Was44]. Boundary-Layer [Seg60]. Boundary-Value [For54, For56]. Bounded [Cam36, Kei62b, Nag67]. Bounds [BP63, Car61, DG48, DR52, Gre48, Her64, Hil60, HP60, PW54, RS57, SS68a, Sie55, Was53, Wei53]. Bouwkamp [Bla47]. Branch [Nav61]. Breakdown [JKS27]. Brittle [Bak68]. Broken [Dra60]. Broken-Line [Dra60]. Brownian [Hod52, Lou65]. Brune [Rez54]. Bubble [Zwi58b]. Bubbles [ZP54, Zwi58a, Zwi58b, Zwi58c]. Buckling [BA68, Car47, HB53, OR40, Ral56, Wei65]. Buildings [Ter29]. Built [Cov58]. Built-Up [Cov58]. Butterworth [Doy58]. Bx [Ada43].

C [Hit44, FG51]. C. [Bla47, Str32]. Calculating [Sal57]. Calculation [Cov58, Cro28, Cro29b, Dwi31, Fet54, Fet57b, Fis63, Fra55, Phi48, Sal58, Tho48]. Calculus [Bus24, Chu54, Fab36]. Cameron [Bee67]. Canonical [Lip23b, Pom68, Tie68, Uhl35]. Cantilever [Hil43, RT46]. Cantor [Sal42]. Capacitance [Dwi31]. Capacity [Bea30, Mur33, PW65]. Capillary [Bla40b]. Card [BG47]. Carleman [Eps38]. Carlo [Cur53]. Cartesian [Sal49b, Sal56, Sal60a]. Case [Car42b, Jen60, Mor65a]. Castigliano [Sti65]. Cauchy [BV65, New60]. Causality [Hop34, Hop35]. Cavitated [tW56a]. Cavitation [GS50]. Cavities [Dwi48, Fis63]. Cavity [BMC63, Kli48]. Cayley [Ric27]. Celles [Kog35]. Celluloid [HA23]. Central [Fle58, Sal43a, Sal49a, Sal63b, Sal43b]. Centripetal [Kei63]. Certain [Abr48, BM66, Blo52, For58, For56, Gil47, Goh49, Hen57, How61, Jon64a, Jon64b, Kag41, Kap36, Lou49, Man55, MS57, Mil49, Mul55, Nag64, Ng67, Nob58, Red60, Ric22a, Roe48, Ros66a, Sal54, Spi57, Wal38, Wie24a, Win37, Wyn56]. ces [Kog35]. Chain [Wu66]. Change [BH64b, Est34]. Channel [Abr49b, Hel61, Ran67]. Channels [Dia59, Yua59]. Chaplygin [Rey67]. Character [Ber67, May39]. Characteristic [FS55, For56]. Characteristics [AH64, Dar39, Fox55, Jon56, Lin54]. Characterization [Fra27a, Kar43, Tho40]. Characterized [sT46, Tsi47]. Charaktersysteme [May39]. Charge [Ros58]. Charged [CC68, Por32, Por60]. Charges [Bus26]. Chebyschev [BH63a]. Chebyshev [Sal47b]. Choked [Hel61]. Cinema [Hed41]. Circle [Bol57b]. Circles [Cam38, Gol58]. Circuit [Bur35, Red49a]. Circuits [Hof46, Jae48]. Circular [Bow56, Buc67, Buc68, GS58, Gri38, Ham52, Han62, Joh63, Lin49, Nag51, Ōku52, Ran68, RS57, RS66a, Sal45a, Sal47a, Sim66, Smi57, Wan67, Wan68b, Wei62]. Civita [Hof34]. Clamped [Mun52]. Class [Ad63, Bar61, Ben61, Gil60, LM65, Mah27, Mur67, Mur68, Noh59, Pay59, Ran64, Rei41, Ric27, Sam54, War52, War53, Wie27, Zad53]. Classes [Fra29a, Gil47]. Classical [Fra28a, SS67, SS68b]. Classification [HR67, McL46, McL47a]. Clay [TM40]. Clifford [Abr50a, Abr54]. Closed [Fra34a, Sai28]. Closure [God41, Kal35a]. Closures [BN67b]. Clouds [Zwi58b]. Coated [Yeh63]. Coaxial [Dwi48]. Codazzi [Sch41, Moo29b]. Coddington [Bil65a, Bil65b]. Coding [Fla54]. Coefficients [Ad63, Ano43, BG47, Bie52, Cam36, Ham52, Hit39, Hud50, Isa47, Kop48, Koz63, Lip24, LS43, LL43, LS44, LS45, Luk52b, Mir63, MW52, RS57, Rut22, Rut27, Rut29a, Rut29b, Sal43a, Sal44a, Sal44b, Sal45c, Sal46a, Sal46b, Sal47c, Sal48a, Sal48b, Sal49a, Sal50a, Sal56, Sal57, Sal63b, Sal43b, Sim66, Spe42c, Wie24d]. Coherency [Wee34a, Wie28a]. Coil [Cha31]. Coils [DL30]. Coincident [Hit24]. Color [Fra34b, Fra37, Kag35, Kag36, Win37]. Colorimetry [MS46]. Coloring [Sha49]. Column [Fan67]. Combination [Cro30]. Combinatorial [Gre40, Tod33]. Combinatory [Ric22a]. Combined [Dei68]. Comment [KB63]. Common [Gil47]. Communication [Fle58]. Commutativity [Zel22a]. Comparison [Dun67, KM62, ML59b, Ser54, Sha41b]. Compatibility [Sav41c, Sti65, Was57]. Complement [HR25]. Complementary [Rei48b]. Complete [Cam47, Har59a, Heu41a, Heu41b, Kap46, Tay39]. Completeness [Fry36]. Complex [Bag53, Ber41, Bur66c, Cam47, Dou40, Hit38, Hit39, Hit44, HS36, Ku26, Lin41, Lin47a, LS44, Lyo24, Moo29a, Pap54, Pea37, Sal45a, Sal47a, Sal48a, Sal49b, Sal50a, Sal50b, Sal52, Sal56, Sal60a, Sam49, Sco66, Tay23, Woo25, Zem61, Sha41b]. Complexes [BM42, Ber42]. Components [HKP36, Pea37]. Composite [Kin34, PW65]. Compound [Cox33]. Compounds [Ric26b, Ric27]. Compressed [Ōku52]. Compressible [BE47, BG47, Ber50, Han57b, Ima49, IH49, Isa47, Kuo43, LRT48, Lud51, Man55, Mil54, Mor54, PK56, TH50, tW56b, Eri55a]. Compression [TM40]. Compressive [Wei62]. Computation [Bla46, Fra58, Ng67, Sam49]. Computational [BG47, Isa47]. Computer [FS55, KC62, SS61]. Computer-Aided [KC62]. Computers [Har57]. Computing [Car65]. Concentration [Eis35]. Concentric [DS30, Hsu42b]. Concept [Krö63]. Concepts [Fle58]. Concerning [De 47, Duf48, Mar63b, Nag65, Spe42c, Tuc66]. Condition [Cob44b, Rut30]. conditioned [Val24a]. Conditions [Dun67, Hod50, Kel67, Mar63a, Sav41c, Sti65, Tho49c, Tin65, sT47, Val27, WY57, Was57]. Conducting [Lam68, MS66, Por32, tW56b]. Conduction [Bol61, EBK55, EBK57, Lot58, Low45, Rai53]. Conductors [Hig42, Man22, PW65]. Cone [She48, Sto48, Sto51, Sto52]. Cones [Ell67]. Confluent [Kar49, Mac49, SS63]. Conformal [Lip23a, Zel23a]. Congruences [Moo28a]. Conical [BHPBT65, Fel56, PB64, PB68]. Conjugate [Est34, LH26]. Connect [Cam60]. Connected [Mul56, Rey51]. Connection [Fic50, MG64a, MG64b]. Consecutive [PPW56]. Considerations [EBK55, EBK57, Was64]. Consistency [Bro50, Fry36, Tho49a]. Consolidation [Car42b, Ort30, Par57, TM40, Ter29]. Constant [Bea30, Cob40, GZ43, Hud50, Key22]. Constants [HA23, Pet57]. Constraints [Det64]. Construction [Bol57a, Gar57, GZ43, Lin57]. Contact [Blo53, De 47, Lip23b, Rob36, Zel22a, Zel27]. Containers [LF67]. Containing [Bow56, KW60, Par57, Tel53, Wan68d, Was44]. Contains [Oon50]. Continued [Cro29b]. Continuous [CM44, Gup59, Hey24b, Tor52, Val26]. Continuum [Fra29a, Krö63]. Contours [Dou31b, Dou32]. Contribution [Gei47, Gui32, Ric24]. Control [Tor52, Tor53]. Controversy [GC65]. Convection [Abr53b, Bus67, Kot68, PVW66, TY57]. Convergence [Bil61, Coo57, Dam59, Erd43, Hil52, IR52, LO51, Leu51, Rut30, Sam54, Str59, WB65, WY54]. Convergent [Kue51, Pit37, Rut27, Sal54, Sam45, Sha55, Szá49, Wyn56]. Converging [Sam52]. Conversion [Rez54]. Convolution [Cam42, Zem61]. Coordinate [Gui60]. Coordinates [Amo60, KR56]. Coriolis [HD64]. Correct [Dou56]. Correction [Fra26a, Kag40]. Corrections [Dou32, Heu41a, McL47a, Rei48a, Sto52, Tsi47]. Correlation [Kre50]. Corresponding [Mor65a]. Corrigenda [SW68]. Cosine [BH63a]. Cosmic [Ban39, Lif42b]. Cotes [Rut22]. Couche [Ric28]. Coulomb [Abr48, Abr50a, Abr54, Cob44b]. Countings [Fry36]. Couple [Coh66]. Couple-Stress [Coh66]. Coupled [Hod52]. Coupling [HF47, Hei51]. Courant [BM62b]. Covering [Sav39]. Crack [Ang59, Buc68, Che66, KW60, Wan68d, Wes64]. Cracked [Fol65]. Cracking [Bak68]. Cracks [Bow56, HS68]. Crank [Str59]. Criteria [Eni61]. Criterion [Cra53, Cro34, Dou59, Fry36, Lel67, Lev46]. Critical [Goh49]. Cross [Bas59, Hig42, Kir57, RT46, Ric25a]. Cross-section [Hig42]. Cubic [Ato59, Lin41]. Curl [Ing24]. Current [Fan67, Kue60, Kue64]. Currents [Fra25a]. Curvature [Bla40a, Bla40b, Bla40c, Cob40, Fra34a, Kap34]. Curvatures [Tho48]. Curve [Sch66]. Curved [HH25, LR48, Lip22, Ran50, Tho47, Was64]. Curves [Cob42, GZ43, Moo29a, Pea37]. Curvilinear [Bas59]. Cusped [Fra27b]. Cut [Smi57]. Cx [Ada43]. Cycle [Rez54]. Cylinder [BG48, GS58, Gri38, KM62, Lee67, Nag51, Pap49, Por60, Ran68, Wan67, Wan68b, WS45, Whi38, Yeh66]. Cylinders [Bas59, Ham52, Jon56, Kel67, Mir66, Por32, TY57]. Cylindrical [Buc67, Buc68, BA68, Cha61, CC68, Cha31, Eni61, Fan67, Hsu42a, Joh63, Kan64, Kue60, Nay65b, Pay64, Rei52a, RS57, RS66a, Sim66].

D [Hec67, Kei63, Ada43, Hit35, Hit44]. Damping [Gri61a, Gri61b, Lou55]. Dams [Wea32]. Dark [MS45]. Data [Gil25]. Debye [Bea27]. Decomposition [Ric29]. Deduction [Wie31]. Definite [Fel56, SB42, You53]. Definition [Fra28b, Fra35b, Rei51]. Deflections [EH66, Wei62]. Deformation [Ber67, Doy57, Hov27, Hov28b, Tho56]. Deformations [Eri55a, JR55, Joh63, O’M63, Rei53, Rei59, Rei61, Val66]. Degeneracy [Mor65a]. Degree [Ada43, Sal49b, SCKH52, WY54]. Delay [Noh59]. Delta [Spi57]. Dependent [Aba60, Zad51]. Deposits [Ort30]. Depressions [Bla40b]. Derivation [Cro29a, KR56, PD55, Rei63b, Sti65, Str30]. Derivative [Cro59, Dwi46, Obe58]. Derivatives [BP67, Car58, Cha34, Cow65, Cro60, Sal44a, Sal63b, Tsu61]. Derived [Dou31c, EBK55, EBK57, Kog35]. Derivées [Kog35]. Deriving [Cro28, Cro29c, Cro29b]. Design [Dar39, Gui32, Lev46, Red49a, Ste55]. Detached [Gar57, Lin57]. Detection [PS56]. Detector [Koz59]. Determinant [Moo23, Ric22b, Ric25c, Ric32, Wal38]. Determinantal [Ric24, Ric25a]. Determinants [Ced55, Hit25b, Ric22b, Ric25b, Ric26a, Ric26b, Ric27, Ric29, Ric30, Tsa54]. Determination [BE47, BG47, Ber50, Car48, Fox51, Gre48, Hov25, Isa47, Rei59, Rut22, BG47, Isa47]. Determined [Rut23]. Determining [Goh49, Mur67]. Detonators [Gar60]. deux [BM42, Ber42]. Developable [Moo24]. Developments [Kog35]. Développements [Kog35]. Devices [BG47, Isa47]. D’Hermite [Kog35]. Diagonally [BH64a]. Diametrically [Ōku52]. Dielectric [KP67, Pet57, Ste55, Yeh63]. Dielectrics [JKS27, Pet57]. Difference [BM66, Bic48, BH64a, Cra53, DR52, Dou56, Dou59, EBK55, EBK57, For54, For56, GR55, Hei40, LM65, MBC56, MW52, Pea60, Sal53a, Sal45c, Sam54, Str62a, WY54, Wie25c, Zaj64, Zel23a]. Differences [LS45, RC30, Sal43a, Sal44a, Sal44b, Sal46b, Sal49a, Sal62, Sal63b, Sal43b]. Differential [Abr53b, Ad63, Cam36, Car42a, Cha67, Cod51, Cod65, Cou41, Cro64, Cro65, DS58, FLL50, Fic50, FM30, Gea64, Gon41, Hei40, Hil42, Hud50, Hym51, Jen60, KP53, Lan51, Lan59, Lat63, Lev35, Lev43b, Lev43a, Lev49, Lou49, MBC56, MM53, Moo30, MW52, Noh59, OHK50, Pay59, Pea68a, Pea68b, Sal45c, Sal50b, Sav38, Sch42, Sha41a, SK52, Var61, Wad32, Wad35, War52, War53, Was44, Was59, Wie23, Zel23b]. Differential-Space [Wie23]. Differentiation [Hil60, Sal43a, Sal52]. Diffraction [Han62, KP67, Mil49, Obe55, Tin53]. Diffuse [Kuš58]. Diffusion [Boy61, Dou56, Fra58, Gup59, Gup60, PVW66]. Diffusion-Convection [PVW66]. Digital [FS55, Fra58, Har57]. Dimension [Amo60]. Dimensional [AP68, Bar68, BR60, Bol57b, Bus67, Car42b, CW52, Dei68, Dia59, Eni61, ES67, Gea64, Gie51, Gol46, Gri38, Her64, Kap36, LRT48, Lud51, Pau61, Ran66, Ran67, Rob36, Sch31a, Sch31b, She52, SW66, SW68, Sti65, sT46, Tsi47, tW56a, KF56]. Dimensions [BBBP64, Cer61, Cox33, J30, Lou65, Moo22b, Moo25, Moo29a, Ric22b, Ric25a, Ros64, Sav41a, Sta36a, Sta36b, Str28a]. Diode [Sco66]. Dipole [BHPBT65, God41, GK38, KM62, Kue66, Lif42a, PB64]. Dirac [Sch31a, Sch31b, Spi57]. Direct [Sal45a, Sal45b, Sal47a]. Directions [Moo29c, She52, Str28a, Tho49b]. Dirichlet [DR52, ES67, Gre48, Kel67, Mul56, PW23, Rey51, Sal53a, WY54, Wie24b, Zuc66]. Disc [Han62]. Discontinuity [Bur66b, Phi34a]. Discontinuous [Che64, Pin57]. Discrete [CK66, FW64, Fri56, Hua67, Hua68, MA65, Mul56, WY57]. Discretization [BV65]. Discussion [Gro62]. Disk [DL30, Gri61a, Gri61b, Ōku52].

Dislocation [Krö63]. Dislocations [Coh66]. Dispersive [Ben67]. Displacements [Kan65, Kan68, Phi48, Rei46a, Rei46b, Rei48a, Rei59]. Dissipative [LC54]. Distance [Ham52]. Distillation [Pea37]. Distortion [Spe41b]. Distributed [Tor53]. Distribution [Aba60, Bue47, Cam38, CK66, HS68, Hil43, Hit41, Hov34, Lar66, Nag51, Ōku52, Phi34a, Rei55, She52, Str30, Sut37, Tho49b, Wie31]. Distributions [DD61, KM62, Ros58, Syn36, WW38]. Divergent [Sha55]. Diverging [Abr49b]. Divisors [Win42]. Domain [Her64, HS36]. Domains [Ber41, Rey51, Zuc66]. Dominant [BH64a]. Doppler [LP63]. Dot [Ric25a]. Double [Rad59, Rho65a, Rho65b, Sal45c]. Drag [Col51, Wan68b]. Driving [Bru31, DeC56, Gew33]. Driving-Point [DeC56, Gew33, Bru31]. Drop [Bla40a, Bla40c]. Drops [Bla40a, Bla40c]. Dual [MF32, Mor65a, Nob58]. Dualitätstheoreme [May39]. Duality [May39]. Ducts [HF47, Hei51]. Due [Kot68, Por32, Ter29]. Duffing [Lou55]. Dynamen [Spe44b]. Dynamic [BMC63, Hey24b, HB53, Kan65]. Dynamical [Doy55, Fra30, Kan68, Lip23c, Tho46b, Val24b]. Dynamics [ELN57, Kro34, Tho46a, ZP54].

E. [Wal59]. Each [Oon50]. Earth [Fox51]. Eccentric [Hig42]. Eccentric-annular [Hig42]. Economic [Luk59]. Edge [Bow66, Bue47, HE64, Wei62]. Edges [Mun52]. Editor [How31]. Effect [BB23, Bea30, HS68, Joh30, Lew63, Man22, MG54, Phi34a, fS52]. Effects [BAC24, Dei68, LP63, Wan68d]. Eigenfunctions [Car48, Mor65a, Mor65b]. Eigenvalue [Cle57, Cra51, HP60, Nay67, Nay65b, Nay66]. Eigenvalues [Coh64, Coh65, Har59a, Her64, Mir63, Osb66, PPW56, Pay60, Sat66, Sch57]. Eight [Sal45b]. Eight- [Sal45b]. Einstein [Cob40, Phi22b]. Elastic [Ang59, Bas59, BMC63, BR60, Car47, Cla50, Con64, Eps42, Eri55a, Hov26, Hov27, JR55, JR58, Joh63, Kan65, Kan68, KR58, KW60, Mun52, OR40, Par57, Rei44b, Rei52b, Rei53, Rei61, Rei63b, RS66a, Rob57b, Set52, Sim66, Syn56, Wan68d]. Elastic-Plastic [Set52]. Elastica [Mow67]. Elastica-Viscous [Mow67]. Elastically [Sav39]. Elasticity [Bol57a, BP61, Coh66, Rei50, Sav38, Sav41a, Sav41b, Str62b, Was53]. Electric [BHPBT65, Bod34, Bur35, Fra25a, Gui32, Kue66, Lee32, Opa41, Su59]. Electrical [Ced55, Cre35, Doy55, IC44, Kro34, Tsa54]. Electrodynamic [Hov32, PD55]. Electrolytic [KN26]. Electromagnetic [AH64, Mac41, Mir66, Phi33, Pin46, Pin47, Wil57]. Electromagnetism [Soh33]. Electron [AM27b]. Electrons [Has39, Mac41]. Electrostatic [Ros58]. Electrostriction [SW66, SW68]. Element [Moo25]. Elementary [Fra26b, Gau59, Red64]. Elements [Ced57, Tel53]. Eleven [Kay55, Sal45b]. Eleven-Point [Sal45b]. Elliptic [BH64a, Cam47, Car61, Car65, CS41, Heu41a, Heu41b, Jon56, Kap46, Kap48, Kap50, MW52, New60, Yee66]. Elliptical [Che66]. Empirical [Lan38]. Employed [Fic50]. Employing [Sal63a]. Empty [Tie68]. Enclosed [KF56]. End [Pay64]. Energy [Bea27, Rei48b, Ste68, Tho56]. Ensemble [Uhl35]. Entire [Fri35, Mal36]. Entropy [Val67]. Envelopes [BN67a]. Equally [Cam38, Gue53, Por32, Sal55a]. Equally-Spaced [Gue53]. Equation [Abr53b, Ada43, BBBP64, Bar64b, Ben65, Bla40a, Bla40c, Bog51, Boy61, BH64b, Car42a, CW52, Cro64, Cro65, Di 54, Dou56, Dou59, EHY62, Elr58, EBK55, EBK57, FS55, For54, Fri56, Gil25, Gro62, Her64, Hil52, Hil42, Hof34, Hun60, Ing34, Jen60, Jon64a, Jon64b, Kal35b, KSJ22, Ku26, Lan51, LO51, Leu51, Lev43a, Lev49, LZ55, Lou55, LH43, Lyn66, Lyo24, Man22, MS66, Mer50, Mul67, Pea60, Pea68a, Phi22a, Ran64, Rey67, RW56, Ros66a, Rou67, Sal53a, Sat66, Sav37, Shi57, Smi68, Tea45, sT46, Tsi47, WB65, Wes68, Wie25c, Zel23b]. Equations [Ada43, Ad63, Arc62, BM66, Ben61, BP61, Buc67, Cam36, Cam42, Car48, Cha67, Cou41, Cra53, Cro40, Cro61, DS58, EBK55, EBK57, EH66, FLL50, Fic50, For56, FR57, Fra55, FM30, Fra31, Gea64, GR55, Gon41, Har57, Hei40, HC41, HW22, Hit23, Hit38, Hit39, HW47, Hud50, Hym51, KP53, Kli48, KR56, Lan59, Lat63, Law34, Lev35, Lev43b, Lin41, nL43, Lin47a, Lip23b, Lot54, Lot58, Lou49, LU51, Mag42, MBC56, Mil49, Mil54, MS61, Moo29b, Moo30, Mor54, MW52, Mur43, Mur68, Nob58, Noh59, OHK50, Pay59, Pea68b, PD55, PVW66, Pur53, Rei41, Ric32, Rob57a, RS56, Sal45c, Sal50b, Sam54, Sam52, Sav38, Sav41b, Sch42, Sch31a, Sch31b, Sch41, Sha41b, Sim56, SK52, SL56, Tho46a, Tie68, Tin65, Val29, Var61, Wad32, Wad35, Wag51a]. Equations [Wag51b, Wag52, Wag53, WY54, War52, War53, Was44, Was59, Woo25, Zel23a, Pet65]. Equatorial [GK38, Lif42a]. Equi [Bag53]. Equi-momental [Bag53]. Equidistant [Sal45a, Sal47a]. Equilibrating [HB54]. Equilibrium [Bla40a, Bla40c, Gil25, Pet57, Sav41b]. Equimomental [Fra29b]. Equipotential [Fra27a]. Equivalence [WW22]. Equivalent [AK64, Bur37, Cre35, DK56, Hof46]. Errata [Coh65, EBK57, Kan68, Rho65a]. Erratum [Wag52]. Error [Bel51, Fri56, Gie58, Hil60, Jan62, Kay55, Lev46, Lou49, Lyn66, MM53, Rai53, You52]. Errors [BV65, BP67, Cro59]. Establishing [KTY22]. Estimate [fS52]. Estimation [Jan62]. Euclidean [Hit30, J30, Kap36, Sta36a, Sta36b, Wil42]. Euler [EGV38, EGV38, Nav61, Nav62, Ros64]. Euler-Linien [EGV38]. Euler-MacLaurin [Ros64]. Evaluating [Ku26, Lyo24, Mor57, Rus33a, Sha41b, Woo25, Lin41]. Evaluation [Abb49, Abr53a, AM66, Hua67, Hua68, HG68, LR60, Luk55, MS57, Mez58, Tsu61, Wag51a, Wag52]. Evaporation [Cro36, GC36]. Every [Eri55a]. Exact [Con64, GP62, LO51, Leu51, Olm66, OB66, Pau61, Wil66]. Excitation [CD62, GC65, KB63, Kue66, LF67]. Excited [Kue64]. Existence [Cha67, DeC56, Hos36a, How61, Lan51, Lev43b, Val24b]. Expanding [Lam68]. Expansion [BH63a, JK58, Kar49, ME63, ME64, Ric22a, Ric25c, Sch59, Seg60, Smi43, Ste55, Wie22]. Expansions [Abr48, Abr49a, BHL63, Cra59, Dou31c, IR52, Kan65, Kan68, sNR62, Rub64, Sle65, Ste57, Str68, WW22]. Experiments [Sal63a]. Explicit [Hud50, Rut22]. Explosion [BMC63]. Explosive [Gar60]. Exponential [Bel51, BH64b, Kan64, Luk52b, ME63, ME64, Sco66, Win58]. Exposition [Lev47]. Expressed [Abr50a, Abr54, Cam56]. Expression [Hit27, Ric22b, Ric25a]. Expressions [Bel51, MS46]. Extended [Don67]. Extending [RC30]. Extension [AA65, Bol57b, GS58, Hor50a, Hos36b, Hov23, Ike31a, Jon68, Nav61, Nav62, RW68]. Extensive [Ano43]. Exterior [Ran68]. External [BA68]. Extremals [Kar43].

F [Hit35]. Faber [PW60]. Facilitate [Luk52b]. Facilitating [Sal46a, Sal47b, Sal48b]. Factoring [Hil53]. Fading [Val67]. Family [BD41, Cam60, Man55]. Far [Ber68]. Faraday [Phi33]. Few [Car58]. Feynman [Bee67, Cam60]. Field [BHPBT65, Bla40a, Bla40c, Cha61, Cha31, Dra60, Ger39, God41, Hec67, Mac41, Nag51, Por32, PB64, Ran67, Tie68]. Fields [Ber67, Fan67, Pin47, Shi54]. Fifth [Ada43]. Figures [AM27a, Spe44a, Spe44b]. File [Ric25b]. Films [Ben66b]. Filter [Dar39, Gui32, Jae48, Lev46]. Filtering [Lev47]. Filters [Bod34]. Finding [Ada43, Har59a, Hit38, nL43]. Fine [Val25]. Finite [BL64, BG48, BH64a, Bru31, Col51, Dou36a, Dou31c, EH66, FR57, Gew33, GR55, GS58, LM65, Mor65b, Oon50, Ran66, Rei53, Rei61, Rob57a, Set52, TW53, Wei62, Whe55, Wu54]. First [BP63, Car58, Che61, DG48, HP60, Kay55, LH43, Sal50b, Tay24, Wad35]. Fit [Spi64]. Fitting [Gue53, ML59b]. Five [Cro64, LH43, Str35, Zel23b, Sch31b]. Fixed [Ben61, Gol58, Gri38, KF56, Sav39]. Flat [Ald48, CD56, Che66, Hec67, HB53, Kuo53, Nag51, Tho40, TY56]. Flexure [Sav39]. Flight [Cer61]. Flow [Abr53b, AP68, Ald48, Bar68, BE47, Ber50, CW52, Cha61, Dia59, Eic53, Ell67, Elr58, Eni61, Fin55, Gar37, Gil47, GP62, Hec67, Hel61, Hil52, HD64, Ima49, IH49, Jae55, Kot68, Kue51, Kuo43, Kuo53, LR48, Lin49, Lud51, Mar50b, Mil54, Nag51, NP48, Ölç67, Olm65, Olm66, OB66, OH66, PK56, Por42, Ran50, Ran67, RW56, Sea49, She48, fS52, Sto48, Sto51, Sto52, Tho49c, TH50, TY56, Tri56, TW53, sT47, Van58, Wan67, tW56b, Wu66, Yua59, vk47]. Flowing [Syn36]. Flows [Ben66a, Eri52, Gie51, How61, Hsu68, Man55, Mar50a, Mur67, Pri49, Ran66, Ran68, Ser54, sT46, Tsi46, Tsi47, Wil67]. Fluid [Abr49b, Ben66a, BE47, Ber50, Bla40a, Bla40b, Bla40c, CD56, Gil47, GW65, HD64, Ima49, IH49, Kan62, KM62, Kuo43, Kuo53, LRT48, LF67, Mei66, Mil54, Mor54, Par57, Por42, Ran68, TH50, TY57, tW56b]. Fluid-Loaded [Kan62]. Fluids [Ber67, BG47, Isa47]. Fonctions [BM42, Ber42]. Force [Ang59, Bus26, Kei63, Olm66, OB66]. Forced [Cha67, FLL50, Lan51, WB65, Zwi58c]. Forces [HE64, Kei63, Wei62]. Forcing [Lev43b]. Foregoing [Pól52]. Form [Fet54, Mer50, Moo23, Moo29b, MF31b, Rob36]. Formal [Cha34]. Forms [Mur67, Pom68, Tie68]. Formula [Kag35, Kag36, Kop48, KCK51, Mar57, Nav61, Nav62, Ros64, Sal46a, Sal47b, Sch55b]. Formulae [DK56, Est34, LS43, Ng67, Sal60a, Spo55]. Formulas [BR43, Bon61, Cha31, Cob42, Cro28, Cro29a, Cro29c, Cro29b, Kir57, Luk55, MS50, MB50, MG64a, MG64b, Pay59, Sal45a, Sal47a, Sal49b, Sal50b, Sal52, Sal55a, Sal55b, Sal57, Sal60b, Sal61, Sal64, Sav38, Sch42, Tay39]. Formulation [BW26, MS45, Mor65a]. Foundations [IC44, JR58, Pin46, Val68]. Four [Cla37, Cox33, Fra37, Gew33, Hit30, J30, Moo22b, Mun52, Pay53, Sav39, Sta36b, Str28a, Tay23, Wil42, Win37, Zel23b, Sch31a]. Four-Space [Hit30, Tay23]. Four-Terminal [Gew33, Pay53]. Fourier [Ber29, Fej34, Hei35, HS36, Lee32, Lev35, Lif42a, Lit29, LL43, Mor65b, Rus33a, Sal57, Wie24d, Wie29b]. Fractional [CT50, Fab36, McL47b, Red60]. Frame [FS55]. Fredholm [Wag52, Car48, Tea45, Wag51a, Wag51b]. Free [BH63b, Fis63, GS50, Hel61, Ste68, tW56a, Wu66]. Free-Energy [Ste68]. Free-Molecule [Wu66]. French [Kog35]. Frenet [Cob42]. Frequencies [Dwi48]. Frequency [Bru31, SK53, Ste54a, Ste54b]. Fresnel [Moo40]. Frobenius [Ric24]. Frontal [Gil54]. Fully [tW56a]. Function [Abb49, Abr48, Bel51, Bru31, Cam38, Car58, Cla37, Cow65, Cro51, Don67, Eps39, Est34, Fem58, Fet57b, FG51, Gau59, Gew33, GD49, Gui60, HS48, Hun60, Kay55, Lev40, LH42, Lud51, Mac49, Moo37, Sal45a, Sal47a, SS63, Sie55, Spi57, Str65, Tay39, Wie24d, Win58, Yee66, Luk52a]. Functional [Gou28]. Functionals [BM62b]. Functions [Abr48, Abr49a, Abr50a, Abr50b, Abr54, Ava35, BHL63, Bar64b, Bea27, Ber41, BY50, Bla46, Bla47, Bog51, Bon61, Bou47, Bur64, Bur66a, Bur66c, CM44, Car65, CT50, Cer61, Cha57, CS41, Chu36, Coc67, Coh64, Coh65, CM45, Cor45, Cov58, Cro30, Cro34, Cro60, DeC56, Duf48, Dun67, Dwi46, Erb59, Fel56, Fel58, Fet57a, For58, Fra26b, Fra26c, Fra28b, Fra29a, Fra44, Fri35, Gew33, Gie51, Gil54, Gol57, Gol62, Gol67, Ham52, Hen57, Hor50a, Hor50b, HB54, Hos36a, Ike31c, ISC47, Jae48, Jon68, Kan64, Kar49, Kei62b, Kir57, Lan38, Lee32, Lee67, Lev37, LBA43, Luk50, Luk52b, Luk55, Luk59, Mah27, Mal36, Mar63a, Mar53, MP46, McL46, McL47a, McL47b, ME63, ME64, Mon46, MA65, Muc29, Mul55, Nav61, Nay66, Ng67, Obe55]. Functions [Obe58, Pap54, Par54, Phi34b, Pin46, Rad59, Rai53, Rho65a, Rho65b, Roe48, Rub64, Rus33b, Sal42, Sal47c, Sal53b, Sal64, Sav41a, Sav41c, Sle65, Son65, Spe42b, Spe42c, Spo55, Ste54a, Sti65, Str65, Str68, Van54, WW22, WY57, Wat63, WYA64, WS45, Wes60, Wie25b, Wie27, Wie35, Yeh66, You52, Zem61]. Fundamental [Bol57a, Fra26c, Moo23, Rut29a]. Further [Hov28a, Nav62]. Fused [HA23].

G [Hit44]. G. [Kei63]. Gamma [Cow65, Gau59]. Gaps [Mal36]. Garabedian [Lin57]. Gas [Eri52, Gar37, Key22, KSJ22, KF56, Mar50b, NP48, sT46, Tsi47, Zwi58a, Zwi58b, Zwi58c, Gar35]. Gases [Cro36, GC36, LP63, YL56]. Gauss [HG68, Sch41]. Gaussian [Sal46a, Wie31]. Gegenbauer [ASC63]. General [Bod34, Bro50, Dou36a, Fic50, Gon41, Gri61b, KR56, LC54, LC34, Ölç67, PS56, Pen28, Ric28, Ros30, Sch31a, Sch31b, SS67, SS68b, SK53, Ste54a, Ste54b, Tor53, Val25].

Generalisation [Mor57]. Generalization [KD48, Pra57, Red68, Ric24, Sch55b, Ste68, WP38]. Generalizations [Win58]. Generalized [Blo52, Cob40, Hit28, Kra59, ME64, PD55, Ran65, Spo55, TA63]. Generated [Ang59]. Geodesics [FM31, Fra35a, Moo28b]. Geometric [Fra27a, NP48, Spe44a]. Geometrical [HKP36]. Geometrie [Spe44b]. Geometries [BBBP64]. Geometry [Dou40, Lip22, Moo26, Moo29a, Moo30, MF31a, PB68, Wad35, Zel27]. German [EGV38, May39]. Gibbs [Hos36b]. Gilbert [Wal59]. Given [Car58, Mac41, Pri49]. Giving [Est34]. Glass [HA23]. Gordon [Har57]. Goursat [Dam59]. Gradient [Bur66b]. Graph [Har59a, Kag41]. Graphen [EGV38]. Graphical [Ada43, Tea45, TA63]. Graphs [EGV38, Gou58, Kag37, Kir57, TA63]. Grassmannian [Moo26]. Gratings [MS66]. Gravitation [Phi22b, Soh33]. Gravitational [Bla40a, Bla40c, Kei63, Tie68]. Gravity [Mei66]. Green [Cov58, Est34, Kei62b, Rad59]. Grid [LS44, Sal48a, Sal56]. Ground [Dwi31]. Grounded [CC68, Por60]. Group [Har59b, Oon50, Wie29a, Wil66, Zel22b]. Groups [J30, Zel23b]. Guide [Hsu42a, Kli48]. Gyromagnetic [AH64]. Gyrotropie [Nag67].

H [Hec67]. Haber [Gil25]. Hadamard [Duf48]. Half [BMC63, HG68, Olm66, OB66, OH66]. Half-Space [BMC63]. Hamilton [Lip23b]. Hamiltonian [Pau61]. Hardening [HP48]. Hardy [Ike31d, Lev40]. Harmonic [BY50, DD61, Dun67, Fra28b, Mar63a, PW54, Sav37, WY57, Wie26a, Wie26b, Wie29a, DG48]. Harmonics [Ato59, BP61, Cha34, Nay65b]. Harnack [Red68]. Having [Blo52, Fra29a, GC36]. Heat [Abr53b, Bea27, Bea30, Bol61, Dei68, Dou59, EBK55, EBK57, Fri56, Hil52, Jae55, KP53, Kot68, Leu51, Lot58, Low45, Nay67, Ölç67, Rai53, TY57, tW56b]. Heat-Conducting [tW56b]. Heat-Flow [Hil52]. Heated [Bus67]. Heaviside [Win58, Zad51]. Heawood [Kag35, Kag36]. Helicoidal [O’M63, RW68]. Helium [Gri61a, Gri61b]. Helmholtz [BH64b, MS66, Smi68]. Heredity [Gei47]. Hermite [BP67, HG68, IR52, Kog35, Kop48, Mer50, Rus33b]. Hermitian [Wie29b]. Heterogeneous [Rob57b]. Heuman [Fem58]. Heuristic [Lev47]. High [Col51, Dia59, Dou56, Eis35, Hel61, Hov29, Lot52]. High-Pressure [Hov29]. Higher [Doy57, Dut44, Hsu68, Kar43, Lin41, Luk55, Ric27, Ric30, Sal49b, Var61, Woo25]. Higher-Order [Lin41]. Hilbert [CM44]. Hitchcock [Pra57]. Hitting [Lou65]. Hole [Bow56]. Hollow [Hsu42a]. Homogeneous [Cro61, Dei68, Gea64, Hov23, Hud50, Wad35, Zel23a, MS66]. Hopf [PB68, Wil66]. Horizontal [Olm66]. Horns [BAC24]. Howard [Mos58]. Humidity [BB23]. Hydrodynamic [Di 54, GP62, How58, KC62, RS66b, Zel22b]. Hydrofoils [Wu54, tW56a]. Hydrogen [KN26]. Hyperbolic [Bol57b, Mur68, Ral56, Rey67, Sal53b]. Hyperboloid [Van54]. Hyperellipsoidal [Mor65b]. Hyperelliptic [Cro51].

Hypergeometric [Cro65, Erb59, Hen57, Kar49, Mac49, MS57, SS63]. Hypernom [Ada43]. Hyperosculatory [Sal60a]. Hyperquaternions [Moo22a]. Hypersonic [Ell67, PK56, She48, fS52, Tsi46]. Hypersonic-Flow [Ell67]. Hyperspace [Moo24]. Hypersurface [Moo23]. Hypersurfaces [Cob40, Moo29b]. Hypocycloid [Fra27b].

Ice [KTY22]. Ideal [Car42a, Rez54]. Identity [Bie52, Yee66]. II [Rei48a, Sto52, Bea30, Bla40b, Buc68, Cod65, Gar37, Gol67, Gri61a, Gri61b, IH49, Isa47, Jon64b, Rei46b, Rob57a, Ste54a, Sto51, Zwi58b]. III [Bla40c, Ste54b, Zwi58c]. Ikehara [WP38]. Imaginary [Mon46]. Imanuel [Tuc66]. Immersed [Moo29c, TY57]. Impedance [Bol57b, Bru31, DeC56, FG55, Mar67]. Implicit [Cro34, Hos36a, Var61]. Improved [MBC56]. Improvement [Hit44]. Impulsively [Wan67, Wan68b]. Incidence [Hel61]. Including [Dar39]. Incomplete [Car61, CT50, Gau59, Kap48, Tsu61]. Incompressible [Bar68, Ber67, Ber68, Kuo53, Ran67, TY57]. Increase [Bil61]. Incremental [HW54]. Indefinite [For58]. Independent [Fry36, Mag42, Sav37]. Indeterminate [Kin34]. Index [Bur66b, Spo55, Wat63]. Individual [Ter29]. Inductance [DL30, DS30, Hig42]. Inequalities [Gau59, Pay60, SS63, Zyg42]. Inequality [DD47, Jon68, Lev41, PW60, Red68, SB42, Son65]. Inextensibility [BR60]. Inextensional [JR55, Rei61]. Infinite [Bow56, Bow66, Buc68, CD56, Cod51, Cod65, EGV38, Fan67, Fle58, Fra29a, Han62, Hec67, Ike31b, Kan65, Kan68, Kin65, KM62, Man55, Mul55, Pap49, Sim66, Smi57, Whi38, Bue47, RS66a, Sal55a]. Infinitely [Kue60]. Infinitesimal [De 47, Lip23b]. Infinity [Lud51]. Influence [RS57, Sim66]. Information [Rei51]. Inhomogeneous [Che64, LP63, Smi68, SL56]. Initial [BA68, Tin65]. Injection [Tor52, Tor53]. Inputs [Shi57]. Insertion [Dar39]. Instability [HD64, Lev49]. Insulated [fS52]. Integrable [Val24b]. Integral [Abb49, Abr51, Abr53a, Bar64b, Bel51, Ben65, BG47, Bra54, CN64, Cha57, Cro40, Cro61, Don67, EHY62, Gla64, Hen57, HC41, HW22, Hit23, Hua67, Hua68, HG68, Ike31c, Isa47, Jon64a, Jon64b, Kal35b, Kar49, KR51, Lat63, LM65, Lev41, Lot54, Mag42, Man22, ME63, ME64, Mil49, Nob58, Pea60, Rei41, Rus33a, Sam52, Sat66, Sav38, Sav41a, Sav41c, Shi57, SCKH52, Spo55, Tea45, Wag51a, Wag51b, Wag52, Wag53, Wie22]. Integrals [Abr50b, Bar61, BH63a, Bee67, Bon66, Cam47, Cam60, Car61, Car65, CT50, Cro28, Cro29b, Cro30, Fel56, Fet54, Fet57a, For58, Fra30, Heu41a, Heu41b, Hor50a, Hor50b, Jae48, JK58, Kap46, Kap48, Kap50, Kay55, Lee67, Lin52, Lit29, LA43, Luk50, MS57, Mez58, Mor57, Mul55, Nag64, Ng67, Pin57, Sal61, SB42, Tsu61, Wie25c, Yeh66]. Integraph [Hed41]. Integrating [Gou28, Lot52, Sal45c]. Integration [Bir37, Bla40a, Bla40c, BR43, Coo56, Cro29a, Fic50, GD49, Ham52, Jen60, Kap52, Lot53, Lot58, LS43, LS45, Luk52b, MS50, Sal46b, Sal49a, Sal50b, Sal63a, Sal64, WYA64, Bla40a, Bla40c]. Integro [Hil42]. Integro-Differential [Hil42]. Integrodifferential [Ben61]. Interacting [BN67b]. Interaction [AP68, Has39, Kei63, Nag67, Ros58]. Interactions [BL64]. Interesting [MS57]. Interfaces [Bla40a, Bla40b, Bla40c, Wil57]. Interior [Ran68, Wes68]. Intermittent [Gol59]. Internal [Bow56, Sav41b]. Interpolating [Sal47c]. Interpolation [Amo60, Ano43, BG60, BP67, Coo56, Cro29c, Cro60, Lan38, LS44, Luk52b, MB50, Rut27, RC30, Sal44b, Sal45a, Sal45b, Sal47a, Sal48a, Sal48b, Sal49b, Sal50a, Sal56, Sal60a, Sal60b, Sal64, Sal43b, Sch66, Sch55b, dB62]. Interpolatory [Isa47]. Interpretation [De 26, HKP36, Val29]. Interreflection [Hed41]. Interreflections [Whi38]. Intersection [sT47]. Intersections [Gea64]. Interval [Fra29a, LM65, Lev49]. Intervals [Sal55a]. Intrinsic [Zel27]. Introduction [De 26, Ric30]. Invariance [Ced55]. Invariant [Cro29a, Doy55, Lou61]. Invariants [Doy57, Hit28, LC34, TA63, Zel23a]. Inverse [Cam56, Gol57, HS48, Mir66, Ric26a, Rut32, Sal44b, Sal45a, Sal45b, Sal53b, Sal58, Sal43b]. Inversion [Sal61]. Inverting [RS56, SS61]. Investigation [Sut37]. Inviscid [HD64, Syn36]. Involved [Ric22a]. Involving [Bog51, Cam40, Cha57, Fel56, Fel58, Fet57a, For58, GC65, Hen57, Hor50b, Jon68, Lee67, Luk50, Mul55, Nay66, Pom68, Spo55, Yee66, Yeh66]. Ionized [LP63]. Iron [KN26]. Irregular [Abr54, Wie26a, Wie26b]. Irreversible [Lip23c]. Isolated [Zwi58a]. Isometric [Bol57b]. Isotropic [Ber67, Eri55a, Tho56]. Iteration [Cra55, Cur53, Lot54, Wag51b]. Iterative [Gro62, Sam54, Sam49, Sam45, War52, War53].

J [Bla47]. Jacobian [Yee66]. Jet [AP68, Dra60, How58]. Jets [GS50, sNR62]. Joseph [Wie24c]. Joukovsky [Est35]. Joule [Bea30]. Joule-Thomson [Bea30].

Kasner [De 47]. Keel [Hov26]. Kempe [Gol46]. Kernel [LM65, Mor65a]. Kernels [Kog35, LM65]. Kind [Bur66c, Che61, Tea45, Wag51b]. Kinetic [Cro36, GC36, Kue64, YL56]. Klein [Har57]. Koebe [Dou31a]. Krahn [PW60]. Kron [Cob38].

L [Str32]. Labelled [TA63]. Ladder [Doy58, FG51]. Lagrange [Cro59, Rut30, Sch55b]. Lagrangean [Ano43, Cro60, Rut29a, Rut29b]. Lagrangian [BR43, Eni61, Ham52]. LaGuerre [Tay39, Kog35, Lee32, Pin46]. Lamina [ES67]. Laminar [Abr53b, Ald48, Ber68]. Landau [Ike31a]. LaPlace [Bla40a, Bla40c, BBBP64, Bar61, Cam56, Rad59, Sal53a, Sal58, Sav37, SS61]. Large [Bla40a, Bla40b, Bla40c, CM45, Fet57b, Gri61a, Mur67, Sch55a, Val66]. Lateral [Kel67]. Laterally [HB53]. Lattice [Gil54, Kei62a, Spe42a]. Lattices [Lin49]. Law [HW54, Str30, vk47, Pea37, Phi33]. Laws [Blo52, BW26, Tsi46]. Layer [Ald48, Bus67, HW67, Hsu68, Joh63, Kue51, Pea68a, Pea68b, Seg60, Tin65, Tor52, Tor53]. Layers [Cra59, Ste57]. Least [Cro61, HC41, Hov25, Lot56, MA65, ML59b, NG47, Sch57]. Lebedev [Smi68]. Legendre [BM66, Bon61, Bur66c, CT50, Chu54, Fel58, Sie55, Spo55, Tsu61, Wat63]. Legendrian [Cam47]. Length [Fra35b]. Lenses [Ste55]. Lenticular [KM62]. Letter [How31]. Levi [Hof34]. Levi-Civita [Hof34]. Levinson [Lan51]. Liapounov [ELN57]. Lichtenberg [AM27a]. Lie [Gon41, Zel23b]. Lie-Groups [Zel23b]. Liénard [WB65]. Lift [Est35, Rob48]. Light [Ger39, Kuš58, Wee34b]. Like [PW60, Wes68]. Limited [Bar64b]. Limiters [Ste54b]. Limiting [Rut29b]. Limits [For56]. Lin [Hil53]. Lindelof [Lev41]. Line [Dra60, Gar60, Har59b, Red62, Rey67]. Line-Group [Har59b]. Linear [Ad63, Amo60, Ben66a, Boy61, Cam36, CD62, Cro51, Cro30, Cro64, Cro65, Cur53, EBK55, EBK57, FLL50, Fic50, FM30, Gea64, GW65, Gui60, Hei40, Hud50, Kal35b, KP53, Koz63, KB63, Lan59, Law34, Luk52b, Lyn66, Mar63b, MW52, Mur43, Pap58, Pea68b, Pom68, Pur53, Red60, Ric32, RW56, RS56, SE59, Sch42, Sea49, SL56, Val65, Wad32, Wad35, Was59, You53, Zel23a, HE64, Kin65, Lev43a, Pet65, Sha55]. Linear-Perturbation [Sea49]. Linearization [Mil54]. Linearized [Hec67]. Linearly [Mar67]. Lines [Duf52, EGV38, Fra27a, Fra27b, Hey24a, Hsu42b, Kap34, Sha49, SK52]. Linien [EGV38]. Linkage [Gol46]. Linkages [Gol47]. Liouville [Wag52, Mul67, Wag51a]. Liouville-Neumann [Wag52]. Lipka [Wie24c]. Lipschitz [WY57]. Liquid [Ben66b, Gri38, Mow67, Syn36, TY56, Zwi58a, Zwi58b, Zwi58c]. Liquids [ZP54]. Littlewood [Ike31d]. Load [BW61]. Loaded [For54, HB53, Jen60, Kan62, Ral56]. Loading [HW54]. Loads [SS68a]. Localities [Hit41]. Locally [Dei68]. Location [Lou61]. Logarithmic [Lar66]. Lommel [Fet57a, Hor50a]. Lommel-Type [Fet57a]. Long [Ben66a, Ben66b, Kue60, Lou49, Mur33]. Long-Run [Lou49]. Longitudinal [Hov34]. Loop [Fan67]. Lorentz [Ros30]. Loss [Dar39, Gui60]. Lossless [Doy58]. Louis [Mos58]. Low [TY57, YL56]. Low-Speed [TY57]. Lower [DG48, DR52, Gre48, Her64, Wei53]. Lubrication [BH63b]. Lucas [Bar64a].

M [Hec67]. Mach [YL56]. Machine [Gou28]. Machinery [Kro34]. Machines [BG47]. MacLaurin [Ros64, Nav61, Nav62]. Macromechanics [Val27]. Macroscopic [PD55]. Magnetic [Cha61, Cha31, Dra60, God41, GK38, Hec67, Jon56, Lif42a, PB64, Ran67]. Magneto [GP62]. Magneto-Hydrodynamic [GP62]. Magnetostatic [Jon56]. Manifold [Lou61, Str28a]. Manifolds [Kap36, Sch41]. Many [Sal53b]. Mapping [Dou31a, Kag37]. Marcinkiewicz [Ben65]. Markov [Fla54, Kei62a, Wu66]. Marx [Tuc66]. Masses [Goh49]. Matched [Kan65, Kan68]. Matching [LC54]. Material [Eri55a, Fab25, Par57]. Materials [HP48, Hov23, Tho56]. Mathematical [Hal35, Han57a, Lev47, MM53, MS45, Ort30, Pin46, Sha41a, Wee34b]. Mathieu [McL47a, BHL63, Bla46, Bur64, Lat63, McL46, McL47b, Rho65a, Rho65b, Rub64]. Matric [Fra31]. Matrices [Bur35, Bur37, Ced57, Eav53, Kin34, Mar56, Pal33, PVW66, Ric26a, RS56, Wal38, Wee34a, Wie28a, You53, Zem61]. Matrix [AK64, Bar64a, Coo57, Har59a, Hit28, Ing34, Law34, Ric28, Ric32, Rut32]. Maxima [Cou41]. Maximum [CK66, FW64, Str62a, Tho44]. Maxwell [Val29]. Mean [Lev46, Spe42b]. Means [Bar64b, Dou31c, Fej34, Hov25, Isa47, Kog35, Lee32, LH26, Sav38, Sti65]. Measurable [Fra44]. Measure [CM44]. Mechanical [Cro29a, KCK51]. Mechanics [De 47, Fra28a, Fra29c, Shi54, Zel22a]. Mechanism [JKS27]. Media [Che64, Hey24b, Nag67, Syn56]. Medium [AH64, CT67, Han57b, Kuš58, Rob57b]. Melting [Bol61, KTY22]. Membrane [Ava35, HP60, Rei47]. Membranes [PW60, Rei63a]. Memory [Val67, Wie24c]. Meniscus [Bla40b]. Meridian [Bla40a, Bla40c]. Metals [Lin24]. Methane [KSJ22, KTS22]. Method [Ada43, AH64, Ben61, Bol57a, Bol61, Bol57b, BM62b, Cle57, Cra51, Cre35, Cro28, Cro29c, Cro29b, DS58, ELN57, Fox55, Goh49, Har59a, Hil53, HW22, Hit23, Hit25a, Hit29, Hit44, HW47, HKP36, Hsu68, Hua67, Hua68, JK58, Ku26, Lin41, nL43, Lin54, Luk58, Lyo24, LH26, Mag42, Mir63, ML59b, Mur67, Nay65a, New60, Pau59, Pur53, Rei48b, RS56, Rus33a, Sal53a, Sal54, Sav41a, Sav41c, Sch57, SS61, SL56, Tea45, Wie28b, Wil66, Win45, Woo25, Wyn56, You53, Ima49, IH49, Pet65]. Methods [Bil61, Cur53, DD47, Fic50, Fra28a, Her64, RC30, Sha41b, Ste55, Str62a, Var61, Wee34b]. Meyer [Mar50a, Mar50b, Ran50]. Microwave [Nag67, Ste55]. Milne [Kuš55, Wil66]. Minima [Cou41]. Minimal [Dou36a, Moo25]. Minimization [FG55]. Minimizing [BM62b]. Minimum [Tel53, Ste68]. Minkowski [Fra35b]. Mixed [Kel67, Red64]. Mixing [Tin59, TR67]. Mixtures [Pea37]. Model [Hov26]. Models [Gar60, Lin24]. Moderate [Kuo53]. Moderately [Bla40a, Bla40b]. Modification [Hof34]. Modified [Cle57, Coc67, Don67, HG68, Jon68, Nay66, Son65]. Modulated [PS56]. Modulation [SK53, Ste54a, Ste54b]. Modulus [Cam47]. Molecule [GC36, Wu66]. Møller [Coo57]. Moment [Est35]. momental [Bag53]. Monatomic [Cro36]. Monoenergetic [Mac41]. Monotonic [Sal42]. Monotonicity [Coc67]. Monte [Cur53]. Moore [Str32]. Morley [Fra27b]. Motion [CD56, Gri38, Hod52, KF56, LRT48, Lip22, Lou65, Opa41, Ste68, Wie26a, Wie26b, Zel24]. Motions [Gil47]. Movements [Gup59]. Moving [Ang59, Bus26, PK56]. Moyennes [Kog35]. Mud [Ort30]. Multiple [HR25, Hit28, JK58, Kap50, Ric25a, Sal62, SK52]. Multiplication [Ric25b]. Multiply [Rey51]. Multipole [Ran68]. Multivalents [Gei47]. Mutual [Ced55, DL30, DS30, Has39].

N [Hil53, Mos58, Wal59]. Natural [Ber35, Dwi48]. Naturally [Was64]. Navier [Mur67]. Near [HW67, Hun61, Kap52, Rey67, sT47]. Necessary [Rut30]. Negative [Fra34a, Ike31c, BH64a]. Neighborhood [Mac41]. Neighborhoods [CM44]. Neither [BH64a]. Neo [Hay55]. Net [For54, Gew33, Kap34]. Net-Work [Gew33]. Nets [PW23]. Network [Ars53, Bru31, Ced55, Doy55, FG51, Fra25a, IC44, Oon50, Red49b, Sha49, Tsa54]. Networks [AK64, BP32, Bur37, Doy58, DK56, Gui32, Gui49, Har59b, Lee32, LH26, Oon50, Pay53, RS42, Tel53]. Neumann [Wag52, Cam56, Gie58, Wag51a]. Neutron [Sat66]. Newtonian [Blo52]. Nicolson [Str59]. Nine [Sal45b]. Nine- [Sal45b]. Nodal [Duf52]. Noise [PS56]. Non [Ben66a, Boy61, BH64a, Coh64, Coh65, Cro51, Det64, FLL50, Fin55, GW65, HE64, Hov23, Jon56, Kag41, Kal35b, KP53, Kei62a, Kin65, Kro34, Lan59, Lev37, Lev43a, LRT48, Mar63b, MS66, Mil54, Moo27, Osb66, Pap58, Pea68b, Rei52a, RW56, Sha55, SCKH52, SK52, Val24b, Val24a, WYA64]. Non-analytic [WYA64]. Non-conditioned-periodic [Val24a]. Non-Homogeneous [Hov23, MS66]. Non-Integrable [Val24b]. Non-Integral [SCKH52]. Non-Linear [Ben66a, Boy61, Cro51, FLL50, GW65, Kal35b, KP53, Lan59, Mar63b, Pap58, Pea68b, RW56, HE64, Kin65, Lev43a, Sha55]. Non-Magnetic [Jon56]. Non-negative [BH64a]. Non-Riemannian [Kro34, Moo27]. Non-self-adjoint [Coh64, Coh65]. Non-Separating [Kag41]. Non-simple [Det64]. Non-Stationary [Kei62a]. Non-Steady [LRT48, Mil54]. Non-Uniform [Fin55, Rei52a, SK52]. Non-vanishing [Lev37]. Nonintegral [Spo55]. Nonlinear [Arc62, Ben65, BN62, Ben67, BN67a, Ber67, Lot54, Mei66, Mur68, Nay65a, Noh59, Ros66a, Sim56]. Nonstationary [Shi57]. Nonuniqueness [OH66]. Nor [BH64a]. Normal [Dei68, Gui60, Moo22b, War52, War53]. Normalization [Chu36, Tsu61]. Not-So-Slender [Wan68c]. Notched [Lin47b]. Notches [Bow66]. Note [AK64, Ars53, Bag53, Bla47, Blo53, Bra54, Bur66a, Bus24, CK66, Cov58, Cra59, Deb25, Di 54, Eav53, Eis35, Elr58, Fra37, Gla64, Gue53, Hey24a, Hil53, Hil42, How61, Hua68, Ike31b, Ike31d, Ing24, Kag35, Kan64, Kar49, KR58, Ku26, Kuh57, Lan51, Lan59, Lem25, Lin57, Lot56, Lyo24, ML59a, McM39, Mon46, Moo22b, Moo23, Moo24, Moo27, Moo28a, Nag64, Olm65, Pay60, Phi22b, Rei44a, RT46, Rei47, Rei48b, Rei61, RW56, Rob48, Roe48, Ros30, Sal46b, She52, SCKH53, Spe42b, Sta36b, TR67, Val24a, Val29, Wal38, Wan68b, War53, Was57, Whe55, Wie25a, Wie25b, Woo25, Wyn56, Yeh66, Zaj64]. Notes [Luk50, Sha68, Val24b, Zyg42]. Notions [Wie24a]. Noyaux [Kog35]. Nozzle [TH50]. Number [Dou36a, Gil68, Mur67, RS42, Tel53, YL56]. Numbers [Erb59, Ike31a, Kuo53, McM39, Sut37, TY56]. Numerical [Arc62, BHL63, BG47, BY50, Dam59, DR52, EHY62, Eni61, FR57, Fra58, Gar57, Gie58, GD49, Hil52, Hil60, Hit23, Hud50, Hym51, Isa47, Jae55, Kap52, Lin47a, Lin57, Lot58, Lou49, LS43, LS45, Lyn66, Mag42, New60, OHK50, Rob57a, Rou67, Sal43a, Sal46b, Sal50b, Sal52, Sal58, Sal63a, Ste55, Tea45, Wag51a, Wag52, Wag53, WYA64, You52, Bla40a, Bla40b]. Numerous [Hit41].

O. [Wie25a]. Obey [Pea37]. Obreschkoff [Sal64]. Obreschkoff-Type [Sal64]. Observations [Gue53]. Obstacle [BE47]. Obtained [Sch57]. Obtaining [Cro60, Sal53b, Sal63b]. Occurring [Abr53b, Obe55, RS66b, Sco66, Shi57, Bar61]. Ohmic [Oon50]. One [Cer61, Che61, Cro34, GC36, Her64, HR25, KF56, Oon50, Spe42b, sT46, Tsi47, Kei63, Wie27]. One-Dimensional [Her64, sT46, Tsi47, KF56]. One-valent [Spe42b]. Only [Oon50]. Onsager [Val67]. Operating [Doy58]. Operational [BM62a, Bus24, Chu54, Pen28]. Operator [Bar64a, BG47, Isa47, Mor65b, Mur43, Ran64, RS66b]. Operators [Bic48, Cur53, Fra29c, LM65, Mul56, Osb66, Pom68, Zad51, Zad53, Pet65]. Optical [MS66, Par49]. Optimization [Shi57]. Orbits [Ban39, God41, GK38, Lif42a, Lif42b, Val24b]. Order [Bla47, Bou47, Cha67, Coc67, Cro64, DD61, Dav57, Dou56, Doy57, For54, HR67, Kar43, Kar49, Lan51, Lev43b, Lev43a, Lin41, McL47b, Obe58, Pap54, Sal45c, Sal50b, Str59, Str65, Tsu61, Var61, Wad32, Wad35, Woo25, Cro65]. Ordered [Hit25b, Ric25b]. Orders [CT50, Zel23b]. Ordinary [Cro64, Cro65, Hov23, Hua67, Lev35, Lev49, Pea68b, War52, War53, Was44]. Ordinate [Hua68]. Orientable [Kag41]. Oriented [SS61]. Originating [Bow56, Sal61]. Orthogonal [Fra29a, Kap34, KR56, Pal33, Sal61, WW22]. Orthogonality [Rho65a, Rho65b, Van54]. Orthotropic [Bak68]. Oscillating [Gri61a, Gri61b, Tur49]. Oscillation [Cro51, Nay65a, PVW66]. Oscillations [Pap58]. Osculatory [Coo56, Sal55b, Sal56]. Oseen [Olm65, Olm66, OB66, OH66]. Oval [BE47]. Oval-Shaped [BE47]. Overdifferentiation [Sal63a]. Overly [Zaj64]. Overly-Stable [Zaj64]. Overvoltage [Kno26]. Own [Ral56]. Oxygen [Eis35]. Ozone [Fab25].

P [Hit35]. Padé [Luk58]. Paper [Ber42, Bla47, Cra59, Dou32, Hit29, Hua68, Ike31d, Lin57, McL47a, Mos58, Pól52, Sto52, Tay24, Tsi47, Tuc66, War53, Wie25a, Wie26b, Rei48a]. Papers [Str35]. Parabola [ML59b]. Parabolic [Lee67, LF67, Mur68, Nag51, Var61, WS45, Yeh66]. Paraboloid [Van58]. Paraboloidal [Pin46, Pin47, Ral56]. Parallel [AP68, Dia59, Dra60, Har59b, HD64, Man22, Moo28b, Por32, Por60, RS42, Syn36, Tin59, TR67, Zuc66]. Parallelism [Lip23a]. Parameter [HD64, Was44]. Parametric [CD62, GC65, KB63, MF31b, Tor53]. Part [Kei63, Mah27, Wie27, Zwi58a, Zwi58b, Zwi58c]. Partial [Ad63, BH63b, Cra53, Fej34, For54, FM30, Gea64, GR55, Hym51, KP53, Lev35, OHK50, Pay59, Var61, Wad32, Wad35, Zel23b]. Partially [Bue47]. Particle [Mez58, Opa41]. Particular [Cow65, Fis35]. Partitioning [AK64, Her64]. Past [BE47, CD56, Cha61, Dia59, Gri38, Hec67, Hel61, Kan62, Kuo53, Man55, Olm66, OB66, OH66, PK56, Sto48, Sto51, Sto52, TY56, TW53, Wan67]. Paths [HW54]. Pattern [Hel61]. Patterns [Ber50, Gil47]. Pekeris [AA65]. Penetration [CT67, KN26]. Perfect [Kag38, Moo25]. Perfectly [Eri55a, MS66]. Periodic [Ava35, BW26, Cam36, Cam38, Cha67, FLL50, Fra26b, Fra26c, God41, GK38, Lan51, Lan59, Lev43b, Lif42b, Lou55, Muc29, Roe48, Sal64, Val24a].

Periodicity [Win42]. Permanent [BL64, Zwi58a, Zwi58b, Zwi58c]. Perron [Wie25a]. Perturbation [Det64, Fox55, Jan62, Lin54, Mir63, Nay65a, Ran66, Sea49]. Perturbations [Mur68, tW56b]. Perturbed [Lou61]. Perversion [J30]. Pfaffian [Moo30]. Pfaffians [FM31, MF31a, MF31b, MF32]. Phase [BH64b, JK58, KSJ22, Red49a]. Phases [Win45]. Phenomena [BW26]. Phonograph [BAC24]. Photoelastic [HA23]. Photometry [MS46]. Phragmen [Lev41]. Physical [Bla40a, Bla40b, Bla40c, Cro40]. Physically [MS57]. Piecewise [Amo60]. Piecewise-Linear [Amo60]. Pipe [Hov27, Hov28a, Hov29, Kuo43]. Pipe-Bends [Hov27]. Pipes [Hov28b]. Pitching [MG54]. Place [Bea27, Cla37]. Places [Sal53b]. Planck [Moo37]. Plane [Ber67, Bue47, Cob44a, Cob44b, Dou40, ES67, Gil47, Gol46, GK38, Hil42, Hod50, Hov28b, Kar43, Kot68, Lif42a, LS44, Man55, Mar67, Mar50a, Mar50b, NP48, Olm65, Olm66, OB66, OH66, Sal50b, Sal52, Tay23, Tho49c, Wil57]. Planes [Moo22b, Por60, Syn36]. Plasma [Has39, Kue64, Kue66]. Plastic [Car42a, HP48, Hod50, Hod57, Phi48, Set52, Tri56]. Plasticity [Cob44a, Cob44b, Eri55b, Krö63, SS67, SS68b]. Plate [Ald48, Bow56, Bue47, CD56, Duf52, EH66, HE64, HS68, Hec67, Hod57, KW60, Kuo53, Lin47b, Mun52, Nag51, Pay60, Smi57, TY56, Wan68d, Wes68]. Plate-Like [Wes68]. Plateau [Dou31a, Dou31b, Dou32, Dou36b]. Plates [BR60, Con64, Eps42, Low45, OR40, Rei37, Rei44b, Wei62]. Point [Ben61, BR43, Cro34, DeC56, Gew33, HR67, KTY22, Lev49, Pau59, Pea37, Sal45b, SS68a, Was59, Bru31]. Points [Cro64, Cro65, Hit24, Nay67, Rut23, Spe42a]. Pointwise [BP63]. Poisson [Hun60, Str30]. Polar [Sal50a]. Polarized [Wee34b]. Poles [Dar39, LC54]. Polyadic [Hit27, Hos43]. Polyadics [HR25, Hit25b, Ric25a]. Polygons [Dou40, Gol51, Gol55]. Polyhedra [Gre40]. Polyhedral [Fra35a]. Polynomial [Ada43, Cro40, Cro59, HC41, Rut22, Rut23, Sco66]. Polynomials [Bar64a, Cro29a, Gue53, Hil53, Hit35, IR52, Rut30, Sal61, Sav38, Str62a, Tay39, Tsu61, Wie29b]. Polytopes [Cox33]. Polytropic [Mar50b]. Popov [Lel67]. Porous [Bar68, HW67, Par57, Yua59]. Position [Fox51, Uhl35]. Positive [Fra34a, Gol62, Gol67, MW52, You53, Zem61]. Positive-Real [Gol67, Zem61]. Possessing [Kap34]. Possible [Eri55a]. Post [BA68]. Post-Buckling [BA68]. Postulates [Ars53]. Potential [CC68, Est35, Lud51, Phi34a, Por60, Val68, Wie24a]. Potentials [Ran65]. Poussin [Dou31c, Rut32, RD35]. Power [BHL63, Fra29a, LH26, Rub64, Sim56]. Prager [Kuh57]. Prandtl [Mar50a, Mar50b, Ran50]. Preceding [Mos58]. Prediction [Car58, Lev46, Lev47]. Prescribed [Bru31, Dar39, Gew33, Red49a]. Presence [Cha61, Gup60, Hec67]. Pressure [Bea30, BA68, Hov29, Wea32]. Pressurized [Rei63a, Wes64]. Primary [Ban39, Lif42b]. Prime [Sut37, Win42]. Principal [Hey24a, Lif42b, Moo29c, Str28a]. Principle [CK66, FW64, HP48, Hov25, Sti65]. Principles [Ber35]. Probabilities [Lou65]. Probability [Ber35, Hop34, Hop35, Jes35, Str35, Uhl35]. Problem [Abr53b, AA65, BP63, BH64a, BH64b, Che61, Cob44a, Cob44b, Con64, Dam59, DG48, DR52, Dou31a, Dou31b, Dou32, Dou36b, ES67, FW64, Fis35, Fra34b, Fra37, Gie58, Gre48, Han57a, HH25, Hil42, Kar43, Kuh57, Kuš55, Low45, Mir66, Mul56, Nay67, Nay65b, New60, Ölç67, OR40, PW23, Pra57, PB68, Rai53, Ral56, Rey51, Sal53a, Seg60, Sle60, Spi64, SK53, Ste54a, Ste54b, Tho40, Tod33, WY54, Was59, Wei59, Wie24b, Wil66, Win37, YL56, Zuc66, Zwi58c]. Problems [BV65, Bol61, Buc67, Buc68, CK66, Cle57, Coh64, Cra51, Cro36, Cro40, For54, For56, Fra58, GC36, Gil60, GR55, GC65, GP62, Hal35, Hed41, Her64, Jes35, Mar56, Mez58, Nay65a, Nay66, PW54, Pom68, Ran64, Red64, RS66a, Roe48, SW66, SW68, Str62b, Was53, Was44, Coh65]. Procedure [Cro60, Cro61, HC41, Lot52, Lot53, Str59]. Procedures [Cra55, War52, War53]. Process [Ort30, Sam45]. Processes [KR60]. Produce [Dar39]. Product [Bir37, Gou28, Hit41, Kin34, Kir57, Rai53, Ric22a, Ric26b, SK53, Ste54a, Ste54b]. Products [Cha57, Hit27, Hos43, Ike31b, Kin65, Ric25a, Ric27]. Profiles [BD41]. Progressing [Lew63]. Progressive [Ter29]. Projection [Sar56]. Projective [Moo30]. Prolate [Mar53, Sle60, Sle65, Str68, Yeh63]. Proof [Eps38, Hov23]. Proofs [De 47]. Propagation [AA65, AH64, BN67a, Bur66c, Che64, Fox55, Han57a, Han57b, Kan62, Pau59, Rob57b, Val65]. Properties [Bea30, Bla40a, Bla40b, Bla40c, Bon66, Cam36, Cro29a, Fej34, Fem58, Gil54, Hov26, KTS22, Mac49, Mah27, NP48, Par54, Ran65, Rho65a, Rho65b, Wie27]. Property [Blo52, Hey24a, Pet57]. Pseudo [Hey24b]. Pseudo-Continuous [Hey24b]. Publication [Ano43]. Punch [BG47]. Pure [Rei63a]. Purely [Mon46].

Quadratic [Cam42, Fra30, Koz59, Syn36, Wie24d, Zel23b]. Quadrature [BY50, HG68, Kan64, Kop48, Sal45c, Sal46a, Sal47b, Sal55a, Sal55b, Sal62, You52]. Quadratures [Cam40, KCK51]. Qualitative [Fra28b]. Quanta [Str28b]. Quantics [Hit25a, Hit29]. Quantization [BW26, Val24a]. Quantum [De 26, Fra28a, Fra29c, Jan62, Wie28a]. Quartic [Ku26, Lin47a, Lyo24, Sal48a, Sha41b]. Quartz [HA23]. Quasi [Gea64, Lyn66, Wie25b]. Quasi-Analytic [Wie25b]. Quasi-Linear [Gea64, Lyn66]. Question [Duf48]. Queuing [Kei62b]. Quintic [Ada43, Sal48a].

R [FG51]. R-C [FG51]. Raabe [Bar61]. Racah [Bie52]. Radial [Bow56, Bur64, Jae55, sNR62]. Radiating [Aba60]. Radiation [Ban39, BHPBT65, BG48, Hor50a, Kei63, Kel67, Pap49, Par54, PB64, PB68]. Radiations [Fab25]. Radiative [Gro62]. Radius [Bla40a, Bla40b, Bla40c]. Radix [Sal53b]. Ramanujan [Bra54]. Random [BN67b, Car58, CD62, Cer61, Che64, CT67, GC65, Kei62b, KR60, Koz63, KB63, Wie35]. Range [HG68, RD35]. Rank [Hit28, Ric32]. Ranks [Ric28]. Raoult [Pea37]. Rapidly [Sam52]. Rate [Bil61]. Ratio [PPW56]. Rational [Ing34]. Ray [Val26]. Rayleigh [DD47, Kan62, YL56]. Rays [Lif42b, Deb25]. RC [Gui49]. RC-Networks [Gui49]. Reactance [Dar39]. Reaction [Gar35, Gar37]. Reactions [Kei63]. Reactor [ELN57]. Real [CM44, Gol62, Gol67, Ike31c, Lin41, Zem61]. Reciprocal [Pom68]. Rectangle [Gie58]. Rectangular [Con64, Mun52, RT46, Sav39]. Rectifiers [Ste54b]. Rectilinear [Hey24a]. Recurrence [Mar53, Sch42]. Recursive [Ng67]. Reduced [Wes68]. Reducible [Eav53, Tho46b]. Reduction [Che61, Dav57, HR67, Har59a]. Reduction-of-Order [Dav57]. Reductions [Win37]. Redundancy [Fla54]. Reflection [Bur66b, Kuš58, Tin53, Wil57]. Reflections [Moo40]. Reflector [Pin46, Pin47]. Refraction [Wil57]. Refractive [Bur66b]. Region [Bow66, Mor65b, Ran66, Rey67, Wes68]. Regions [Fra34a]. Regression [Su59]. Regular [Abr54, Amo60, Bas59, Cox33, Cro64, Cro65]. Related [Bar61, BM66, BH64b, Eps39, Han57a, Lat63, LBA43]. Relating [DK56, Gau59, Jes35, Lev37]. Relation [Pea37, Red62]. Relations [Bro50, Ced55, Cha57, KR58, Mar53, Rei52b, Tho49a, Tho56, Val67]. Relationship [Mul67]. Relativistic [De 26, Str28b]. Relativity [Kei63, Sch31a, Sch31b, Val25]. Relaxation [Cle57, Cra51]. Remark [Ber42, Mar63b, Rei37, Tuc66]. Remarks [Hei51, Hop35, Hym51, Mos58, Pól52, Red49b, Red64, Spe42c, Win45]. Rengels [Spe41a]. Repeated [Jae48, Kay55, KCK51, Ran64, Sal49a]. Representation [BM62a, Gil25, Hen57, KR60, Pay59, Pea60, Spe44b, Tay23]. Representations [Don67, Luk59, Sha68]. Reproduction [Joh30]. Repulsion [DL30]. Research [Hov28a]. Resistance [Oon50]. Resistances [Doy58, Dwi31]. Resolvent [LM65]. Resonances [BN62]. Respect [Coc67, Obe58, Tsu61]. Response [BMC63]. Rest [Wan67]. Resulting [For54]. Results [BHL63, Dou36b, Han57a]. Retarded [Kei63]. Reverberation [Joh30]. Revolution [Arc62, Lar66, Rei47, Van58]. Reynolds [Gil68, Kuo53, Mur67, TY56]. Rhombic [HP60]. Ricci [Lip24, Sch41]. Richards [Gol62, Gol67]. Richardson [You53]. Riemann [Tho40]. Riemannian [Kro34, Moo26, Moo27, Moo28a, Moo29a, Moo29c, Sch41, Str28a]. Right [Ham52, KP67]. Right-Angled [KP67]. Rigid [Bag53, Hun61]. Rigidity [Mar57, Wei53]. Rings [Car47]. Ritz [DD47, Hov32]. Rivet [BW61]. Rivet-Attached [BW61]. Rods [Rei52a]. Root [HW47, Lev46]. Root-Squaring [HW47]. Roots [Ada43, Bog51, CT50, Dwi48, Fra55, Hit38, Hit44, Ku26, Lin41, nL43, Lin47a, LZ55, LU51, Lyo24, Sam49, Sha41b, Wat63, Woo25]. Rotating [GW65, HD64, Kro34, Mow67]. Rotation [J30, Lip24]. Rotational [Mar50a, Mar50b, Sea49]. Rotationally [O’M63]. Rotations [Hit30, Sta36b]. Rotors [Gol51, Gol55, Gol58, Gol59]. RS [Hov26]. RS-1 [Hov26]. Ruled [Wil42]. Run [Lou49].

S [Hit35]. S. [Hil53]. Saddle [Pau59]. Sails [Bar68]. Saint [Sav41c]. Salzer [SS61, Wyn56]. Sampling [Bar64b, Kra59]. Satisfied [Bie52]. Saturated [Car42a]. Scalar [Han62, Sle60]. Scale [KTY22]. Scattering [Deb25, Kuš55, Kuš58, Lam68, Mar67, Mir66, Red62, Sle60]. Schlicht [Spe42c]. Schlomilch [Gla64, Hay55]. Schrödinger [Her64]. Schroedinger [Har57]. Schwarz [DD47]. Science [Ber35]. Scott [Ric26b]. Second [Ald48, Bur66c, Cha67, Cro64, Cro65, ELN57, For54, HR67, Hit29, Lan51, Lev43b, Lev43a, Moo23, Sal45c, Sal50b, Tea45, Wag51b, Wie26b]. Second-order [Cro65]. Secondary [Eic53, TM40]. Section [RT46, Hig42]. Sections [Bas59, Mar57]. Sedenions [Hit30]. Sehrödinger [Rou67]. Self [HB54, Kot68, Osb66, Coh64, Coh65]. Self-Adjoint [Osb66]. Self-Convection [Kot68]. Self-Equilibrating [HB54]. Semi [Bow66, Bue47, CD56, Fri56, Hec67, KM62, Man55, RS66a, Sal55a, Sim66]. Semi-Discrete [Fri56]. Semi-Infinite [Bow66, CD56, Hec67, KM62, Man55, Sim66, Bue47, RS66a, Sal55a]. Semiinfinite [Kuš58]. Sensitivity [Lot56]. Separating [Kag41]. Separation [Lev49, Val24b]. Sequence [Mar63b, Rut30]. Sequence-to-Sequence [Mar63b]. Sequences [Sha55]. Sequential [BN67b]. Series [BHL63, BM66, Bil61, Cam56, Car61, Cha34, Eps39, Erd43, Fej34, Gla64, Har59b, Hay55, Hen57, ISC47, Kog35, Lit29, MS61, Obe55, Pit37, RS42, Rub64, Sal54, Sim56, Spi57, Szá49, Wag51a, Whe55, Wyn56, Kog35, Wag52]. Series-Parallel [Har59b, RS42]. Serrin [Red68]. Serving [Cam60]. Sessile [Bla40a, Bla40c]. Set [BV65, Chu36, Cro34, Kal35a, MA65, Tay39]. Sets [Eav53, Str28a]. Settlement [Ter29]. Seven [BR43]. Seven-Point [BR43]. Several [Hit41]. Sextic [Sal48a]. Shallow [JR55, Lar66, Rei46a, Rei46b, Rei48a, Rei59, Rei61, Sim56, Wei59, Wei65]. Shaped [BE47]. Shear [CW52, Che66, Dei68, Elr58, RT46]. Shearing [Tho44]. Sheath [BHPBT65, PB64]. Sheet [BW61, Mei66]. Shell [Fol65, KR56, Lar66, Ral56, Rei47, Ros66b]. Shells [Arc62, Buc67, Buc68, BA68, Cla50, Eps42, Jen60, JR55, JR58, Joh63, KR58, O’M63, Pay64, Rei46a, Rei46b, Rei48a, Rei52b, RS57, Rei59, Rei61, Rei63b, RS66a, RW68, SS68a, SS68b, Sim66, Sim56, Tao59, Wan68a, Wei59, Wei65]. Shock [Bro50, Gar57, KF56, Lin57, Ran50, She52, Tho47, Tho48, Tho49a, Tho49c, Tho49b, sT47]. Shock-Motion [KF56]. Shocks [LR48, Tin53]. Short [DS30, Hor50b, LH43]. Shrink [Spi64]. Sides [Sav39]. Significance [MA65]. Similarity [Gil68, Tsi46, vk47]. Simple [Car42b, Kal35b, Luk55, Mah27, Mur67, Pea37, Sal54, Sal63a, Wie27, Det64]. Simplest [Moo29b]. Simplicial [Amo60]. Simplified [EH66]. Simson [Fra27b]. Simultaneous [Car58, Pur53, SR52, SL56]. Simultaneously [Eav53]. Sine [BH63a]. Single [Tay24]. Singular [Cro64, Cro65, Eri55b, Gea64, Jon64a, Jon64b, Mir63, Mur68, Rei41, Sal42, She52, Tho49b, WW38]. Singularity [Kap52, Kot68, Nav61, Olm66, OB66]. Sitter [Lem25]. Six [Bea27, Fra34b]. Sixspace [Kap36]. Sixteen [Wee34a]. Sixth [Woo25]. Size [Bla40a, Bla40c]. Sizes [Rei55]. Skeleton [Eav53]. Skin [Man22]. Slab [Sav39]. Slender [Ell67, LRT48, She48, Wan68c, Was64]. Slices [Rei55]. Sliding [Car42a]. Slightly [Fin55, Sto48, Sto51, Sto52]. Slit [Zuc66]. Slot [Pap49]. Slowly [Sal54, Sha55, Szá49, Wyn56]. Small [Bon61, Gri61a, Rei46a, Rei46b, Rei48a, TY56, tW56b, ZP54, Zwi58a, Zwi58b, Zwi58c]. Smooth [BG60]. Soil [Car42a, Shi54]. Soils [Car42b]. Solenoids [DS30]. Solid [Rei55, sT47, Gar35, Gar37]. Solidification [Bol61]. Solids [Aba60, Val65]. Soluble [BBBP64]. Solution [Abr53b, Arc62, AH64, Bar64b, CW52, Con64, Cro40, Dam59, DG48, DR52, Dou56, DS58, Elr58, FR57, Gie58, Gol57, Gre48, Gro62, Har57, Hei40, Hit23, Hsu68, Hud50, Hun60, Hym51, KP67, LO51, Leu51, Lev35, Lin47a, Lot54, Lou49, Lyn66, MS66, Mer50, Mow67, Mur43, New60, OHK50, Olm66, OB66, PB68, Ric32, RW56, Rob57a, Roe48, Sal53a, Sav37, Sav38, Seg60, Sim56, Sle60, Smi68, SK53, Ste54a, Ste54b, Tea45, Tor53, Wag51b, Wag53, WY54, Was44, Wie25c, Wil66, Zuc66, Pet65]. Solutions [Ad63, Bol57a, Boy61, BH64b, Cam36, Cha67, Cro65, Di 54, EHY62, Eni61, FLL50, Gea64, GP62, Gri61b, Hil52, Ing34, Lan51, Lan59, LO51, Leu51, Lev43b, Lot56, Lou55, MS61, Mor54, Mul67, Mur67, Nay67, Pau61, Pay59, Pea60, RS66a, Rey67, Ros66a, Ros66b, Rou67, Sam54, Sam52, Sav41b, WB65, WY54, Was53, Wes68]. Solved [Rob36]. Solving [Cro61, HC41, HW47, LH26, Mag42, Pur53, RS56, SL56, Var61, War52, War53, You53]. Some [BBBP64, Bea30, Bog51, Bon61, Bon66, Boy61, Buc67, Buc68, Bur64, Car61, Cha31, Che66, Dou36b, DK56, Fel56, Fel58, Fem58, Fic50, Gar60, Gau59, GP62, Hal35, Hei51, Hor50a, Hor50b, ISC47, Jes35, Kan64, Kli48, Lat63, Luk50, Luk55, Nay66, NP48, Pau61, Pin57, Rad59, Ran65, Ran66, Rho65a, Rho65b, Sal63a, Sav41b, SB42, Sle65, SW66, SW68, Spe42c, Was64, Zel22b, Bic48]. Sommerfeld [Pap54, Val25]. Sonic [Fis63]. Sound [BB23, BG48]. Sounds [Joh30]. Source [Kot68, Kue64]. Sources [Hit41]. Souriau [FS55]. Souriau-Frame [FS55]. Space [BMC63, Ben65, Bol57b, CM44, Cob42, Dut44, God41, Hit30, Hos36b, J30, Kan65, Kan68, Lip22, Moo22b, Moo26, Moo27, Moo28a, Moo29a, Moo29c, Spe44a, Sta36a, Sta36b, Tay23, Tie68, Wie23, Wil42]. Spaced [Cam38, Gue53]. Spaces [Cob40, Gou58, Tho40]. Span [Wu54]. Special [BG47, Dar39, Hun60, Isa47, Jen60]. Specific [Bea27]. Spectra [Eis35]. Spectrum [Mah27, RS66b, Val26, Wie27]. Speed [TY57]. Speeds [Col51, Dia59, Hel61, MG54, PK56, Rob48]. Sphere [CC68, Lam68]. Spheres [Mow67, Rei55]. Spherical [Ato59, BM62a, BMC63, BP61, DD61, Fol65, FR57, Gol51, Gol55, Rei46a, Rei46b, Rei48a, Rei59, Rob57a, Sim56, Wei59, Wei65]. Spheroid [Sle60]. Spheroidal [Abr49a, Bla47, Bou47, Bur66a, Cha57, CS41, Mar53, Rho65a, Rho65b, Sle65, Str68, Yeh63]. Spin [GW65, Pau61]. Spin-up [GW65]. Spinor [Pay53]. Spiral [Kue51]. Spline [Sch66, dB62]. Square [Cro61, HC41, Lev46, LS44, Moo25, MA65, Sal48a]. Squares [Lot56, ML59b, NG47, Sch57]. Squaring [HW47]. Stability [Ben61, Bil65a, Bil65b, Bus67, Cod51, Cod65, Cra53, Dam59, Di 54, Dou59, Dra60, Eni61, EBK55, EBK57, Fra58, GS58, How58, Koz63, KC62, Lel67, Lev49, Lif42b, Mar56, Mor54, Mun52, RS66b, Sco66, Syn36, Wei59, Wil67]. Stable [FLL50, Hun61, Rou67, Var61, Zaj64]. Standing [BN62]. Standpoint [Val25]. Star [Wes64]. Started [Wan67, Wan68b]. State [KSJ22, Phi22a, Str62b, sT46, Tsi47]. States [Hov26]. Static [Mac41]. Stationary [Cro34, JK58, Kei62a, Win45]. Statistical [Val29]. Statistically [Ste68]. Statistics [Hop34, Jan62]. Steady [Gri38, Kel67, LRT48, Mar50b, Mil54, Ste68, Str62b, Tho49c]. Step [Cra55, Pay64]. Stewartson [Cra59]. Stiffened [Bue47]. Stiffener [BW61]. Stirling [Dou31c, Erb59]. Stochastic [SE59, Zad53]. Stokes [Cha61, Mur67, Ran64]. Straight [Kuo43, Lin49]. Strain [Dei68, HW54, HP48, KR58, Rei52b, Tho56]. Strain-Hardening [HP48]. Strata [Ter29]. Stream [Fra27a, Gie51, Hel61, TY57]. Streamline [tW56a]. Streamlines [Fis63, Pri49]. Streams [Tin59, TR67]. Streets [Bil65a, Bil65b, Cod51, Cod65]. Strength [Cha31]. Stress [Bue47, Che66, Coh66, DD61, Doy57, HS68, Hey24a, Hil43, Hil42, Hod50, Hov34, KR58, Lar66, Ōku52, Pay64, Phi34b, Rei44a, Rei52b, Shi54, Sti65, Tho44, Tho56, Tri56]. Stress-Strain [KR58, Tho56]. Stresses [Fol65, Hey24b, Hov25, Lin47b, RT46, Rei46a, Rei46b, Rei48a, Rei59]. Strip [Bak68, Han62]. Strongly [Pap58]. Structure [Dou36a, Val25]. Structures [Hov34, Tin53]. Struve [Abr50b, Hor50a, Hor50b, MP46]. Study [Fri35, Hym51, Kei62b, Mah27, OHK50, Rei55, Sco66, Wee34a, Wie27, Spe44b]. Studying [Ben61, Cre35]. Sub [Fra28b]. Sub- [Fra28b]. Subdeterminants [Ced57]. Subdivision [Kag38]. Subjected [Pay64, Wei62]. Sublimation [Cro36, GC36]. Subsonic [Col51, Dia59, Fis63, Hel61, Lin49, Lud51, Ser54, TR67, Van58]. Subspace [Moo29c]. Successive [Lin41, Ste55]. Suction [Tor52, Tor53]. Sufficient [Rut30]. Sum [GR55, Hit27, Ric22b, Rut27]. Summability [Bil61, Lit29]. Summation [BM66, Dou31c, Nav61, Nav62, Rut32, RD35, Spi57, Szá49, Whe55]. Summing [Sal54, Wyn56]. Sums [Cam40, Fej34, Luk52b, MS57]. Super [Duf48, Fra28b]. Super-Biharmonic [Duf48]. Super-Harmonic [Fra28b]. Supersonic [CW52, Elr58, Fin55, MG54, Rob48, Sto48, Sto51, Sto52]. Sure [Koz63]. Surface [BG60, Car42a, Fox51, Kag41, Mar67, Moo22b, Phi34a, Tor52, Tor53]. Surfaces [Dou36a, Eri55b, Fra34a, Fra35a, Kag36, Kag37, Kag38, Moo24, Moo27, Sai28, Tho44, Wil42]. Survey [LC34]. Sweptback [Tur49]. Switching [Gil54]. Symmetric [Ber50, Cha61, O’M63, Ran65, Sch55b, Wei62, Wil67]. Symmetrical [HKP36, Mar57, Sim56, TH50]. Symmetrically [Jen60]. Symmetry [Bla40a, Bla40c, Rei44a]. Synchrotron [Par54]. Synthesis [Bru31, Dar39, FG55, Gew33, Gui49, Lee32, Oon50, Tel53]. System [EH66, FLL50, Gar35, Gar37, Gui60, HD64, Kar43, KP53, Lou61, Moo30, PS56, Ric32, Was59]. Systematic [KP67]. Systems [CD62, Cod51, Cod65, Fra29b, Fra30, FM30, Gar37, HR67, Hod52, Koz63, KB63, Lip23c, May39, Muc29, SE59, SK52, Tho46b, Val24b, Val24a, Wad32, War52, War53, You53].

Table [Abr51, Cla37, Cor45, Dwi46, Dwi48, Hor50b, HG68, Kap46, Kap48, Kay55, Kop48, LH43, LS43, LL43, LBA43, LS45, Luk58, MP46, Moo37, Moo40, Pro44, Rus33b, Rut29a, Sal44a, Sal44b, Sal45c, Sal53b, Sal43b]. Tables [Abr50b, Ano43, BHL63, Bea27, Bla40a, Bla40b, CT50, CM45, Heu41a, Heu41b, HB54, Lin52, LA43, RC30, Sal47c, Sal47b, Sal58, Sal61, Sal63b, Sch55a, Ste54a]. Tabular [Hig42]. Tabulated [Sal45a, Sal47a]. Take [Moo29b]. Taken [Ham52]. Tangent [Gol58]. Tauberian [Fri35, Ike31d, Lit29, Wie28b]. Tautochronous [Zel24]. Taylor [Ric25c]. Tchebysheff [SB42]. Technique [KP67, Mor57]. Telegraph [CW52]. Telegraphist [Elr58]. Temperature [Aba60, CT67, Jae55, Kel67, KTY22]. Ten [Sal45b]. Ten- [Sal45b]. Tension [Lin47b, Sch66, Smi57]. Tensor [Hit27, Hit28, Rei44a, Spe44b]. Tensors [Doy57, Hof46]. Term [Kog35, Lev43b, Sch42]. Terme [Kog35]. Terminal [Gew33, Har59b, Oon50, Pay53, RS42, Bru31]. Terms [Abr50a, Abr54, Kar49, ME63, ME64, Sal44a, WW22]. Ternary [Zel23b]. Testimony [Tay24]. Tests [Hov29, MA65]. Tetrahedra [Spe42a]. Their [Coc67, Her64, Kli48, McL46, McL47a, Zel23a, Cro60, Pet65]. Theorem [ASC63, Ber29, Cha34, Cou41, Dav57, DeC56, Dou31a, Eps38, Fra25b, Fra26c, Hos36a, Ike31a, KS52, Kra59, Lan51, Lev37, Lev40, Lit29, Mir66, Pin51, Pit37, Rei44a, Rei50, Rei53, Ric24, SS67, SS68b, Sha49, Spe41a, Wal59, WP38, AA65, Hei35]. Theorems [Bar64b, BM62a, De 47, Dun67, Fel58, Fri35, Ike31d, May39, Ser54, Wan68a, Wie28b]. Theoretic [Gil54, Har59a]. Theory [AM27b, Ban39, Ber41, BG47, Bil65a, Bil65b, BR60, Bod34, Bol57a, Bur35, Bur66c, Car42b, Cla50, Cra59, Cro36, De 26, Doy55, Ell67, Eps42, Fle58, Fox55, Fra26b, Gei47, GC36, GS50, Hil42, Hit25a, Hit25b, Hit29, Hov23, Hov32, Hsu42a, Hsu42b, Ike31a, IC44, Isa47, Jae48, Jan62, JR58, Joh63, Kei62b, Kei63, Kli48, Kno26, KR56, Krö63, Kue64, LC54, Law34, LC34, Lev47, Lif42b, Lin54, ML59a, Mil49, Mir66, Nag67, Obe55, Ort30, Pap54, Pay53, Phi22b, Phi33, Pin46, Pin51, Red49b, Red62, Rei37, Rei44b, Rei47, Rei52b, Rei63b, RS66b, Sat66, Sav38, Sav41a, Sav41b, Sch31a, Sch31b, Sea49, Sha41a, Shi57, Soh33, SK52, Ste57, Str28b, Str35, Su59, Tay24, TM40, Tri56, Tur49, Val25, Val26, Wan68c, Wie24a, Wie29a]. Theory [Wu54, tW56a, Wu66, YL56, Zel22b, Bar61]. Thermal [Pet57]. Thermodynamic [Bea30, KTS22, Val68]. Thermodynamics [Val66]. Thermometer [Key22]. these [Kog35]. Thickness [Con64, HS68, MG54, OR40, Rei37, Wan68d]. Thin [Buc67, Buc68, Cla50, Fin55, JR58, KR58, Low45, Mar57, Mei66, Rei52b, Rei63b, SS68a, SS68b, Wan68a]. Thin-Symmetrical-Sections [Mar57]. Third [Che61, Wad32]. Thomson [Bea30]. Those [Bar61, Kog35]. Three [BBBP64, BR60, Bol57b, Car42b, Cer61, ES67, Fet57a, Fra27b, Gie51, Gol46, Kap36, Lou65, Moo25, Ric25a, Rob36, Ros64, Sav37, Sav41a, Sti65, Wad32, Wee34b]. Three-Dimensional [BR60, Bol57b, ES67, Gie51, Gol46, Kap36, Rob36, Sti65]. Time [Aba60, BP32, BN67b, Joh30, Wil67, Zad51]. Time-Dependent [Zad51]. Time-Symmetric [Wil67]. Timoshenko [ML59a]. Topological [Dou36a, Doy55, Sai28]. Topology [Gre40]. Toroidal [Cla50, Jen60, Rei63a, Tao59, Wes60]. Torsion [Bas59, GS58, HH25, Phi48, Rei52a, RW68, Set52]. Torsional [Mar57, Wei53]. Trajectories [Mac41]. Trajectory [Mac41]. Transcendental [LH43, Luk55, Luk59, McM39]. Transfer [BW61, Dei68, FG51, Gew33, Gro62, KP53, Nay67, Red62]. Transform [Ber29, Fel58, Hei35, Kin65, Lev35, Mar63b, Mor65b]. Transformation [Blo53, Gol62, Gol67, Gui60, KD48, Red60, Ros30, Tho46a]. Transformations [Bol57b, De 47, Hit24, Law34, Lip23b, Rob36, Sai28, Sha55, Spe41b, Sta36a, Zel22a, Zel23a, Zel27, Zem61]. Transformer [Rez54]. Transforms [Cam56, Chu54, HS36, Kan64, Lee32, Rad59, Sal58, SS61, Smi68]. Transients [Cre35, Jae48]. Transition [Ros66b]. Translation [Mah27, Wie27]. Transmission [BP32, Hsu42a, Hsu42b, Red62, SK52]. Transmission-Line [Red62]. Transonic [RW56, TH50, TW53, vk47]. Transport [Sat66]. Transportation [FW64, Gil60, Kuh57, Pra57]. Transversality [Dut44]. Transverse [BR60, HE64, Hec67, Pap49, Ran67]. Treated [Mar56]. Treating [Nay65a]. Trefftz [DD47]. Triangle [Fra27b]. Triangles [Fra27b]. Triangular [Eav53, Rob48]. Triangulation [Kag36]. Trigonometric [Cam40, Erd43, Lan38, Sal48b, Sal60b, SR52, Str62a]. Trigonometrical [Pit37, Wie25c]. True [Pea37]. Truncation [Fri56, Gie58]. Tube [Abr53b, Cha61, CC68, Hsu42a, Nay67]. Tubes [Bla40b]. Tubular [Gol47]. tuply [Kap34]. Turbulence [Dei68]. Turbulent [Ste68, Yua59]. Turning [HR67, Nay67, Was59]. Twin [Sut37]. Twisted [BD60, Was64]. Two [AP68, Bar68, Ber41, Bru31, Bus67, Car42b, Cer61, CW52, Dei68, Dia59, Dou31b, Dou32, DL30, Eni61, Fra26b, Fra26c, Gea64, Gri38, Har59b, HF47, Hei51, Hit24, Jon56, KM62, LRT48, Lud51, Mag42, Moo25, Nay67, Pau61, Por32, Por60, Rai53, Ran66, Ran67, Ric22b, RS42, Ros58, Sch42, She52, SW66, SW68, Spo55, SK53, Ste54a, Ste54b, Tin59, TR67, Wan68a, Was59, Wil66, tW56a, Zyg42, Pet65, Mah27]. Two-Dimensional [AP68, Bar68, Bus67, Dei68, Dia59, Eni61, Gri38, LRT48, Pau61, She52, SW66, SW68, tW56a]. Two-Frequency [SK53, Ste54a, Ste54b]. Two-term [Sch42]. Two-Terminal [Har59b, RS42, Bru31]. Type [BH64a, Don67, Dwi48, Fet57a, Fla54, Gon41, Lat63, Lev41, Pea68a, Pea68b, Red64, Ric22a, Sal42, Sal64, Wie22].

unbounded [Pet65]. Undergoing [LF67]. Undetermined [Mir63]. Unendlicher [EGV38]. Uniform [BD60, Fin55, Rei52a, SK52, Str68]. Uniformly [Seg60]. Uniformly-Valid [Seg60]. Uniqueness [Dun67, Eri52, Mar63a, Mir66, Pri49]. Unitary [Cob42]. United [Hov26].

Universe [Lem25]. Unmechanical [Val24b]. Unstable [Leu51]. Unsteady [CD56, Eni61, Seg60, tW56b]. Unsymmetric [Joh63]. Unsymmetrical [Rei59, SS68a]. Unusual [Dou59]. Uplift [Wea32]. Upper [Cro59, DG48, DR52, Gre48, Wei53]. Use [BG47, Bil61, Cre35, Isa47, Kei62b, Roe48, Sal46a, Sal47b, Pin51]. Using [AK64, Pau59, Sch66].

Vacuum [Rey67, Sco66]. valent [Spe42b]. Valid [Seg60]. Validity [Val27]. Vallée [RD35, Dou31c, Rut32]. Vallée-Poussin [Dou31c]. Value [BP63, BH64b, Che61, Cob44a, Cob44b, Coh64, Coh65, DG48, DR52, Fis35, For54, For56, GR55, Nag65, Ran64, RS66a, Ric32, Roe48, Tho40, Was53, Was44]. Valued [KR60]. Values [Cam38, Car58, Cow65, Fet57b, For56, Jae55, Lin52, Rut29b]. Vanes [BAC24]. Vanishing [Moo23, Lev37]. Vapor [ZP54]. Variable [Ad63, Con64, HD64, OR40, Rei37, Wat63]. Variables [Ber41, BM42, Ber42, Fra26b, Fra26c, Mag42, Sal47c, Sav37, Val24b]. variants [Pet65]. Variate [Mor57]. Variation [Kar43, Wie24d]. Variational [DS58, HP48, KS52, Pom68, Rei50, Rei53, SS67, Sle60, Wan68a]. Varieties [Moo25]. Various [EBK55, EBK57]. Varying [Mar67]. Vector [Gou58, HW22, Hos36b, Hos36a, KR60, Pur53, RS56]. Vector-Valued [KR60]. Vectors [Cre35, LH26]. Velocities [Fra30]. Velocity [BB23, Fox51, Gar35, Gar37, Nag51, Pay64, Shi54, Syn36]. Velocity-Distributions [Syn36]. Venant [Sav41c]. Vertical [Lew63, LF67, OB66]. Vibrating [Ava35, Duf52, For54, Muc29]. Vibration [Mar56, Mun52, Zwi58c]. Vibrations [BD60, Eps42, Ros66b]. View [Har59a]. Viscoelastic [Val65, Val66]. Viscosity [Phi22a, fS52]. Viscous [Abr49b, Ber67, CD56, GP62, Gri38, Han57b, How61, Kuo43, Kuo53, Mei66, Mow67, PK56, Ran66, Ran67, Ran68, TY56, tW56b]. Visoelastic [Val68]. Visual [MS45]. Visualization [Ber41]. Volterra [Wag53]. Volume [Key22]. Volumes [Bla40b]. Vortex [Bil65a, Bil65b, Cod51, Cod65, Hun61]. Vortices [Por42].

W.K.B. [Ima49, IH49]. Waal [Tsi47, sT46]. Wake [Ber68]. Wakes [Hun61]. Walks [Kei62a, Kei62b]. Wall [HW67, Kel67]. Walls [Dia59, Kue51, KF56, Yua59]. Warm [Kue64, Kue66]. Water [Bar61]. Wave [Abr48, Abr49a, Abr50a, Abr54, AA65, AM27b, AH64, BN67a, Bla47, Bod34, Bou47, Bur66a, Bur66c, Che64, CS41, Coo57, Cre35, Fox55, Han57a, Han57b, Hof34, LO51, Lyn66, Mar67, Mar53, Pap54, Pau59, Pau61, Pin51, Ran50, Rob57b, Ros66a, Sle65, Str68, sT47, Wes68, Wes60]. Wave-Matrix [Coo57]. Waves [Ang59, Bar61, BM62a, BN62, BL64, Ben66a, Ben66b, Ben67, BN67b, Bro50, Bur66b, CT67, FR57, Gar57, Kan62, Kue64, Lew63, Lin57, Mei66, Obe55, Rob57a, Syn56, Tho47, Tho48, Tho49a, Tho49c, Val65, Wil57]. Way [Hit28]. Weak [Dei68, Tin53]. Weakly [BN62]. Wedge [Col51, Fin55, Hel61, KP67, PK56, PW60, fS52, TW53]. Wedge-Like [PW60]. Wedges [Dia59, Ell67, Obe55]. Weierstrass [Fra25b]. Weight [GD49, Ral56]. Weighted [Sal55a]. Welds [Hov34]. Well [BV65]. Well-Set [BV65]. Which [BBBP64, Dar39, Moo29b, Oon50, Pea37, BH64a, Rey67, Wan67]. Whittaker [Tay39]. Whose [Ced57, Moo25, Bru31]. Wideband [LC54]. Width [GZ43]. Wiener [CM44, Cam60, Lev46, Lev47, PB68, Wil66]. Wing [BD41, Rob48, Wan68c]. Wings [Bar68, Tur49]. Wintner [Ike31d]. Wire [Mur33]. Within [Eic53, LS44]. Without [LS45, Rez54, RC30, Sal45c, FG55]. Witness [Tay24]. WKB [MG64a, MG64b]. Wood [BAC24]. Work [Cob38, Gew33, Hov25].

x [Ada43, Deb25, Val26]. X-Ray [Val26]. X-rays [Deb25].

Yawing [Sto48, Sto51, Sto52]. Yield [Cob44b, Hod50, SS68a]. Yielding [Bak68].

Zero [Bla47, Bou47, DD61]. Zeros [Coh64, Cro60, Gol57, Ike31c, Kir57, Lev40, LH43, Sco66, SCKH52, SCKH53, Str65, Coh65]. Zeta [Lev40]. Zonal [Cha34].

References

Aggarwal:1965:EBP

[AA65] H. R. Aggarwal and C. M. Ablow. Extension of the Bateman–Pekeris Theorem with application to a wave propagation problem. Journal of Mathematics and Physics (MIT), 44(1–4):267–270, April 1965. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

       wiley.com/doi/epdf/10.1002/sapm1965441267.

Abarbanel:1960:TDT

[Aba60] Saul S. Abarbanel. Time dependent temperature distribution in radiating solids. Journal of Mathematics and Physics (MIT), 39(1–4): 246–257, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

       //onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391246.

Abbott:1949:EIB

[Abb49] Wilton R. Abbott. Evaluation of an integral of a Bessel function. Journal of Mathematics and Physics (MIT), 28(1–4):192–194, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

       wiley.com/doi/epdf/10.1002/sapm1949281192.

Abramowitz:1948:BFE

[Abr48] Milton Abramowitz. Bessel function expansions of certain Coulomb wave functions. Journal of Mathematics and Physics (MIT), 27(1–4): 157–158, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271157.

Abramowitz:1949:AES

[Abr49a] Milton Abramowitz. Asymptotic expansions of spheroidal wave functions. Journal of Mathematics and Physics (MIT), 28(1–4):195–199, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281195.

Abramowitz:1949:BVF

[Abr49b] Milton Abramowitz. On the backflow of a viscous fluid in a diverging channel. Journal of Mathematics and Physics (MIT), 28(1–4):1– 21, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19492811.

Abramowitz:1950:CWF

[Abr50a] Milton Abramowitz. Coulomb wave functions expressed in terms of Bessel–Clifford and Bessel functions. Journal of Mathematics and Physics (MIT), 29(1–4):303–308, April 1950. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1950291303.

Abramowitz:1950:TIS

[Abr50b] Milton Abramowitz. Tables of integrals of Struve functions. Journal of Mathematics and Physics (MIT), 29(1–4):49–51, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195029149.

Abramowitz:1951:TI

[Abr51] Milton Abramowitz. Table of the integral. Journal of Mathematics and Physics (MIT), 30(1–4):162–163, April 1951. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1951301162.

Abramowitz:1953:EI

[Abr53a] Milton Abramowitz. Evaluation of the integral. Journal of Mathematics and Physics (MIT), 32(1–4):188–192, April 1953. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1953321188.

Abramowitz:1953:SDE

[Abr53b] Milton Abramowitz. On the solution of the differential equation occurring in the problem of heat convection in laminar flow through a tube. Journal of Mathematics and Physics (MIT), 32(1–4):184–187, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321184.

Abramowitz:1954:RIC

[Abr54] Milton Abramowitz. Regular and irregular Coulomb wave functions expressed in terms of Bessel–Clifford functions. Journal of Mathematics and Physics (MIT), 33(1–4):111–116, April 1954. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1954331111.

Ames:1963:SCL

[Ad63] W. F. Ames and H. de la Cuesta S. Solutions of a class of linear partial differential equations with variable coefficients. Journal of Mathematics and Physics (MIT), 42(1–4):301–306, April 1963. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1963421301.

Adams:1943:QHE

[Ada43] Douglas P. Adams. The quintic “Hypernom” for the equation x^{5} + Ax^{3} + Bx^{2} + Cx + D = 0; a graphical method for finding the roots of polynomial equations through the fifth degree. Journal of Mathematics and Physics (MIT), 22(1–4):78–92, April 1943. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm194322178.

Auld:1964:MCS

[AH64] B. A. Auld and W. B. Hatfield. Method of characteristics solution for electromagnetic wave propagation in a gyromagnetic medium. Journal of Mathematics and Physics (MIT), 43(1–4):34–37, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196443134.

Ahmed:1964:NEN

[AK64] Nasir Ahmed and S. Karni. A note on equivalent networks using matrix partitioning. Journal of Mathematics and Physics (MIT), 43(1–4):336– 338, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431336.

Alden:1948:SAL

[Ald48] Henry Leonard Alden. Second approximation to the laminar boundary layer flow over a flat plate. Journal of Mathematics and Physics (MIT), 27(1–4):91–104, April 1948. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194827191.

Allis:1927:LF

[AM27a] William Phelps Allis and Hans Müller. Lichtenberg’s figures. Journal of Mathematics and Physics (MIT), 6(1–4):62–68, April 1927. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19276162.

Allis:1927:WTE

[AM27b] William Phelps Allis and Hans Müller. A wave theory of the electron. Journal of Mathematics and Physics (MIT), 6(1–4):119–132, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761119.

Anderson:1966:E

[AM66] Donald G. Anderson and Hilliard K. Macomber. Evaluation of. Journal of Mathematics and Physics (MIT), 45(1–4):109–120, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451109.

Amos:1960:SCP

[Amo60] Nathan Amos. Simplicial coordinates and piecewise-linear interpolation in a regular n-dimension. Journal of Mathematics and Physics (MIT), 39(1–4):198–210, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391198.

Ang:1959:EWG

[Ang59] Dang Dinh Ang. Elastic waves generated by a force moving along a crack. Journal of Mathematics and Physics (MIT), 38(1–4):246– 256, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381246.

Anonymous:1943:APN

[Ano43] Anonymous. Announcement of the publication of new and more extensive tables of Lagrangean interpolation coefficients. Journal of Mathematics and Physics (MIT), 22(1–4):1, April 1943. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19432211.

Ackerberg:1968:ITD

[AP68] Robert C. Ackerberg and Alexander Pal. On the interaction of a two-dimensional jet with a parallel flow. Journal of Mathematics and Physics (MIT), 47(1–4):32–56, April 1968. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196847132.

Archer:1962:NSN

[Arc62] Robert R. Archer. On the numerical solution of the nonlinear equations for shells of revolution. Journal of Mathematics and Physics (MIT), 41(1–4):165–178, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411165.

Arsove:1953:NNP

[Ars53] Maynard G. Arsove. A note on the network postulates. Journal of Mathematics and Physics (MIT), 32(1–4):203–206, April 1953. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1953321203.

Al-Salam:1963:GAT

[ASC63] W. A. Al-Salam and L. Carlitz. The Gegenbauer addition theorem. Journal of Mathematics and Physics (MIT), 42(1–4):147–156, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1963421147.

Atoji:1959:CSH

[Ato59] Masao Atoji. On the cubic spherical harmonics. Journal of Mathematics and Physics (MIT), 38(1–4):73–74, April 1959. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195938173.

Avakian:1935:APF

[Ava35] Arra Steve Avakian. Almost periodic functions and the vibrating membrane. Journal of Mathematics and Physics (MIT), 14(1–4):350–378, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141350.

Banos:1939:AOT

[Ban39] Alfredo Baños, Jr. On asymptotic orbits in the theory of primary cosmic radiation. Journal of Mathematics and Physics (MIT), 18(1–4): 211–238, April 1939. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1939181211.

Budiansky:1968:IPB

[BA68] Bernard Budiansky and John C. Amazigo. Initial post-buckling behavior of cylindrical shells under external pressure. Journal of Mathematics and Physics (MIT), 47(1–4):223–235, April 1968. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1968471223.

Barss:1924:EWV

[BAC24] W. R. Barss, J. A. Abbott, and H. M. Chatto. Effects of wood vanes in phonograph horns. Journal of Mathematics and Physics (MIT), 3(1): 52–62, January 1924. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19243152.

Bagchi:1953:NEM

[Bag53] Hari Das Bagchi. Note on an equi-momental complex of a rigid body. Journal of Mathematics and Physics (MIT), 32(1–4):307–311, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321307.

Baker:1968:YBC

[Bak68] Bill R. Baker. Yielding and brittle cracking of an orthotropic strip. Journal of Mathematics and Physics (MIT), 47(1–4):236–248, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471236.

Barakat:1961:CIR

[Bar61] Richard Barakat. On a class of integrals related to those of Raabe and Laplace occurring in the theory of water waves. Journal of Mathematics and Physics (MIT), 40(1–4):244–248, April 1961. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1961401244.

Barakat:1964:MOL

[Bar64a] Richard Barakat. The matrix operator e^{X} and the Lucas polynomials. Journal of Mathematics and Physics (MIT), 43(1–4):332–335, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431332.

Barakat:1964:SAI

[Bar64b] Richard Barakat. Solution of an Abel integral equation for bandlimited functions by means of sampling theorems. Journal of Mathematics and Physics (MIT), 43(1–4):325–331, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1964431325.

Barakat:1968:IFA

[Bar68] Richard Barakat. Incompressible flow around porous two-dimensional sails and wings. Journal of Mathematics and Physics (MIT), 47(1–4): 327–349, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471327.

Bassali:1959:TEC

[Bas59] W. A. Bassali. The torsion of elastic cylinders with regular curvilinear cross sections. Journal of Mathematics and Physics (MIT), 38(1–4): 232–245, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381232.

Barss:1923:EHV

[BB23] W. R. Barss and J. E. Bastille. The effect of humidity on the velocity of sound in air. Journal of Mathematics and Physics (MIT), 2(1–4):210– 215, October 1923. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192321210.

Ball:1964:SNG

[BBBP64] R. Ball, D. Bartlett, R. Bayer, and F. Partovi. Some new geometries for which Laplace’s equation in three dimensions is soluble. Journal of Mathematics and Physics (MIT), 43(1–4):144–157, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431144.

Burington:1941:NFW

[BD41] Richard S. Burington and James M. Dobbie. A new family of wing profiles. Journal of Mathematics and Physics (MIT), 20(1–4):388– 401, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201388.

Budiansky:1960:BVU

[BD60] Bernard Budiansky and R. C. Diprima. Bending vibrations of uniform twisted beams. Journal of Mathematics and Physics (MIT), 39(1–4): 237–245, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391237.

Bergman:1947:DCF

[BE47] Stefan Bergman and Bernard Epstein. Determination of a compressible fluid flow past an oval-shaped obstacle. Journal of Mathematics and Physics (MIT), 26(1–4):195–222, April 1947. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1947261195.

Beattie:1927:SPT

[Bea27] James A. Beattie. Six place tables of the Debye energy and specific heat functions. Journal of Mathematics and Physics (MIT), 6(1–4): 1–32, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1927611.

Beattie:1930:STP

[Bea30] James A. Beattie. Some of the thermodynamic properties of ammonia: II. the Joule-Thomson effect and the heat capacity at constant pressure. Journal of Mathematics and Physics (MIT), 9(1–4):11–21, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19309111.

Beekman:1967:FCI

[Bee67] J. A. Beekman. Feynman–Cameron integrals. Journal of Mathematics and Physics (MIT), 46(1–4):253–266, April 1967. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1967461253.

Bellman:1951:AEE

[Bel51] Richard Bellman. On approximate expressions for the exponential integral and the error function. Journal of Mathematics and Physics (MIT), 30(1–4):226–231, April 1951. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1951301226.

Benes:1961:FPM

[Ben61] V. E. Beneś. A fixed point method for studying the stability of a class of integrodifferential equations. Journal of Mathematics and Physics (MIT), 40(1–4):55–67, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196140155.

Benes:1965:NIE

[Ben65] V. E. Benes. A nonlinear integral equation in the Marcinkiewicz space. Journal of Mathematics and Physics (MIT), 44(1–4):24–35, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196544124.

Benney:1966:LNL

[Ben66a] D. J. Benney. Long non-linear waves in fluid flows. Journal of Mathematics and Physics (MIT), 45(1–4):52–63, April 1966. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196645152.

Benney:1966:LWL

[Ben66b] D. J. Benney. Long waves on liquid films. Journal of Mathematics and Physics (MIT), 45(1–4):150–155, April 1966. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1966451150.

Benney:1967:ABN

[Ben67] D. J. Benney. The asymptotic behavior of nonlinear dispersive waves. Journal of Mathematics and Physics (MIT), 46(1–4):115–132, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461115.

Berry:1929:FTT

[Ber29] Andrew C. Berry. The Fourier transform theorem. Journal of Mathematics and Physics (MIT), 8(1–4):106–118, April 1929. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192981106.

Bernstein:1935:PPN

[Ber35] Felix Bernstein. Principles of probability in natural science. Journal of Mathematics and Physics (MIT), 14(1–4):28–35, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193514128.

Bergman:1941:VDT

[Ber41] Stefan Bergman. The visualization of domains of the theory of functions of two complex variables. Journal of Mathematics and Physics (MIT), 20(1–4):107–117, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1941201107.

Bergman:1942:RPF

[Ber42] Stefan Bergman. A remark on the paper “Sur Les Fonctions Analytiques de deux Variables Complexes”. Journal of Mathematics and Physics (MIT), 21(1–4):141–143, April 1942. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1942211141. See [BM42].

Bergman:1950:DAS

[Ber50] S. Bergman. Determination of axially symmetric flow patterns of a compressible fluid. Journal of Mathematics and Physics (MIT), 29(1–4): 133–145, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291133.

Berg:1967:CPD

[Ber67] C. A. Berg. The character of plane deformation fields in isotropic incompressible nonlinear viscous fluids. Journal of Mathematics and Physics (MIT), 46(1–4):448–458, April 1967. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1967461448.

Berger:1968:ILA

[Ber68] Stanley A. Berger. The incompressible laminar axisymmetric far wake. Journal of Mathematics and Physics (MIT), 47(1–4):292–309, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471292.

Bergman:1947:NDU

[BG47] Stefan Bergman and Leonard Greenstone. Numerical determination by use of special computational devices of an integral operator in the theory of compressible fluids: I. Determination of the coefficients of the integral operator by the use of punch card machines. Journal of Mathematics and Physics (MIT), 26(1–4):1–9, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19472611.

Bordoni:1948:SRF

[BG48] Piero Giorgio Bordoni and Wolf Gross. Sound radiation from a finite cylinder. Journal of Mathematics and Physics (MIT), 27(1–4):241– 252, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271241.

Birkhoff:1960:SSI

[BG60] Garrett Birkhoff and Henry L. Garabedian. Smooth surface interpolation. Journal of Mathematics and Physics (MIT), 39(1–4):258–268, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391258.

Barakat:1963:CES

[BH63a] Richard Barakat and Agnes Houston. Chebyschev expansion of the sine and cosine integrals. Journal of Mathematics and Physics (MIT), 42(1–4):331–333, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421331.

Birkhoff:1963:FBP

[BH63b] Garrett Birkhoff and Donald F. Hays. Free boundaries in partial lubrication. Journal of Mathematics and Physics (MIT), 42(1–4):126– 138, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1963421126.

Bramble:1964:FDA

[BH64a] J. H. Bramble and B. E. Hubbard. On a finite difference analogue of an elliptic boundary problem which is neither diagonally dominant nor of non-negative type. Journal of Mathematics and Physics (MIT), 43(1–4):117–132, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431117.

Bueckner:1964:ESB

[BH64b] H. Bueckner and G. Horvay. Exponential solutions of a boundary value problem of the Helmholtz equation related to change in phase. Journal of Mathematics and Physics (MIT), 43(1–4):227–233, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431227.

Barakat:1963:PSE

[BHL63] Richard Barakat, Agnes Houston, and Elgie Levin. Power series expansions of Mathieu functions with tables of numerical results. Journal of Mathematics and Physics (MIT), 42(1–4):200–247, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1963421200.

Banos:1965:RFE

[BHPBT65] A. Baños, Jr., R. H. Huddlestone, D. C. Pridmore-Brown, and E. C. Taylor. The radiation field of an electric dipole antenna in a conical sheath. Journal of Mathematics and Physics (MIT), 44(1–4): 189–213, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441189.

Bickley:1948:DAO

[Bic48] W. G. Bickley. Difference and associated operators, with some applications. Journal of Mathematics and Physics (MIT), 27(1–4):183–192, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271183.

Biedenharn:1952:ISR

[Bie52] L. C. Biedenharn. An identity satisfied by the Racah coefficients. Journal of Mathematics and Physics (MIT), 31(1–4):287–293, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1952311287.

Bilodeau:1961:USM

[Bil61] G. G. Bilodeau. On the use of summability methods to increase the rate of convergence of series. Journal of Mathematics and Physics (MIT), 40(1–4):289–299, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1961401289.

Bils:1965:CSTa

[Bil65a] O. Bils. On Coddington’s stability theory of vortex streets. Journal of Mathematics and Physics (MIT), 44(1–4):86–88, April 1965. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196544186.

Bils:1965:CSTb

[Bil65b] O. Bils. On Coddington’s stability theory of vortex streets. Journal of Mathematics and Physics (MIT), 44(1–4):95–98, April 1965. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196544195.

Birkhoff:1937:PI

[Bir37] Garrett Birkhoff. On product integration. Journal of Mathematics and Physics (MIT), 16(1–4):104–132, April 1937. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1937161104.

Benney:1964:IPW

[BL64] D. J. Benney and J. C. Luke. On the interactions of permanent waves of finite amplitude. Journal of Mathematics and Physics (MIT), 43(1–4): 309–313, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431309.

Blaisdell:1940:PPFa

[Bla40a] B. Edwin Blaisdell. The physical properties of fluid interfaces of large radius of curvature. I. Integration of LaPlace’s equation for the equilibrium meridian of a fluid drop of axial symmetry in a gravitational field. Numerical integration and tables for sessile drops of moderately large size. Journal of Mathematics and Physics (MIT), 19(1–4):186– 216, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1940191186.

Blaisdell:1940:PPFb

[Bla40b] B. Edwin Blaisdell. The physical properties of fluid interfaces of large radius of curvature. II. Numerical tables for capillary depressions and meniscus volumes in moderately large tubes. Journal of Mathematics and Physics (MIT), 19(1–4):217–227, April 1940. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1940191217.

Blaisdell:1940:PPI

[Bla40c] B. Edwin Blaisdell. The physical properties of interfaces of large radius of curvature. III. Integration of LaPlace’s equation for the equilibrium Meridian of a fluid drop of axial symmetry in a gravitational field. approximate analytic integration for sessile drops of large size. Journal of Mathematics and Physics (MIT), 19(1–4):228–245, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1940191228.

Blanch:1946:CMF

[Bla46] Gertrude Blanch. On the computation of Mathieu functions. Journal of Mathematics and Physics (MIT), 25(1–4):1–20, April 1946. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19462511.

Blanch:1947:NCJ

[Bla47] Gertrude Blanch. Note on C. J. Bouwkamp’s paper “On Spheroidal Wave Functions of Order Zero”. Journal of Mathematics and Physics (MIT), 26(1–4):93, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194726193. See [Bou47].

Block:1952:LAH

[Blo52] H. D. Block. Laws of attraction having a certain generalized Newtonian property. Journal of Mathematics and Physics (MIT), 31(1–4):151– 153, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311151.

Block:1953:NCT

[Blo53] H. D. Block. A note on contact transformation. Journal of Mathematics and Physics (MIT), 32(1–4):207–208, April 1953. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1953321207.

Bergman:1942:FAD

[BM42] S. Bergman and J. Marcinkiewicz. Sur les fonctions analytiques de deux variables complexes. Journal of Mathematics and Physics (MIT), 21(1–4):125–140, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1942211125. See remark [Ber42].

Ben-Menahem:1962:ORA

[BM62a] Aru Ben-Menahem. An operational representation of the addition theorems for spherical waves. Journal of Mathematics and Physics (MIT), 41(1–4):201–204, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411201.

Butler:1962:MCM

[BM62b] Terence Butler and Abram V. Martin. On a method of Courant for minimizing functionals. Journal of Mathematics and Physics (MIT), 41(1–4):291–299, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411291.

Ben-Menahem:1966:SCL

[BM66] Ari Ben-Menahem. Summation of certain Legendre series and related difference equations. Journal of Mathematics and Physics (MIT), 45(1–4):224–228, April 1966. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1966451224.

Ben-Menahem:1963:DRE

[BMC63] Ari Ben-Menahem and Armando Cisternas. The dynamic response of an elastic half-space to an explosion in a spherical cavity. Journal of Mathematics and Physics (MIT), 42(1–4):112–125, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1963421112.

Benney:1962:ARW

[BN62] D. J. Benney and A. M. Niell. Apparent resonances of weakly nonlinear standing waves. Journal of Mathematics and Physics (MIT), 41(1–4): 254–263, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411254.

Benney:1967:PNW

[BN67a] D. J. Benney and A. C. Newell. The propagation of nonlinear wave envelopes. Journal of Mathematics and Physics (MIT), 46(1–4):133– 139, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461133.

Benney:1967:STC

[BN67b] D. J. Benney and A. C. Newell. Sequential time closures for interacting random waves. Journal of Mathematics and Physics (MIT), 46(1–4): 363–393, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461363.

Bode:1934:GTE

[Bod34] H. W. Bode. A general theory of electric wave filters. Journal of Mathematics and Physics (MIT), 13(1–4):275–362, April 1934. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1934131275.

Bogert:1951:SRE

[Bog51] B. P. Bogert. Some roots of an equation involving Bessel functions. Journal of Mathematics and Physics (MIT), 30(1–4):102–105, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1951301102.

Boley:1957:MCF

[Bol57a] Bruno A. Boley. A method for the construction of fundamental solutions in elasticity theory. Journal of Mathematics and Physics (MIT), 36(1–4):261–268, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361261.

Bolinder:1957:ITE

[Bol57b] E. Folke Bolinder. Impedance transformations by extension of the isometric circle method to the three-dimensional hyperbolic space. Journal of Mathematics and Physics (MIT), 36(1–4):49–61, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195736149.

Boley:1961:MHC

[Bol61] Bruno A. Boley. A method of heat conduction analysis of melting and solidification problems. Journal of Mathematics and Physics (MIT), 40(1–4):300–313, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1961401300.

Bonham:1961:SAF

[Bon61] Russell A. Bonham. Some asymptotic formulas for the Legendre functions P_{ν}^{μ}(cos θ) and Bessel functions J_{ν}(x) for small θ and x. Journal of Mathematics and Physics (MIT), 40(1–4):282–288, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1961401282.

Bonham:1966:SPI

[Bon66] R. A. Bonham. On some properties of the integrals and. Journal of Mathematics and Physics (MIT), 45(1–4):331–334, April 1966. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451331.

Bouwkamp:1947:SWF

[Bou47] C. J. Bouwkamp. On spheroidal wave functions of order zero. Journal of Mathematics and Physics (MIT), 26(1–4):79–92, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194726179. See note [Bla47].

Bowie:1956:AIP

[Bow56] O. L. Bowie. Analysis of an infinite plate containing radial cracks originating at the boundary of an internal circular hole. Journal of Mathematics and Physics (MIT), 35(1–4):60–71, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195635160.

Bowie:1966:AEN

[Bow66] O. L. Bowie. Analysis of edge notches in a semi-infinite region. Journal of Mathematics and Physics (MIT), 45(1–4):356–366, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451356.

Boyer:1961:SSN

[Boy61] R. H. Boyer. On some solutions of a non-linear diffusion equation. Journal of Mathematics and Physics (MIT), 40(1–4):41–45, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196140141.

Barnes:1932:TAT

[BP32] John L. Barnes and Clarence T. Prendergast. On the time admittances of transmission networks. Journal of Mathematics and Physics (MIT), 11(1–4):27–72, April 1932. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193211127.

Bramble:1961:ASH

[BP61] J. H. Bramble and L. E. Payne. An analogue of the spherical harmonics for the equations of elasticity. Journal of Mathematics and Physics (MIT), 40(1–4):163–171, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1961401163.

Bramble:1963:PBF

[BP63] J. H. Bramble and L. E. Payne. Pointwise bounds in the first biharmonic boundary value problem. Journal of Mathematics and Physics (MIT), 42(1–4):278–286, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421278.

Birkhoff:1967:HIE

[BP67] Garrett Birkhoff and Arthur Priver. Hermite interpolation errors for derivatives. Journal of Mathematics and Physics (MIT), 46(1–4):440– 447, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461440.

Blanch:1943:SPL

[BR43] G. Blanch and I. Rhodes. Seven-point Lagrangian integration formulas. Journal of Mathematics and Physics (MIT), 22(1–4):204–207, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1943221204.

Boal:1960:TDT

[BR60] J. L. Boal and E. Reissner. Three-dimensional theory of elastic plates with transverse inextensibility. Journal of Mathematics and Physics (MIT), 39(1–4):161–181, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391161.

Brachman:1954:NIR

[Bra54] Malcolm K. Brachman. Note on an integral of Ramanujan. Journal of Mathematics and Physics (MIT), 33(1–4):374–375, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331374.

Brown:1950:GCR

[Bro50] W. F. Brown. The general consistency relations for shock waves. Journal of Mathematics and Physics (MIT), 29(1–4):252–262, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291252.

Brune:1931:SFT

[Bru31] Otto Brune. Synthesis of a finite two-terminal network whose drivingpoint impedance is a prescribed function of frequency. Journal of Mathematics and Physics (MIT), 10(1–4):191–236, April 1931. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1931101191.

Buchwald:1967:SPT

[Buc67] V. T. Buchwald. Some problems of thin circular cylindrical shells. i. the equations. Journal of Mathematics and Physics (MIT), 46(1–4): 237–252, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461237.

Buchwald:1968:SPT

[Buc68] V. T. Buchwald. Some problems of thin circular cylindrical shells II. The infinite crack. Journal of Mathematics and Physics (MIT), 47(1– 4):57–66, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196847157.

Buell:1947:DPS

[Bue47] E. L. Buell. On the distribution of plane stress in a semi-infinite plate with partially stiffened edge. Journal of Mathematics and Physics (MIT), 26(1–4):223–233, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1947261223.

Burington:1935:MEC

[Bur35] Richard Stevens Burington. Matrices in electric circuit theory. Journal of Mathematics and Physics (MIT), 14(1–4):325–349, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141325.

Burington:1937:RME

[Bur37] Richard Stevens Burington. R-matrices and equivalent networks. I. Journal of Mathematics and Physics (MIT), 16(1–4):85–103, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193716185.

Burke:1964:ASR

[Bur64] J. E. Burke. Approximations for some radial Mathieu functions. Journal of Mathematics and Physics (MIT), 43(1–4):234–240, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431234.

Burke:1966:NSW

[Bur66a] J. E. Burke. Note on spheroidal wave functions. Journal of Mathematics and Physics (MIT), 45(1–4):425–431, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451425.

Burman:1966:RWD

[Bur66b] R. Burman. The reflection of waves at a discontinuity in refractive index gradient. Journal of Mathematics and Physics (MIT), 45(1–4): 162–168, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451162.

Burridge:1966:LFS

[Bur66c] R. Burridge. The Legendre functions of a second kind with complex argument in the theory of wave propagation. Journal of Mathematics and Physics (MIT), 45(1–4):322–330, April 1966. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1966451322.

Bush:1924:NOC

[Bus24] V. Bush. Note on operational calculus. Journal of Mathematics and Physics (MIT), 3(2):95–107, March 1924. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19243295.

Bush:1926:FBM

[Bus26] V. Bush. The force between moving charges. Journal of Mathematics and Physics (MIT), 5(1–4):129–157, January 1926. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192651129.

Busse:1967:STD

[Bus67] F. H. Busse. On the stability of two-dimensional convection in a layer heated from below. Journal of Mathematics and Physics (MIT), 46(1–4):140–150, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461140.

Birkhoff:1965:DEW

[BV65] Garrett Birkhoff and Richard S. Varga. Discretization errors for well-set Cauchy problems. I. Journal of Mathematics and Physics (MIT), 44(1– 4):1–23, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19654411.

Born:1926:NFL

[BW26] Max Born and Norbert Wiener. A new formulation of the laws of quantization of periodic and aperiodic phenomena. Journal of Mathematics and Physics (MIT), 5(1–4):84–98, January 1926. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19265184.

Budiansky:1961:TLS

[BW61] Bernard Budiansky and Tai Te Wu. Transfer of load to a sheet from a rivet-attached stiffener. Journal of Mathematics and Physics (MIT), 40(1–4):142–162, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1961401142.

Birkhoff:1950:NQA

[BY50] Garrett Birkhoff and David Young. Numerical quadrature of analytic and harmonic functions. Journal of Mathematics and Physics (MIT), 29(1–4):217–221, April 1950. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1950291217.

Cameron:1936:APP

[Cam36] Robert H. Cameron. Almost periodic properties of bounded solutions of linear differential equations with almost periodic coefficients. Journal of Mathematics and Physics (MIT), 15(1–4):73–81, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193615173.

Cameron:1938:DVA

[Cam38] R H. Cameron. The distribution of values of an analytic almost periodic function in equally spaced circles. Journal of Mathematics and Physics (MIT), 17(1–4):214–217, April 1938. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1938171214.

Cameron:1940:QIT

[Cam40] R. H. Cameron. Quadratures involving trigonometric sums. Journal of Mathematics and Physics (MIT), 19(1–4):161–166, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1940191161.

Cameron:1942:QCE

[Cam42] R. H. Cameron. Quadratic convolution equations. Journal of Mathematics and Physics (MIT), 21(1–4):57–62, April 1942. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194221157.

Cambi:1947:CEI

[Cam47] Enzo Cambi. Complete elliptic integrals of complex Legendrian modulus. Journal of Mathematics and Physics (MIT), 26(1–4):234–245, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261234.

Cambi:1956:ILT

[Cam56] Enzo Cambi. Inverse Laplace transforms expressed as Neumann series. Journal of Mathematics and Physics (MIT), 35(1–4):114–122, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351114.

Cameron:1960:FIS

[Cam60] R. H. Cameron. A family of integrals serving to connect the Wiener and Feynman integrals. Journal of Mathematics and Physics (MIT), 39(1–4):126–140, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391126.

Carrillo:1942:DES

[Car42a] Nabor Carrillo. Differential equation of a sliding surface in an ideal saturated plastic soil. Journal of Mathematics and Physics (MIT), 21(1– 4):6–9, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19422116.

Carrillo:1942:STT

[Car42b] Nabor Carrillo. Simple two and three dimensional case in the theory of consolidation of soils. Journal of Mathematics and Physics (MIT), 21(1–4):1–5, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19422111.

Carrier:1947:BER

[Car47] G. F. Carrier. On the buckling of elastic rings. Journal of Mathematics and Physics (MIT), 26(1–4):94–103, April 1947. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194726194.

Carrier:1948:DEF

[Car48] G. F. Carrier. On the determination of the eigenfunctions of Fredholm equations. Journal of Mathematics and Physics (MIT), 27(1–4):82– 83, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194827182.

Cartwright:1958:PRF

[Car58] D. E. Cartwright. The prediction of a random function, given simultaneous values of its first few derivatives. Journal of Mathematics and Physics (MIT), 37(1–4):229–245, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1958371229.

Carlson:1961:SSB

[Car61] B. C. Carlson. Some series and bounds for incomplete elliptic integrals. Journal of Mathematics and Physics (MIT), 40(1–4):125–134, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401125.

Carlson:1965:CEI

[Car65] B. C. Carlson. On computing elliptic integrals and functions. Journal of Mathematics and Physics (MIT), 44(1–4):36–51, April 1965. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196544136.

Chang:1968:PCS

[CC68] I. C. Chang and I. Dee Chang. Potential of a charged sphere inside a grounded cylindrical tube. Journal of Mathematics and Physics (MIT), 47(1–4):366–370, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471366.

Carrier:1956:UMV

[CD56] G. F. Carrier and R. C. DiPrima. On the unsteady motion of a viscous fluid past a semi-infinite flat plate. Journal of Mathematics and Physics (MIT), 35(1–4):359–383, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351359.

Caughey:1962:BLS

[CD62] T. K. Caughey and J. K. Dienes. The behavior of linear systems with random parametric excitation. Journal of Mathematics and Physics (MIT), 41(1–4):300–318, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1962411300. See comment [KB63].

Cederbaum:1955:IMR

[Ced55] I. Cederbaum. Invariance and mutual relations of electrical network determinants. Journal of Mathematics and Physics (MIT), 34(1–4): 236–244, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341236.

Cederbaum:1957:MAW

[Ced57] I. Cederbaum. Matrices all of whose elements and subdeterminants are 1, −1, or 0. Journal of Mathematics and Physics (MIT), 36(1–4):

351–361, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361351.

Cerulus:1961:ARF

[Cer61] F. Cerulus. Approximations to random flight functions in one, two, and three dimensions. Journal of Mathematics and Physics (MIT), 40(1–4): 277–281, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401277.

Chao:1931:SFS

[Cha31] Robert F. H. Chao. Some formulas for the strength of the magnetic field of a cylindrical coil. Journal of Mathematics and Physics (MIT), 10(1– 4):13–18, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193110113.

Chao:1934:FTD

[Cha34] Robert F. H. Chao. A formal theorem on the derivatives of a series of zonal harmonics. Journal of Mathematics and Physics (MIT), 13(1–4): 41–45, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193413141.

Chako:1957:IRI

[Cha57] Nicholas Chako. On integral relations involving products of spheroidal functions. Journal of Mathematics and Physics (MIT), 36(1–4):62– 73, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195736162.

Chang:1961:SFP

[Cha61] I-Dee Chang. Stokes flow past an axially symmetric body in a cylindrical tube in the presence of a magnetic field. Journal of Mathematics and Physics (MIT), 40(1–4):205–219, April 1961. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1961401205.

Chang:1967:EFP

[Cha67] K. W. Chang. The existence of forced periodic solutions for second order differential equations. Journal of Mathematics and Physics (MIT), 46(1–4):93–98, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746193.

Chester:1961:RBV

[Che61] Clive R. Chester. Reduction of a boundary value problem of the third kind to one of the first kind. Journal of Mathematics and Physics (MIT), 40(1–4):68–71, April 1961. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196140168.

Chen:1964:WPI

[Che64] Yung Ming Chen. Wave propagation in inhomogeneous and discontinuous random media. Journal of Mathematics and Physics (MIT), 43(1–4):314–324, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431314.

Chen:1966:SAF

[Che66] W. T. Chen. Some aspects of a flat elliptical crack under shear stress. Journal of Mathematics and Physics (MIT), 45(1–4):213–223, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451213.

Chuang:1936:NSF

[Chu36] C T. Chuang. On the normalization of the set of functions x^{n}sechx. Journal of Mathematics and Physics (MIT), 15(1–4):264–267, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151264.

Churchill:1954:OCL

[Chu54] R. V. Churchill. The operational calculus of Legendre transforms. Journal of Mathematics and Physics (MIT), 33(1–4):165–178, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331165.

Charnes:1966:NDM

[CK66] A. Charnes and K. Kortanek. A note on the discrete maximum principle and distribution problems. Journal of Mathematics and Physics (MIT), 45(1–4):121–126, April 1966. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1966451121.

Clapp:1937:FPT

[Cla37] M. M. Clapp. Four place table of the Bessel function. Journal of Mathematics and Physics (MIT), 16(1–4):76, April 1937. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193716176.

Clark:1950:TTE

[Cla50] R. A. Clark. On the theory of thin elastic toroidal shells. Journal of Mathematics and Physics (MIT), 29(1–4):146–178, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291146.

Clendenin:1957:MRM

[Cle57] W. W. Clendenin. Modified relaxation method for eigenvalue problems. Journal of Mathematics and Physics (MIT), 36(1–4):74–81, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195736174.

Cameron:1944:WMH

[CM44] R. H. Cameron and W. T. Martin. The Wiener measure of Hilbert neighborhoods in the space of real continuous functions. Journal of Mathematics and Physics (MIT), 23(1–4):195–209, April 1944. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1944231195.

Corrington:1945:TBFa

[CM45] Murlan S. Corrington and William Miehle. Tables of Bessel functions J_{n}(x) for large arguments. Journal of Mathematics and Physics (MIT), 24(1–4):30–50, April 1945. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194524130.

Chahine:1964:I

[CN64] M. T. Chahine and R. Narasimha. The integral. Journal of Mathematics and Physics (MIT), 43(1–4):163–168, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1964431163.

Coburn:1938:NAK

[Cob38] Nathaniel Coburn. A new approach to Kron’s work. Journal of Mathematics and Physics (MIT), 17(1–4):112–122, April 1938. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1938171112.

Coburn:1940:GEH

[Cob40] N. Coburn. Generalized Einstein hypersurfaces of spaces of constant curvature. Journal of Mathematics and Physics (MIT), 19(1–4):140– 152, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1940191140.

Coburn:1942:FFC

[Cob42] N. Coburn. Frenet formulas for curves in unitary space. Journal of Mathematics and Physics (MIT), 21(1–4):10–18, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194221110.

Coburn:1944:BVPa

[Cob44a] N. Coburn. Boundary value problem in plane plasticity. Journal of Mathematics and Physics (MIT), 23(1–4):61–68, April 1944. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194423161.

Coburn:1944:BVPb

[Cob44b] N. Coburn. A boundary value problem in plane plasticity for the Coulomb yield condition. Journal of Mathematics and Physics (MIT), 23(1–4):117–125, April 1944. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1944231117.

Cochran:1967:MMB

[Coc67] James Alan Cochran. The monotonicity of modified Bessel functions with respect to their order. Journal of Mathematics and Physics (MIT), 46(1–4):220–222, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461220.

Coddington:1951:SID

[Cod51] E A. Coddington. The stability of infinite differential systems associated with vortex streets. Journal of Mathematics and Physics (MIT), 30(1–4):171–199, April 1951. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1951301171.

Coddington:1965:SID

[Cod65] E. A. Coddington. The stability of infinite differential systems associated with vortex streets, II. Journal of Mathematics and Physics (MIT), 44(1–4):89–94, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196544189.

Cohen:1964:ZBF

[Coh64] Donald S. Cohen. Zeros of Bessel functions and eigenvalues of nonself-adjoint boundary value problems. Journal of Mathematics and Physics (MIT), 43(1–4):133–139, April 1964. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1964431133. See errata [Coh65].

Cohen:1965:EZB

[Coh65] Donald S. Cohen. Errata: Zeros of Bessel Functions and Eigenvalues of Non-self-adjoint Boundary Value Problems. Journal of Mathematics and Physics (MIT), 44(1–4):410, April 1965. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1965441410. See [Coh64].

Cohen:1966:DCS

[Coh66] H. Cohen. Dislocations in couple-stress elasticity. Journal of Mathematics and Physics (MIT), 45(1–4):35–44, April 1966. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196645135.

Cole:1951:DFW

[Col51] Julian D. Cole. Drag of a finite wedge at high subsonic speeds. Journal of Mathematics and Physics (MIT), 30(1–4):79–93, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195130179.

Contri:1964:ESP

[Con64] L. Contri. Exact solution of a problem of rectangular elastic plates of variable thickness. Journal of Mathematics and Physics (MIT), 43(1– 4):72–74, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196443172.

Cooke:1956:OII

[Coo56] J. C. Cooke. Osculatory interpolation and integration. Journal of Mathematics and Physics (MIT), 35(1–4):394–400, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1956351394.

Cook:1957:CMW

[Coo57] J. M. Cook. Convergence to the Møller wave-matrix. Journal of Mathematics and Physics (MIT), 36(1–4):82–87, April 1957. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195736182.

Corrington:1945:TBFb

[Cor45] Murlan S. Corrington. Table of Bessel functions J_{n}(1000). Journal of Mathematics and Physics (MIT), 24(1–4):144–147, April 1945. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1945241144.

Court:1941:TMM

[Cou41] Louis M. Court. A theorem on maxima and minima with an application to differential equations. Journal of Mathematics and Physics (MIT), 20(1–4):99–106, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194120199.

Covert:1958:NAC

[Cov58] Eugene E. Covert. A note on an approximate calculation of Green’s functions for built-up bodies. Journal of Mathematics and Physics (MIT), 37(1–4):58–65, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195837158.

Cowan:1965:PVD

[Cow65] R. W. Cowan. On particular values of the derivatives of the gamma function. Journal of Mathematics and Physics (MIT), 44(1–4):271– 274, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441271.

Coxeter:1933:RCP

[Cox33] H. S. M. Coxeter. Regular compound polytopes in more than four dimensions. Journal of Mathematics and Physics (MIT), 12(1–4):334– 345, April 1933. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1933121334.

Crandall:1951:RME

[Cra51] Stephen H. Crandall. On a relaxation method for eigenvalue problems. Journal of Mathematics and Physics (MIT), 30(1–4):140–145, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1951301140.

Crandall:1953:SCP

[Cra53] Stephen H. Crandall. On a stability criterion for partial difference equations. Journal of Mathematics and Physics (MIT), 32(1–4):80– 81, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195332180.

Craig:1955:SIP

[Cra55] Edward J. Craig. The N-step iteration procedures. Journal of Mathematics and Physics (MIT), 34(1–4):64–73, April 1955. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195534164.

Crane:1959:NSP

[Cra59] L. J. Crane. A note on Stewartson’s paper “On Asymptotic Expansions in the Theory of Boundary Layers”. Journal of Mathematics and Physics (MIT), 38(1–4):172–174, April 1959. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1959381172. See [Ste57].

Creedy:1935:EWM

[Cre35] F. Creedy. The equivalent wave method — the use of vectors in studying electrical transients. Journal of Mathematics and Physics (MIT), 14(1–4):291–324, April 1935. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1935141291.

Crout:1928:MDF

[Cro28] Prescott D. Crout. A method for deriving formulas for the approximate calculation of integrals. Journal of Mathematics and Physics (MIT), 7(1–4):126–159, April 1928. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192871126.

Crout:1929:AIA

[Cro29a] Prescott D. Crout. An application of the invariant area properties of algebraic polynomials to the derivation of formulas for mechanical integration. Journal of Mathematics and Physics (MIT), 8(1–4):200– 215, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192981200.

Crout:1929:MDFb

[Cro29b] Prescott D. Crout. A method for deriving formulas for the approximate calculation of integrals (continued). Journal of Mathematics and Physics (MIT), 8(1–4):119–128, April 1929. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192981119.

Crout:1929:MDFa

[Cro29c] Prescott D. Crout. A method for deriving formulas of interpolation. Journal of Mathematics and Physics (MIT), 8(1–4):18–55, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19298118.

Crout:1930:AFI

[Cro30] Prescott D. Crout. The approximation of functions and integrals by a linear combination of functions. Journal of Mathematics and Physics (MIT), 9(1–4):278–314, April 1930. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193091278.

Crout:1934:CSP

[Cro34] Prescott D. Crout. The criterion for a stationary point of one of a set of implicit functions. Journal of Mathematics and Physics (MIT), 13(1–4):387–398, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1934131387.

Crout:1936:AKT

[Cro36] Prescott D. Crout. An application of kinetic theory to the problems of evaporation and sublimation of monatomic gases. Journal of Mathematics and Physics (MIT), 15(1–4):1–54, April 1936. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19361511.

Crout:1940:APA

[Cro40] Prescott D. Crout. An application of polynomial approximation to the solution of integral equations arising in physical problems. Journal of Mathematics and Physics (MIT), 19(1–4):34–92, April 1940. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194019134.

Crossley:1951:HFN

[Cro51] F. R. Erskine Crossley. A hyperelliptic function as a non-linear oscillation. Journal of Mathematics and Physics (MIT), 30(1–4):214– 225, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1951301214.

Crout:1959:UBD

[Cro59] Prescott D. Crout. An upper bound on derivative errors in Lagrange polynomial approximation. Journal of Mathematics and Physics (MIT), 38(1–4):327–330, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1959381327.

Crout:1960:POZ

[Cro60] Prescott D. Crout. A procedure for obtaining the zeros of functions or their derivatives by Lagrangean interpolation. Journal of Mathematics and Physics (MIT), 39(1–4):141–150, April 1960. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1960391141.

Crout:1961:LSP

[Cro61] Prescott D. Crout. Least square procedure for solving homogeneous integral equations. Journal of Mathematics and Physics (MIT), 40(1–4): 103–115, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401103.

Crowson:1964:ASO

[Cro64] Henry L. Crowson. An analysis of a second order linear ordinary differential equation with five regular singular points. Journal of Mathematics and Physics (MIT), 43(1–4):38–44, April 1964. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196443138.

Crowson:1965:HSS

[Cro65] Henry L. Crowson. Hypergeometric solutions of a second-order linear ordinary differential equation with n-regular singular points. Journal of Mathematics and Physics (MIT), 44(1–4):384–390, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441384.

Chu:1941:ESW

[CS41] L. J. Chu and J. A. Stratton. Elliptic and spheroidal wave functions. Journal of Mathematics and Physics (MIT), 20(1–4):259–309, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201259.

Carrus:1950:TRI

[CT50] Pierre A. Carrus and Charlotte G. Treuenfels. Tables of roots and incomplete integrals of associated Legendre functions of fractional orders. Journal of Mathematics and Physics (MIT), 29(1–4):282–299, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291282.

Chen:1967:PTW

[CT67] Y. M. Chen and C. L. Tien. Penetration of temperature waves in a random medium. Journal of Mathematics and Physics (MIT), 46(1–4): 188–194, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461188.

Curtiss:1953:MCM

[Cur53] J. H. Curtiss. “Monte Carlo” methods for the iteration of linear operators. Journal of Mathematics and Physics (MIT), 32(1–4):209–232, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321209.

Chang:1952:STE

[CW52] Chieh-Chien Chang and Jack Werner. A solution of the telegraph equation with application to two dimensional supersonic shear flow. Journal of Mathematics and Physics (MIT), 31(1–4):91–101, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195231191.

Dames:1959:SCN

[Dam59] Ralph T. Dames. Stability and convergence for a numerical solution of the Goursat problem. Journal of Mathematics and Physics (MIT), 38(1–4):42–67, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195938142.

Darlington:1939:SRP

[Dar39] S. Darlington. Synthesis of reactance 4-poles which produce prescribed insertion loss characteristics: Including special applications to filter design. Journal of Mathematics and Physics (MIT), 18(1–4):257–353, April 1939. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1939181257.

Davis:1957:ROT

[Dav57] Robert B. Davis. A reduction-of-order theorem. Journal of Mathematics and Physics (MIT), 36(1–4):164–166, April 1957. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1957361164.

deBoor:1962:BSI

[dB62] Carl de Boor. Bicubic spline interpolation. Journal of Mathematics and Physics (MIT), 41(1–4):212–218, April 1962. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1962411212.

Diaz:1947:SIM

[DD47] J. B. Diaz and J. B. Diaz. Schwarz’ inequality and the methods of Rayleigh–Ritz and Trefftz. Journal of Mathematics and Physics (MIT), 26(1–4):133–136, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1947261133.

Davids:1961:ZOS

[DD61] Norman Davids and Donald E. Dawson. Zero order spherical harmonic stress distributions. Journal of Mathematics and Physics (MIT), 40(1– 4):46–54, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196140146.

Donder:1926:IRI

[De 26] Th. De Donder. Introduction to the relativistic interpretation of the quantum theory. Journal of Mathematics and Physics (MIT), 5(1–4):251–261, January 1926. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192651251.

DeCicco:1947:NPT

[De 47] John De Cicco. New proofs of the theorems of Kasner concerning the infinitesimal contact transformations of mechanics. Journal of Mathematics and Physics (MIT), 26(1–4):104–109, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1947261104.

Debye:1925:NSX

[Deb25] P. Debye. Note on the scattering of X-rays. Journal of Mathematics and Physics (MIT), 4(1–4):133–147, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192541133.

DeClaris:1956:ETD

[DeC56] N. DeClaris. An existence theorem for driving-point impedance functions. Journal of Mathematics and Physics (MIT), 35(1–4):83–88, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195635183.

Deissler:1968:ECT

[Dei68] R. G. Deissler. Effects of combined two-dimensional shear and normal strain on weak locally homogeneous turbulence and heat transfer. Journal of Mathematics and Physics (MIT), 47(1–4):310–326, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1968471310.

Dettman:1964:PNS

[Det64] John W. Dettman. Perturbation by non-simple constraints. Journal of Mathematics and Physics (MIT), 43(1–4):304–308, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431304.

Diaz:1948:ULB

[DG48] J. B. Diaz and H. J. Greenberg. Upper and lower bounds for the solution of the first bi-harmonic boundary value problem. Journal of Mathematics and Physics (MIT), 27(1–4):193–201, April 1948. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1948271193.

DiPrima:1954:NAS

[Di 54] Richard C. Di Prima. A note on the asymptotic solutions of the equation of hydrodynamic stability. Journal of Mathematics and Physics (MIT), 33(1–4):249–257, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1954331249.

Diaz:1959:TDF

[Dia59] J. B. Diaz. Two-dimensional flow at high subsonic speeds past wedges in channels with parallel walls. Journal of Mathematics and Physics (MIT), 38(1–4):75–76, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195938175.

Duffin:1956:FRS

[DK56] R. J. Duffin and Elsa Keitzer. Formulae relating some equivalent networks. Journal of Mathematics and Physics (MIT), 35(1–4):72–82, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195635172.

Dwight:1930:MIR

[DL30] H. B. Dwight and T. Y. Lu. Mutual inductance and repulsion of two adjacent disk coils. Journal of Mathematics and Physics (MIT), 9(1–4): 315–319, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193091315.

Donaldson:1967:IRE

[Don67] James A. Donaldson. Integral representations of the extended Airy integral type for the modified Bessel function. Journal of Mathematics and Physics (MIT), 46(1–4):111–114, April 1967. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1967461111.

Douglas:1931:MTK

[Dou31a] Jesse Douglas. The mapping theorem of Koebe and the problem of plateau. Journal of Mathematics and Physics (MIT), 10(1–4):106– 130, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1931101106.

Douglas:1931:PPT

[Dou31b] Jesse Douglas. The problem of plateau for two contours. Journal of Mathematics and Physics (MIT), 10(1–4):315–359, April 1931. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1931101315.  See correction

[Dou32].

Douglass:1931:SED

[Dou31c] Raymond D. Douglass. Stirling expansions derived by means of finite De La Vallée-Poussin summation. Journal of Mathematics and Physics (MIT), 10(1–4):131–156, April 1931. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1931101131.

Douglas:1932:CPP

[Dou32] Jesse Douglas. Corrections to the paper “The Problem of Plateau for Two Contours”. Journal of Mathematics and Physics (MIT), 11(1–4):

212, April 1932. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1932111212.

See [Dou31b].

Douglas:1936:MSG

[Dou36a] Jesse Douglas. Minimal surfaces of general topological structure with any finite number of assigned boundaries. Journal of Mathematics and Physics (MIT), 15(1–4):105–123, April 1936. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1936151105.

Douglas:1936:SNR

[Dou36b] Jesse Douglas. Some new results in the problem of plateau. Journal of Mathematics and Physics (MIT), 15(1–4):55–64, April 1936. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193615155.

Douglas:1940:GPC

[Dou40] Jesse Douglas. Geometry of polygons in the complex plane. Journal of Mathematics and Physics (MIT), 19(1–4):93–130, April 1940. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194019193.

Douglas:1956:SDE

[Dou56] Jim Douglas, Jr. The solution of the diffusion equation by a high order correct difference equation. Journal of Mathematics and Physics (MIT), 35(1–4):145–151, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351145.

Douglas:1959:USC

[Dou59] Jim Douglas, Jr. An unusual stability criterion for a difference analogue of the heat equation. Journal of Mathematics and Physics (MIT), 38(1–4):150–152, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1959381150.

Doyle:1955:TDI

[Doy55] Thomas C. Doyle. Topological and dynamical invariant theory of an electrical network. Journal of Mathematics and Physics (MIT), 34(1– 4):81–94, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195534181.

Doyle:1957:HOI

[Doy57] Thomas C. Doyle. Higher order invariants of stress or deformation tensors. Journal of Mathematics and Physics (MIT), 36(1–4):297–305, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361297.

Doyle:1958:LBL

[Doy58] Worthie Doyle. Lossless Butterworth ladder networks operating between arbitrary resistances. Journal of Mathematics and Physics (MIT), 37(1–4):29–37, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195837129.

Diaz:1952:ULB

[DR52] J. B. Diaz and R. C. Roberts. Upper and lower bounds for the numerical solution of the Dirichlet difference boundary value problem. Journal of Mathematics and Physics (MIT), 31(1–4):184–191, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1952311184.

Drazin:1960:SBL

[Dra60] Philip G. Drazin. Stability of a broken-line jet in a parallel magnetic field. Journal of Mathematics and Physics (MIT), 39(1–4):49– 53, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196039149.

Dwight:1930:MIS

[DS30] H. B. Dwight and P. W. Sayles. Mutual inductance of short concentric solenoids. Journal of Mathematics and Physics (MIT), 9(1–4):162– 165, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193091162.

Duffin:1958:ASD

[DS58] R. J. Duffin and W. D. Serbyn. Approximate solution of differential equations by a variational method. Journal of Mathematics and Physics (MIT), 37(1–4):162–168, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1958371162.

Duffin:1948:QHC

[Duf48] R. J. Duffin. On a question of Hadamard concerning super-biharmonic functions. Journal of Mathematics and Physics (MIT), 27(1–4):253– 258, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271253.

Duffin:1952:NLV

[Duf52] R. J. Duffin. Nodal lines of a vibrating plate. Journal of Mathematics and Physics (MIT), 31(1–4):294–299, April 1952. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1952311294.

Dunninger:1967:UCT

[Dun67] Dennis R. Dunninger. Uniqueness and comparison theorems for harmonic functions under boundary conditions. Journal of Mathematics and Physics (MIT), 46(1–4):299–310, April 1967. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1967461299.

Dutka:1944:THS

[Dut44] Jacques Dutka. Transversality in higher space. Journal of Mathematics and Physics (MIT), 23(1–4):126–133, April 1944. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1944231126.

Dwight:1931:CRG

[Dwi31] Herbert Bristol Dwight. The calculation of resistances to ground and of capacitance. Journal of Mathematics and Physics (MIT), 10(1–4): 50–74, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193110150.

Dwight:1946:TBF

[Dwi46] H. B. Dwight. Table of Bessel functions and derivative J_{2}, J_{1}′, J_{2}′, N_{2}, N_{1}′, N_{2}′. Journal of Mathematics and Physics (MIT), 25(1–4):93–95, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194625193.

Dwight:1948:TRN

[Dwi48] H. B. Dwight. Table of roots for natural frequencies in coaxial type cavities. Journal of Mathematics and Physics (MIT), 27(1–4):84–89, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194827184.

Eaves:1953:NSM

[Eav53] J. C. Eaves. A note on sets of matrices simultaneously reducible to the triangular skeleton. Journal of Mathematics and Physics (MIT), 32(1–4):302–306, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321302.

Evans:1955:SCV

[EBK55] G. W. Evans, R. Brousseau, and R. Keirstead. Stability considerations for various difference equations derived for the linear heat conduction equation. Journal of Mathematics and Physics (MIT), 34(1–4):267–285, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1955341267. See errata [EBK57].

Evans:1957:ESC

[EBK57] G. W. Evans, R. Brousseau, and R. Keirstead. Errata for “Stability Considerations for Various Difference Equations Derived for the Linear Heat Conduction Equation”. Journal of Mathematics and Physics (MIT), 36(1–4):294–295, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361294. See [EBK55].

Erdos:1938:ELU

[EGV38] P. Erdős, T. Grunwald, and E. Vázsonyi. Über Euler-Linien Unendlicher Graphen. (German) [On Euler lines of infinite graphs]. Journal of Mathematics and Physics (MIT), 17(1–4):59–75, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193817159.

Evans:1966:SSE

[EH66] D. J. Evans and V. G. Hart. On a system of simplified equations for finite axisymmetric deflections of an annular plate. Journal of Mathematics and Physics (MIT), 45(1–4):367–369, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451367.

Edels:1962:NSA

[EHY62] H. Edels, K. Hearne, and A. Young. Numerical solutions of the Abel integral equation. Journal of Mathematics and Physics (MIT), 41(1–4): 62–75, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196241162.

Eichenberger:1953:SFW

[Eic53] Hans P. Eichenberger. Secondary flow within a bend. Journal of Mathematics and Physics (MIT), 32(1–4):34–42, April 1953. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195332134.

Eiseman:1935:NAS

[Eis35] B. J. Eiseman, Jr. Note on the absorption spectra of oxygen at high concentration. Journal of Mathematics and Physics (MIT), 14(1–4): 260–261, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141260.

Ellinwood:1967:AHF

[Ell67] John W. Ellinwood. Asymptotic hypersonic-flow theory for blunted slender cones and wedges. Journal of Mathematics and Physics (MIT), 46(1–4):281–298, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461281.

Ergen:1957:ALS

[ELN57] W. K. Ergen, H. J. Lipkin, and J. A. Nohel. Applications of Liapounov’s second method in reactor dynamics. Journal of Mathematics and Physics (MIT), 36(1–4):36–48, April 1957. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195736136.

Elrod:1958:NST

[Elr58] Harold G. Elrod, Jr. Note on a solution of the telegraphist’s equation applicable to supersonic shear flow. Journal of Mathematics and Physics (MIT), 37(1–4):66–68, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195837166.

Enig:1961:SCN

[Eni61] Julius W. Enig. Stability criteria for numerical solutions in unsteady two-dimensional cylindrical Lagrangian flow. Journal of Mathematics and Physics (MIT), 40(1–4):23–32, April 1961. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196140123.

Epstein:1938:PTC

[Eps38] Benjamin Epstein. A proof of a theorem of Carleman. Journal of Mathematics and Physics (MIT), 17(1–4):208–213, April 1938. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1938171208.

Epstein:1939:FRS

[Eps39] Leo F. Epstein. A function related to the series for. Journal of Mathematics and Physics (MIT), 18(1–4):153–173, April 1939. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1939181153.

Epstein:1942:TEV

[Eps42] Paul S. Epstein. On the theory of elastic vibrations in plates and shells. Journal of Mathematics and Physics (MIT), 21(1–4):198–209, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1942211198.

Erber:1959:SNH

[Erb59] Thomas Erber. Stirling numbers and hypergeometric functions. Journal of Mathematics and Physics (MIT), 38(1–4):331–334, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1959381331.

Erdos:1943:CTS

[Erd43] P. Erdős. On the convergence of trigonometric series. Journal of Mathematics and Physics (MIT), 22(1–4):37–39, April 1943. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194322137.

Ericksen:1952:UGF

[Eri52] J. L. Ericksen. On the uniqueness of gas flows. Journal of Mathematics and Physics (MIT), 31(1–4):63–68, April 1952. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195231163.

Ericksen:1955:DPE

[Eri55a] J. L. Ericksen. Deformations possible in every compressible, isotropic, perfectly elastic material. Journal of Mathematics and Physics (MIT), 34(1–4):126–128, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341126.

Ericksen:1955:SSP

[Eri55b] J. L. Ericksen. Singular surfaces in plasticity. Journal of Mathematics and Physics (MIT), 34(1–4):74–79, April 1955. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195534174.

Epstein:1967:TDD

[ES67] Bernard Epstein and Harvey Senter. The three-dimensional Dirichlet problem associated with a plane lamina. Journal of Mathematics and Physics (MIT), 46(1–4):425–439, April 1967. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1967461425.

Estes:1934:FGC

[Est34] J. G. Estes. Formulae giving the change in Green’s function and in the conjugate function. Journal of Mathematics and Physics (MIT), 13(1–4):249–252, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1934131249.

Estes:1935:LMA

[Est35] J. G. Estes. The lift and the moment of an arbitrary aerofoil-joukovsky potential. Journal of Mathematics and Physics (MIT), 14(1–4):100– 124, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141100.

Fabry:1925:OAM

[Fab25] Charles Fabry. Ozone as an absorbing material for radiations in the atmosphere. Journal of Mathematics and Physics (MIT), 4(1–4):1–20, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1925411.

Fabian:1936:FC

[Fab36] William Fabian. Fractional calculus. Journal of Mathematics and Physics (MIT), 15(1–4):83–89, April 1936. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm193615183.

Fante:1967:FCL

[Fan67] R. L. Fante. The fields of a current loop inside an infinite cylindrical column. Journal of Mathematics and Physics (MIT), 46(1–4):325– 332, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461325.

Fejer:1934:NPA

[Fej34] Leopold Fejér. On new properties of the arithmetical means of the partial sums of Fourier series. Journal of Mathematics and Physics (MIT), 13(1–4):1–17, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19341311.

Felsen:1956:SDI

[Fel56] Leopold B. Felsen. Some definite integrals involving conical functions. Journal of Mathematics and Physics (MIT), 35(1–4):177–178, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351177.

Felsen:1958:SNT

[Fel58] L. B. Felsen. Some new transform theorems involving Legendre functions. Journal of Mathematics and Physics (MIT), 37(1–4):188–191, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371188.

Fempl:1958:SPH

[Fem58] S. Fempl. Some properties of Heuman’s function Λ_{0}. Journal of Mathematics and Physics (MIT), 37(1–4):137–142, April 1958. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1958371137.

Fettis:1954:CIF

[Fet54] Henry E. Fettis. On the calculation of integrals of the form. Journal of Mathematics and Physics (MIT), 33(1–4):283–289, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331283.

Fettis:1957:LTI

[Fet57a] Henry E. Fettis. Lommel-type integrals involving three Bessel functions. Journal of Mathematics and Physics (MIT), 36(1–4):88–95, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195736188.

Fettis:1957:CFJ

[Fet57b] Henry E. Fettis. On the calculation of the function j_{0}(z, θ) for large values of “z”. Journal of Mathematics and Physics (MIT), 36(1–4): 279–283, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361279.

Fialkow:1951:TFR

[FG51] Aaron Fialkow and Irving Gerst. The transfer function of an R-C ladder network. Journal of Mathematics and Physics (MIT), 30(1–4):

49–72, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195130149.

Fialkow:1955:ISM

[FG55] Aaron Fialkow and Irving Gerst. Impedance synthesis without minimization. Journal of Mathematics and Physics (MIT), 34(1–4):160– 168, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341160.

Fichera:1950:SGI

[Fic50] Gaetano Fichera. On some general integration methods employed in connection with linear differential equations. Journal of Mathematics and Physics (MIT), 29(1–4):59–68, April 1950. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195029159.

Finston:1955:TWS

[Fin55] Morton Finston. A thin wedge in a slightly non-uniform supersonic flow. Journal of Mathematics and Physics (MIT), 34(1–4):328–334, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341328.

Fish:1935:PBV

[Fis35] Maurice Julius Fish. A particular boundary value problem. Journal of Mathematics and Physics (MIT), 14(1–4):262–273, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141262.

Fisher:1963:CSC

[Fis63] Donald D. Fisher. Calculation of subsonic cavities with sonic free streamlines. Journal of Mathematics and Physics (MIT), 42(1–4):14– 26, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196342114.

Flanagan:1954:CAM

[Fla54] Joseph E. Flanagan. Coding to achieve Markov type redundancy. Journal of Mathematics and Physics (MIT), 33(1–4):258–268, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331258.

Fleischer:1958:CCC

[Fle58] Isidore Fleischer. The central concepts of communication theory for infinite alphabets. Journal of Mathematics and Physics (MIT), 37(1–4):

223–228, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371223.

Farnell:1950:FPS

[FLL50] A. B. Farnell, C. E. Langenhop, and N. Levinson. Forced periodic solutions of a stable non-linear system of differential equations. Journal of Mathematics and Physics (MIT), 29(1–4):300–302, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291300.

Franklin:1930:SLP

[FM30] P. Franklin and C. L. E. Moore. Systems of linear partial differential equations. Journal of Mathematics and Physics (MIT), 9(1–4): 22–46, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19309122.

Franklin:1931:GP

[FM31] P. Franklin and C. L. E. Moore. Geodesics of Pfaffians. Journal of Mathematics and Physics (MIT), 10(1–4):157–190, April 1931. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1931101157.

Folias:1965:SCS

[Fol65] Efthymios S. Folias. The stresses in a cracked spherical shell. Journal of Mathematics and Physics (MIT), 44(1–4):164–176, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441164.

Fort:1954:LVN

[For54] Tomlinson Fort. The loaded vibrating net and resulting boundary-value problems for a partial difference equation of the second order. Journal of Mathematics and Physics (MIT), 33(1–4):94–104, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195433194.

Fort:1956:LCV

[For56] Tomlinson Fort. Limits of the characteristic values for certain boundary-value problems associated with difference equations. Journal of Mathematics and Physics (MIT), 35(1–4):401–407, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1956351401.

Ford:1958:CII

[For58] F. A. J. Ford. On certain indefinite integrals involving Bessel functions. Journal of Mathematics and Physics (MIT), 37(1–4):157–161, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371157.

Fox:1951:DPV

[Fox51] Charles Fox. The determination of position and velocity on the Earth’s surface. Journal of Mathematics and Physics (MIT), 30(1–4):146–155, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1951301146.

Fox:1955:PTW

[Fox55] Phyllis A. Fox. Perturbation theory of wave propagation based on the method of characteristics. Journal of Mathematics and Physics (MIT), 34(1–4):133–151, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341133.

Fox:1957:NSE

[FR57] P. Fox and A. Ralston. On the numerical solution of the equations for spherical waves of finite amplitude, I. Journal of Mathematics and Physics (MIT), 36(1–4):313–328, April 1957. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1957361313.

Franklin:1925:ECN

[Fra25a] Philip Franklin. The electric currents in a network. Journal of Mathematics and Physics (MIT), 4(1–4):97–102, January 1925. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19254197.

Franklin:1925:WAT

[Fra25b] Philip Franklin. The Weierstrass approximation theorem. Journal of Mathematics and Physics (MIT), 4(1–4):148–152, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192541148.

Franklin:1926:C

[Fra26a] Philip Franklin. A correction. Journal of Mathematics and Physics (MIT), 5(1–4):236–237, January 1926. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192651236.

Franklin:1926:ETA

[Fra26b] Philip Franklin. The elementary theory of almost periodic functions of two variables. Journal of Mathematics and Physics (MIT), 5(1–4): 40–54, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19265140.

Franklin:1926:FTA

[Fra26c] Philip Franklin. The fundamental theorem of almost periodic functions of two variables. Journal of Mathematics and Physics (MIT), 5(1–4):201–236, January 1926. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192651201.

Franklin:1927:GCE

[Fra27a] Philip Franklin. A geometric characterization of equipotential and stream lines. Journal of Mathematics and Physics (MIT), 6(1–4):191– 208, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761191.

Franklin:1927:SLT

[Fra27b] Philip Franklin. The Simson lines of a triangle, the three cusped hypocycloid and the Morley triangles. Journal of Mathematics and Physics (MIT), 6(1–4):50–61, April 1927. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19276150.

Frank:1928:MCQ

[Fra28a] N. H. Frank. On the methods of classical and quantum mechanics. Journal of Mathematics and Physics (MIT), 7(1–4):298–304, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192871298.

Franklin:1928:QDS

[Fra28b] Philip Franklin. A qualitative definition of the sub- and superharmonic functions. Journal of Mathematics and Physics (MIT), 7(1– 4):86–92, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19287186.

Franklin:1929:CFO

[Fra29a] Philip Franklin. Classes of functions, orthogonal on an infinite interval, having the power of the continuum. Journal of Mathematics and Physics (MIT), 8(1–4):74–79, April 1929. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19298174.

Franklin:1929:ES

[Fra29b] Philip Franklin. Equimomental systems. Journal of Mathematics and Physics (MIT), 8(1–4):129–140, April 1929. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192981129.

Franklin:1929:OQM

[Fra29c] Philip Franklin. The operators of quantum mechanics. Journal of Mathematics and Physics (MIT), 8(1–4):235–245, April 1929. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192981235.

Franklin:1930:DSI

[Fra30] Philip Franklin. Dynamical systems with integrals quadratic in the velocities. Journal of Mathematics and Physics (MIT), 9(1–4):333– 344, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193091333.

Franklin:1931:AME

[Fra31] Philip Franklin. Algebraic matric equations. Journal of Mathematics and Physics (MIT), 10(1–4):289–314, April 1931. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1931101289.

Franklin:1934:RPN

[Fra34a] Philip Franklin. Regions of positive and negative curvature on closed surfaces. Journal of Mathematics and Physics (MIT), 13(1–4):253– 260, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1934131253.

Franklin:1934:SCP

[Fra34b] Philip Franklin. A six color problem. Journal of Mathematics and Physics (MIT), 13(1–4):363–369, April 1934. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1934131363.

Franklin:1935:GPS

[Fra35a] Philip Franklin. Geodesics on polyhedral surfaces. Journal of Mathematics and Physics (MIT), 14(1–4):274–278, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1935141274.

Franklin:1935:MDL

[Fra35b] Philip Franklin. On Minkowski’s definition of length and area. Journal of Mathematics and Physics (MIT), 14(1–4):179–185, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141179.

Franklin:1937:NFC

[Fra37] Philip Franklin. Note on the four color problem. Journal of Mathematics and Physics (MIT), 16(1–4):172–184, April 1937. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1937161172.

Franklin:1944:MF

[Fra44] Philip Franklin. Measurable functions. Journal of Mathematics and Physics (MIT), 23(1–4):24–44, April 1944. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm194423124.

Frank:1955:CRE

[Fra55] Evelyn Frank. On the calculation of the roots of equations. Journal of Mathematics and Physics (MIT), 34(1–4):187–197, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1955341187.

Franklin:1958:NSD

[Fra58] J. N. Franklin. Numerical stability in digital and analog computation for diffusion problems. Journal of Mathematics and Physics (MIT), 37(1–4):305–315, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1958371305.

Friedman:1935:ATT

[Fri35] Bernard Friedman. Applications of Tauberian theorems to the study of entire functions. Journal of Mathematics and Physics (MIT), 14(1–4): 232–240, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141232.

Friedmann:1956:TES

[Fri56] Norman E. Friedmann. The truncation error in a semi-discrete analog of the heat equation. Journal of Mathematics and Physics (MIT), 35(1–4):299–308, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351299.

Fry:1936:CIC

[Fry36] Thornton C. Fry. Consistency of independent countings as a criterion for completeness. Journal of Mathematics and Physics (MIT), 15(1–4): 211–217, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151211.

Shen:1952:EVE

[fS52] Shan fu Shen. An estimate of viscosity effect on the hypersonic flow over an insulated wedge. Journal of Mathematics and Physics (MIT), 31(1–4):192–205, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1952311192.

Forsythe:1955:SFC

[FS55] George E. Forsythe and Louise W. Straus. The Souriau-frame characteristic equation algorithm on a digital computer. Journal of Mathematics and Physics (MIT), 34(1–4):152–156, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341152.

Fan:1964:ADM

[FW64] Liang-Tseng Fan and Chid-Sen Wang. The application of the discrete maximum principle to a transportation problem. Journal of Mathematics and Physics (MIT), 43(1–4):255–260, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1964431255.

Gardner:1935:RVS

[Gar35] George S. Gardner. Reaction velocity in the system: SOLID_{I} + GAS → SOLID_{II}. Journal of Mathematics and Physics (MIT), 14(1–4):241– 259, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141241.

Gardner:1937:RVS

[Gar37] George S. Gardner. Reaction velocity in the system: Reaction velocity in the system: Solid_{I} + gas → Solid_{II}: II. flow systems. Journal of Mathematics and Physics (MIT), 16(1–4):77–84, April 1937. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193716177.

Garabedian:1957:NCD

[Gar57] P. R. Garabedian. Numerical construction of detached shock waves. Journal of Mathematics and Physics (MIT), 36(1–4):192– 205, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361192.

See note [Lin57].

Garsia:1960:SME

[Gar60] A. Garsia. Some models of explosive line detonators. Journal of Mathematics and Physics (MIT), 39(1–4):54–57, April 1960. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196039154.

Gautschi:1959:SEI

[Gau59] Walter Gautschi. Some elementary inequalities relating to the gamma and incomplete gamma function. Journal of Mathematics and Physics (MIT), 38(1–4):77–81, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195938177.

Giddings:1936:AKT

[GC36] H. A. Giddings and P. D. Crout. An application of kinetic theory to the problems of evaporation and sublimation of gases having more than one atom per molecule. Journal of Mathematics and Physics (MIT), 15(1–4):124–178, April 1936. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1936151124.

Gray:1965:CPI

[GC65] A. H. Gray, Jr. and T. K. Caughey. A controversy in problems involving random parametric excitation. Journal of Mathematics and Physics (MIT), 44(1–4):288–296, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1965441288.

Greenwood:1949:NIW

[GD49] Robert E. Greenwood and Masil B. Danford. Numerical integration with a weight function x. Journal of Mathematics and Physics (MIT), 28(1–4):99–106, April 1949. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194928199.

Gear:1964:SSB

[Gea64] C. W. Gear. Singular solutions at boundary intersections in two dimensional quasi-linear homogeneous partial differential equations. Journal of Mathematics and Physics (MIT), 43(1–4):100–110, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431100.

Geiringer:1947:CHT

[Gei47] Hilda Geiringer. Contribution to the heredity theory of multivalents. Journal of Mathematics and Physics (MIT), 26(1–4):246–278, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261246.

Gershun:1939:LF

[Ger39] A. Gershun. The light field. Journal of Mathematics and Physics (MIT), 18(1–4):51–151, April 1939. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193918151.

Gewertz:1933:SFF

[Gew33] Charles Mason Gewertz. Synthesis of a finite, four-terminal network from its prescribed driving-point functions and transfer function. Journal of Mathematics and Physics (MIT), 12(1–4):1–257, April 1933. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19331211;
https://www.proquest.com/pqdtglobal/docview/301789570/B03D93D9E3C14B46PQ.

Giese:1951:SFT

[Gie51] J. H. Giese. Stream functions for three-dimensional flows. Journal of Mathematics and Physics (MIT), 30(1–4):31–35, April 1951. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195130131.

Giese:1958:TEN

[Gie58] John H. Giese. On the truncation error in a numerical solution of the Neumann problem for a rectangle. Journal of Mathematics and Physics (MIT), 37(1–4):169–177, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1958371169.

Gillespie:1925:RHE

[Gil25] Louis J. Gillespie. The representation of the Haber equilibrium data by an equation. Journal of Mathematics and Physics (MIT), 4(1–4):84–96, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19254184.

Gilbarg:1947:FPC

[Gil47] D. Gilbarg. On the flow patterns common to certain classes of plane fluid motions. Journal of Mathematics and Physics (MIT), 26(1–4): 137–142, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261137.

Gilbert:1954:LTP

[Gil54] E. N. Gilbert. Lattice theoretic properties of frontal switching functions. Journal of Mathematics and Physics (MIT), 33(1–4):57–67, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195433157.

Gill:1960:CTP

[Gil60] Arthur Gill. On a class of transportation problems. Journal of Mathematics and Physics (MIT), 39(1–4):76–82, April 1960. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196039176.

Gill:1968:RNS

[Gil68] A. E. Gill. The Reynolds number similarity argument. Journal of Mathematics and Physics (MIT), 47(1–4):437–441, April 1968. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1968471437.

Graef:1938:POE

[GK38] C. Graef and S. Kusaka. On periodic orbits in the equatorial plane of a magnetic dipole. Journal of Mathematics and Physics (MIT), 17(1–4): 43–54, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193817143.

Glasser:1964:NIA

[Gla64] M. L. Glasser. A note on the integral with an application to schlomilch series. Journal of Mathematics and Physics (MIT), 43(1–4):158–162, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431158.

Godart:1941:SCP

[God41] O. Godart. On space closure of periodic orbits in the field of a magnetic dipole. Journal of Mathematics and Physics (MIT), 20(1–4):207– 217, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201207.

Goheen:1949:MDC

[Goh49] Harry E. Goheen. A method for determining certain critical masses. Journal of Mathematics and Physics (MIT), 28(1–4):107–116, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281107.

Goldberg:1946:TDA

[Gol46] Michael Goldberg. A three-dimensional analog of a plane Kempe linkage. Journal of Mathematics and Physics (MIT), 25(1–4):96–110, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194625196.

Goldberg:1947:TL

[Gol47] Michael Goldberg. Tubular linkages. Journal of Mathematics and Physics (MIT), 26(1–4):10–21, April 1947. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm194726110.

Goldberg:1951:RSP

[Gol51] Michael Goldberg. Rotors in spherical polygons. Journal of Mathematics and Physics (MIT), 30(1–4):235–244, April 1951. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1951301235.

Goldberg:1955:BRS

[Gol55] Michael Goldberg. Basic rotors in spherical polygons. Journal of Mathematics and Physics (MIT), 34(1–4):322–327, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341322.

Gold:1957:IBF

[Gol57] Louis Gold. Inverse Bessel functions: Solution for the zeros. Journal of Mathematics and Physics (MIT), 36(1–4):167–171, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1957361167.

Goldberg:1958:RTF

[Gol58] Michael Goldberg. Rotors tangent to n fixed circles. Journal of Mathematics and Physics (MIT), 37(1–4):69–74, April 1958. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195837169.

Goldberg:1959:IR

[Gol59] Michael Goldberg. Intermittent rotors. Journal of Mathematics and Physics (MIT), 38(1–4):135–140, April 1959. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1959381135.

Goldberg:1962:RTP

[Gol62] J. L. Goldberg. Richards’ transformation and positive real functions. Journal of Mathematics and Physics (MIT), 41(1–4):191–200, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411191.

Goldberg:1967:RTP

[Gol67] J. L. Goldberg. Richards’ transformation and positive-real functions: II. Journal of Mathematics and Physics (MIT), 46(1–4):355–361, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461355.

Gonzalez:1941:GTL

[Gon41] Mario O. González. A general type of Lie’s differential equations. Journal of Mathematics and Physics (MIT), 20(1–4):80–88, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194120180.

Gould:1928:NMI

[Gou28] King E. Gould. A new machine for integrating a functional product. Journal of Mathematics and Physics (MIT), 7(1–4):305–316, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192871305.

Gould:1958:GVS

[Gou58] Roderick Gould. Graphs and vector spaces. Journal of Mathematics and Physics (MIT), 37(1–4):193–214, April 1958. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1958371193.

Greenspan:1962:SES

[GP62] H. P. Greenspan and L. A. Peletier. Some exact solutions of magnetohydrodynamic viscous flow problems. Journal of Mathematics and Physics (MIT), 41(1–4):116–131, April 1962. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1962411116.

Glantz:1955:FSE

[GR55] Herbert Glantz and Eric Reissner. On finite sum equations for boundary value problems of partial difference equations. Journal of Mathematics and Physics (MIT), 34(1–4):286–297, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341286.

Greitzer:1940:CTP

[Gre40] S. L. Greitzer. Combinatorial topology of polyhedra. Journal of Mathematics and Physics (MIT), 19(1–4):260–288, April 1940. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1940191260.

Greenberg:1948:DUL

[Gre48] H. J. Greenberg. The determination of upper and lower bounds for the solution of the Dirichlet problem. Journal of Mathematics and Physics (MIT), 27(1–4):161–182, April 1948. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1948271161.

Griffith:1938:STD

[Gri38] B. A. Griffith. On the steady two-dimensional motion of a viscous liquid past a fixed circular cylinder. Journal of Mathematics and Physics (MIT), 17(1–4):5–42, April 1938. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19381715.

Gribben:1961:DODa

[Gri61a] R. J. Gribben. The damping of an oscillating disk in helium II: I. Behavior at small and large amplitudes. Journal of Mathematics and Physics (MIT), 40(1–4):177–188, April 1961. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1961401177.

Gribben:1961:DODb

[Gri61b] R. J. Gribben. The damping of an oscillating disk in helium II: II. Behavior for general amplitudes based on approximate solutions. Journal of Mathematics and Physics (MIT), 40(1–4):189–204, April 1961.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1961401189.

Gross:1962:DIS

[Gro62] Kenneth I. Gross. Discussion of an iterative solution to an equation of radiative transfer. Journal of Mathematics and Physics (MIT), 41(1–4): 53–61, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196241153.

Gilbarg:1950:FBJ

[GS50] David Gilbarg and James Serrin. Free boundaries and jets in the theory of cavitation. Journal of Mathematics and Physics (MIT), 29(1–4): 1–12, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19502911.

Green:1958:SCC

[GS58] A. E. Green and A. J. M. Spencer. The stability of a circular cylinder under finite extension and torsion. Journal of Mathematics and Physics (MIT), 37(1–4):316–338, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1958371316.

Guest:1953:NFP

[Gue53] P. G. Guest. Note on the fitting of polynomials to equally-spaced observations. Journal of Mathematics and Physics (MIT), 32(1–4): 68–71, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195332168.

Guillemin:1932:RCD

[Gui32] E. A. Guillemin. A recent contribution to the design of electric filter networks. Journal of Mathematics and Physics (MIT), 11(1–4):150– 211, April 1932. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1932111150.

Guillemin:1949:SRN

[Gui49] E. A. Guillemin. Synthesis of RC-networks. Journal of Mathematics and Physics (MIT), 28(1–4):22–42, April 1949. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194928122.

Guillemin:1960:NCT

[Gui60] E. A. Guillemin. The normal coordinate transformation of a linear system with an arbitrary loss function. Journal of Mathematics and Physics (MIT), 39(1–4):97–104, April 1960. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196039197.

Gupta:1959:DCM

[Gup59] H. C. Gupta. Diffusion by continuous movements. Journal of Mathematics and Physics (MIT), 38(1–4):36–41, April 1959. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195938136.

Gupta:1960:DPB

[Gup60] H. C. Gupta. Diffusion in the presence of barriers. Journal of Mathematics and Physics (MIT), 39(1–4):58–63, April 1960. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196039158.

Greenspan:1965:NLS

[GW65] H. P. Greenspan and S. Weinbaum. On non-linear spin-up of a rotating fluid. Journal of Mathematics and Physics (MIT), 44(1–4):66– 85, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196544166.

Gere:1943:CCC

[GZ43] Brewster H. Gere and David Zupnik. On the construction of curves of constant width. Journal of Mathematics and Physics (MIT), 22(1–4): 31–36, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194322131.

Heymans:1923:PCC

[HA23] Paul Heymans and W. P. Allis. The photoelastic constants of celluloid, glass and fused quartz. Journal of Mathematics and Physics (MIT), 2(1–4):216–233, October 1923. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192321216.

Haldane:1935:SPM

[Hal35] J. B. S. Haldane. Some problems of mathematical biology. Journal of Mathematics and Physics (MIT), 14(1–4):125–136, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141125.

Hammersley:1952:LIC

[Ham52] J. M. Hammersley. Lagrangian integration coefficients for distance functions taken over right circular cylinders. Journal of Mathematics and Physics (MIT), 31(1–4):139–150, April 1952. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1952311139.

Hanin:1957:MRR

[Han57a] Meir Hanin. Mathematical results related to a problem in wave propagation. Journal of Mathematics and Physics (MIT), 36(1–4):250–260, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361250.

Hanin:1957:PAW

[Han57b] Meir Hanin. Propagation of an aperiodic wave in a compressible viscous medium. Journal of Mathematics and Physics (MIT), 36(1–4): 234–249, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361234.

Hansen:1962:SDI

[Han62] Erik Bent Hansen. Scalar diffraction by an infinite strip and a circular disc. Journal of Mathematics and Physics (MIT), 41(1–4):229– 245, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411220.

Harmuth:1957:SSK

[Har57] Henning F. Harmuth. On the solution of the Schroedinger and the Klein–Gordon equations by digital computers. Journal of Mathematics and Physics (MIT), 36(1–4):269–278, April 1957. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1957361269.

Harary:1959:GTM

[Har59a] Frank Harary. A graph theoretic method for the complete reduction of a matrix with a view toward finding its eigenvalues. Journal of Mathematics and Physics (MIT), 38(1–4):104–111, April 1959. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1959381104.

Harary:1959:LGT

[Har59b] Frank Harary. On the line-group of two-terminal series-parallel networks. Journal of Mathematics and Physics (MIT), 38(1–4):112–118, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381112.

Haseltine:1939:MIP

[Has39] W. R. Haseltine. The mutual interaction of plasma electrons. Journal of Mathematics and Physics (MIT), 18(1–4):174–201, April 1939. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1939181174.

Hayes:1955:NSS

[Hay55] Wallace D. Hayes. Neo–Schlömilch series. Journal of Mathematics and Physics (MIT), 34(1–4):129–132, April 1955. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1955341129.

Hoff:1953:DAB

[HB53] N. J. Hoff and Victor G. Bruce. Dynamic analysis of the buckling of laterally loaded flat arches. Journal of Mathematics and Physics (MIT), 32(1–4):276–288, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321276.

Horvay:1954:TSE

[HB54] G. Horvay and J. S. Born. Tables of self-equilibrating functions. Journal of Mathematics and Physics (MIT), 33(1–4):360–373, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331360.

Hildebrand:1941:LSP

[HC41] F. B. Hildebrand and P. D. Crout. A least square procedure for solving integral equations by polynomial approximation. Journal of Mathematics and Physics (MIT), 20(1–4):310–335, April 1941. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1941201310.

Howard:1964:IPF

[HD64] Louis N. Howard and Philip G. Drazin. On instability of parallel flow of inviscid fluid in a rotating system with variable Coriolis parameter. Journal of Mathematics and Physics (MIT), 43(1–4):83–99, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196443183.

Hart:1964:NLB

[HE64] V. G. Hart and D. J. Evans. Non-linear bending of an annular plate by transverse edge forces. Journal of Mathematics and Physics (MIT), 43(1–4):275–303, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431275.

Hector:1967:LMH

[Hec67] D. L. Hector. On the linearized m. h. d. flow past a semi-infinite flat plate in the presence of a transverse magnetic field. Journal of Mathematics and Physics (MIT), 46(1–4):408–424, April 1967. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1967461408.

Hedeman:1941:CII

[Hed41] Walter R. Hedeman, Jr. The cinema integraph in interreflection problems. Journal of Mathematics and Physics (MIT), 20(1–4):402–417, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201402.

Heins:1935:AFT

[Hei35] Albert E. Heins. Applications of the Fourier Transform Theorem. Journal of Mathematics and Physics (MIT), 14(1–4):137–142, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141137.

Heins:1940:SLD

[Hei40] Albert E. Heins. On the solution of linear difference differential equations. Journal of Mathematics and Physics (MIT), 19(1–4):153–157, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1940191153.

Heins:1951:SRC

[Hei51] Albert E. Heins. Some remarks on the coupling of two ducts. Journal of Mathematics and Physics (MIT), 30(1–4):164–169, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1951301164.

Helliwell:1961:FPH

[Hel61] J. B. Helliwell. A flow pattern at high subsonic speeds past a wedge at incidence in a free stream and a choked channel. Journal of Mathematics and Physics (MIT), 40(1–4):1–22, April 1961. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19614011.

Henrici:1957:RCI

[Hen57] Peter Henrici. On the representation of a certain integral involving Bessel functions by hypergeometric series. Journal of Mathematics and Physics (MIT), 36(1–4):151–156, April 1957. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1957361151.

Hersch:1964:MOD

[Her64] Joseph Hersch. On the methods of one-dimensional auxiliary problems and of domain partitioning: Their application to lower bounds for the eigenvalues of Schrödinger’s equation. Journal of Mathematics and Physics (MIT), 43(1–4):15–26, April 1964. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196443115.

Heuman:1941:CTC

[Heu41a] Carl Heuman. Corrections in “Tables of Complete Elliptic Integrals”. Journal of Mathematics and Physics (MIT), 20(1–4):336, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1941201336. See [Heu41b].

Heuman:1941:TCE

[Heu41b] Carl Heuman. Tables of complete elliptic integrals. Journal of Mathematics and Physics (MIT), 20(1–4):127–206, April 1941. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1941201127. See corrections [Heu41a].

Heymans:1924:NPR

[Hey24a] Paul Heymans. Note on a property of rectilinear lines of principal stress. Journal of Mathematics and Physics (MIT), 3(3):182–185, April 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192433182.

Heymans:1924:DSP

[Hey24b] Paul Heymans. On dynamic stresses in pseudo-continuous media. Journal of Mathematics and Physics (MIT), 3(4):237–252, May 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192434237.

Heins:1947:CTA

[HF47] A. E. Heins and H. Feshbach. The coupling of two acoustical ducts. Journal of Mathematics and Physics (MIT), 26(1–4):143–155, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261143.

Huang:1968:NTM

[HG68] A. Ben Huang and Don P. Giddens. A new table for a modified (half range) Gauss–Hermite quadrature with an evaluation of the integral. Journal of Mathematics and Physics (MIT), 47(1–4):213–218, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471213.

Heymans:1925:TPC

[HH25] Paul Heymans and W. J. Heymans. The torsion problem of curved beams. Journal of Mathematics and Physics (MIT), 4(1–4):33–61, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19254133.

Higgins:1942:ITC

[Hig42] Thomas James Higgins. The inductance of tabular conductors of eccentric-annular cross-section. Journal of Mathematics and Physics (MIT), 21(1–4):159–177, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1942211159.

Hilderband:1942:NID

[Hil42] F. B. Hilderband. Note on the integro-differential equation of a problem in the theory of plane stress. Journal of Mathematics and Physics (MIT), 21(1–4):19–22, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194221119.

Hildebrand:1943:SDC

[Hil43] F. B. Hildebrand. On the stress distribution in cantilever beams. Journal of Mathematics and Physics (MIT), 22(1–4):188–203, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1943221188.

Hildebrand:1952:CNS

[Hil52] F. B. Hildebrand. On the convergence of numerical solutions of the heat-flow equation. Journal of Mathematics and Physics (MIT), 31(1– 4):35–41, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195231135.

Hildebrand:1953:NLM

[Hil53] F. B. Hildebrand. Note on S. N. Lin’s method of factoring polynomials. Journal of Mathematics and Physics (MIT), 32(1–4):164–170, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321164.

Hildebrand:1960:EBN

[Hil60] F. B. Hildebrand. On error bounds for numerical differentiation. Journal of Mathematics and Physics (MIT), 39(1–4):151–154, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1960391151.

Hitchcock:1923:MNS

[Hit23] Frank L. Hitchcock. A method for the numerical solution of integral equations. Journal of Mathematics and Physics (MIT), 2(1–4):88–104, October 1923. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19232188.

Hitchcock:1924:CPT

[Hit24] Frank Lauren Hitchcock. The coincident points of two algebraic transformations. Journal of Mathematics and Physics (MIT), 3(2):66–71, March 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19243266.

Hitchcock:1925:NMT

[Hit25a] Frank L. Hitchcock. A new method in the theory of quantics. Journal of Mathematics and Physics (MIT), 4(1–4):238–256, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192541238.

Hitchcock:1925:TOD

[Hit25b] Frank L. Hitchcock. A theory of ordered determinants with application to polyadics. Journal of Mathematics and Physics (MIT), 4(1–4):205– 237, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192541205.

Hitchcock:1927:ETP

[Hit27] Frank L. Hitchcock. The expression of a tensor or a polyadic as a sum of products. Journal of Mathematics and Physics (MIT), 6(1–4):

164–189, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761164.

Hitchcock:1928:MIG

[Hit28] Frank L. Hitchcock. Multiple invariants and generalized rank of a P - way matrix or tensor. Journal of Mathematics and Physics (MIT), 7(1– 4):39–79, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19287139.

Hitchcock:1929:NMT

[Hit29] Frank L. Hitchcock. A new method in the theory of quantics. second paper. Journal of Mathematics and Physics (MIT), 8(1–4):81– 105, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19298181.

Hitchcock:1930:ARE

[Hit30] Frank L. Hitchcock. An analysis of rotations in Euclidean four-space by sedenions. Journal of Mathematics and Physics (MIT), 9(1–4):188– 193, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193091188.

Hitchcock:1935:PDF

[Hit35] Frank L. Hitchcock. On S, P, D and F biharmonic polynomials. Journal of Mathematics and Physics (MIT), 14(1–4):143–178, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141143.

Hitchcock:1938:FCR

[Hit38] Frank L. Hitchcock. Finding complex roots of algebraic equations. Journal of Mathematics and Physics (MIT), 17(1–4):55–58, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193817155.

Hitchcock:1939:AEC

[Hit39] Frank L. Hitchcock. Algebraic equations with complex coefficients. Journal of Mathematics and Physics (MIT), 18(1–4):202–210, April 1939. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1939181202.

Hitchcock:1941:DPS

[Hit41] Frank L. Hitchcock. The distribution of a product from several sources to numerous localities. Journal of Mathematics and Physics (MIT), 20(1–4):224–230, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1941201224.

Hitchcock:1944:IGC

[Hit44] Frank L. Hitchcock. An improvement on the G C D method for complex roots. Journal of Mathematics and Physics (MIT), 23(1–4):69– 74, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194423169.

Hoffmann:1936:GIM

[HKP36] B. Hoffmann, Winston E. Kock, and M. H. L. Pryce. A geometrical interpretation of the method of symmetrical components. Journal of Mathematics and Physics (MIT), 15(1–4):187–190, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1936151187.

Hodge:1950:YCP

[Hod50] P. G. Hodge, Jr. Yield conditions in plane plastic stress. Journal of Mathematics and Physics (MIT), 29(1–4):38–48, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195029138.

Hodge:1952:BMC

[Hod52] P. G. Hodge, Jr. The Brownian motion of coupled systems. Journal of Mathematics and Physics (MIT), 31(1–4):120–130, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1952311120.

Hodge:1957:PBA

[Hod57] P. G. Hodge, Jr. Plastic bending of an annular plate. Journal of Mathematics and Physics (MIT), 36(1–4):130–137, April 1957. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1957361130.

Hoffmann:1934:MLC

[Hof34] Banesh Hoffmann. A modification of Levi-Civita’s wave equation. Journal of Mathematics and Physics (MIT), 13(1–4):268–274, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1934131268.

Hoffmann:1946:TEC

[Hof46] Banesh Hoffmann. Tensors and equivalent circuits. Journal of Mathematics and Physics (MIT), 25(1–4):21–25, April 1946. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194625121.

Hopf:1934:CSP

[Hop34] Eberhard Hopf. On causality, statistics and probability. Journal of Mathematics and Physics (MIT), 13(1–4):51–102, April 1934. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193413151.

Hopf:1935:RCP

[Hop35] E. Hopf. Remarks on causality and probability. Journal of Mathematics and Physics (MIT), 14(1–4):4–9, April 1935. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19351414.

Horton:1950:ESL

[Hor50a] C. W. Horton. On the extension of some Lommel integrals to Struve functions with an application to acoustic radiation. Journal of Mathematics and Physics (MIT), 29(1–4):31–37, April 1950. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195029131.

Horton:1950:STS

[Hor50b] C. W. Horton. A short table of Struve functions and of some integrals involving Bessel and Struve functions. Journal of Mathematics and Physics (MIT), 29(1–4):56–58, April 1950. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195029156.

Hostetter:1936:ETI

[Hos36a] I. M. Hostetter. Existence theorem on implicit vector functions. Journal of Mathematics and Physics (MIT), 15(1–4):97–104, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193615197.

Hostetter:1936:EGV

[Hos36b] I. M. Hostetter. An extension of Gibbs’ vector analysis to Nspace. Journal of Mathematics and Physics (MIT), 15(1–4):191–204, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151191.

Hostetter:1943:PP

[Hos43] I. M. Hostetter. Polyadic products. Journal of Mathematics and Physics (MIT), 22(1–4):93–114, April 1943. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm194322193.

Hovgaard:1923:NPT

[Hov23] William Hovgaard. A new proof of the theory of ordinary bending and its extension to beams of non-homogeneous materials. Journal of Mathematics and Physics (MIT), 2(1–4):195–209, October 1923. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192321195.

Hovgaard:1925:DSB

[Hov25] Prof. William Hovgaard. Determination of the stresses in a beam by means of the principle of least work. Journal of Mathematics and Physics (MIT), 4(1–4):103–129, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192541103.

Hovgaard:1926:AEP

[Hov26] Professor William Hovgaard. Adjustment of the elastic properties of a model keel United States Army Airship RS-1. Journal of Mathematics and Physics (MIT), 5(1–4):8–39, January 1926. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1926518.

Hovgaard:1927:EDP

[Hov27] Prof. William Hovgaard. The elastic deformation of pipe-bends. Journal of Mathematics and Physics (MIT), 6(1–4):69–118, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19276169.

Hovgaard:1928:FRP

[Hov28a] Prof William Hovgaard. Further research on pipe bends. Journal of Mathematics and Physics (MIT), 7(1–4):239–297, April 1928. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192871239.

Hovgaard:1928:DPP

[Hov28b] Professor William Hovgaard. Deformation of plane pipes. Journal of Mathematics and Physics (MIT), 7(1–4):198–238, April 1928. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192871198.

Hovgaard:1929:THP

[Hov29] William Hovgaard. Tests on high-pressure pipe bends. Journal of Mathematics and Physics (MIT), 8(1–4):293–344, April 1929. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192981293.

Hovgaard:1932:RET

[Hov32] William Hovgaard. Ritz’s electrodynamic theory. Journal of Mathematics and Physics (MIT), 11(1–4):218–254, April 1932. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1932111218.

Hovgaard:1934:SDL

[Hov34] William Hovgaard. The stress distribution of longitudinal welds and adjoining structures. Journal of Mathematics and Physics (MIT), 13(1–4):195–248, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1934131195.

Howe:1931:LE

[How31] G. W. O. Howe. Letter to the Editor. Journal of Mathematics and Physics (MIT), 10(1–4):237, April 1931. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1931101237.

Howard:1958:HSJ

[How58] Louis N. Howard. Hydrodynamic stability of a jet. Journal of Mathematics and Physics (MIT), 37(1–4):283–298, April 1958. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1958371283.

Howard:1961:NEC

[How61] Louis N. Howard. A note on the existence of certain viscous flows. Journal of Mathematics and Physics (MIT), 40(1–4):172–176, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1961401172.

Hodge:1948:VPP

[HP48] P. Hodge and W. Prager. A variational principle for plastic materials with strain-hardening. Journal of Mathematics and Physics (MIT), 27(1–4):1–10, April 1948. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19482711.

Hooker:1960:BFE

[HP60] W. Hooker and M. H. Protter. Bounds for the first eigenvalue of a rhombic membrane. Journal of Mathematics and Physics (MIT), 39(1– 4):18–34, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196039118.

Hitchcock:1925:MCO

[HR25] Frank L. Hitchcock and Lepine Hall Rice. The multiple complement of one or more polyadics. Journal of Mathematics and Physics (MIT), 4(1–4):179–187, January 1925. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192541179.

Hanson:1967:CRS

[HR67] R. J. Hanson and D. L. Russell. Classification and reduction of second order systems at a turning point. Journal of Mathematics and Physics (MIT), 46(1–4):74–92, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746174.

Hua:1936:FTC

[HS36] Loo-Keng Hua and Shien-Siu Shu. On Fourier transforms in L^{p} in the complex domain. Journal of Mathematics and Physics (MIT), 15(1–4): 249–263, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151249.

Hillman:1948:IF

[HS48] Abraham Hillman and Herbert E. Salzer. The inverse function of z = w^{−1} tanh w and z = w^{−1} coth w. Journal of Mathematics and Physics (MIT), 27(1–4):202–209, April 1948. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1948271202.

Hartranft:1968:EPT

[HS68] R. J. Hartranft and G. C. Sih. Effect of plate thickness on the bending stress distribution around through cracks. Journal of Mathematics and Physics (MIT), 47(1–4):276–291, April 1968. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1968471276.

Hsu:1942:TTCa

[Hsu42a] Chang-Pen Hsu. Transmission theory of a cylindrical hollow tube guide. Journal of Mathematics and Physics (MIT), 21(1–4):23–42, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194221123.

Hsu:1942:TTCb

[Hsu42b] Chang-Pen Hsu. Transmission theory of concentric lines. Journal of Mathematics and Physics (MIT), 21(1–4):43–51, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194221143.

Hsu:1968:MSH

[Hsu68] Chen-Chi Hsu. A method of solution for higher approximation of boundary layer flows. Journal of Mathematics and Physics (MIT), 47(1–4):185–198, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471185.

Huang:1967:EID

[Hua67] A. Ben Huang. Evaluation of the integral by the discrete ordinary method. Journal of Mathematics and Physics (MIT), 46(1–4):107– 110, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461107.

See note [Hua68].

Huang:1968:NPE

[Hua68] A. B. Huang. A note on the paper “Evaluation of the Integral by the Discrete Ordinate Method”. Journal of Mathematics and Physics (MIT), 47(1–4):219, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471219. See [Hua67].

Hudson:1950:ENS

[Hud50] G. E. Hudson. On the explicit numerical solution of linear homogeneous differential equations with constant coefficients. Journal of Mathematics and Physics (MIT), 29(1–4):52–55, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195029152.

Hunter:1960:SFA

[Hun60] Herbert E. Hunter. On a special function arising in a solution of the Poisson equation. Journal of Mathematics and Physics (MIT), 39(1–4): 217–221, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391217.

Hunt:1961:SVW

[Hun61] J. N. Hunt. Stable vortex wakes near rigid boundaries. Journal of Mathematics and Physics (MIT), 40(1–4):33–40, April 1961. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196140133.

Hitchcock:1922:NVM

[HW22] Frank L. Hitchcock and Norbert Wiener. A new vector method in integral equations. Journal of Mathematics and Physics (MIT), 1(1): 1–20, January 1922. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1922111.

Hoel:1947:ARS

[HW47] P. G. Hoel and D. D. Wall. The accuracy of the root-squaring method for solving equations. Journal of Mathematics and Physics (MIT), 26(1–4):156–164, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1947261156.

Handelman:1954:LPI

[HW54] G. H. Handelman and W. H. Warner. Loading paths and the incremental strain law. Journal of Mathematics and Physics (MIT), 33(1–4): 157–164, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331157.

Ho:1967:BLN

[HW67] Dar-Veig Ho and Howell K. Wilson. On the boundary layer near a porous wall. Journal of Mathematics and Physics (MIT), 46(1–4):203– 210, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461203.

Hyman:1951:RSN

[Hym51] M. A. Hyman. Remarks on “A Study of the Numerical Solution of Partial Differential Equations”. Journal of Mathematics and Physics (MIT), 30(1–4):256, April 1951. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1951301256. See [OHK50].

Ingram:1944:FEN

[IC44] W. H. Ingram and C. M. Cramlet. On the foundations of electrical network theory. Journal of Mathematics and Physics (MIT), 23(1–4): 134–155, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1944231134.

Imai:1949:AWKb

[IH49] Isao Imai and Hidenori Hasimoto. Applications of the W.K.B. method to the flow of a compressible fluid, II. Journal of Mathematics and Physics (MIT), 28(1–4):205–214, April 1949. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1949281205.

Ikehara:1931:ELT

[Ike31a] Shikao Ikehara. An extension of Landau’s theorem in the analytical theory of numbers. Journal of Mathematics and Physics (MIT), 10(1– 4):1–12, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19311011.

Ikehara:1931:NIP

[Ike31b] Shikao Ikehara. A note on infinite products. Journal of Mathematics and Physics (MIT), 10(1–4):92–94, April 1931. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193110192.

Ikehara:1931:IFR

[Ike31c] Shikao Ikehara. On integral functions with real negative zeros. Journal of Mathematics and Physics (MIT), 10(1–4):84–91, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193110184.

Ikehara:1931:TTH

[Ike31d] Shikao Ikehara. On Tauberian theorems of Hardy and Littlewood and a note on Wintner’s paper. Journal of Mathematics and Physics (MIT), 10(1–4):75–83, April 1931. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193110175.

Imai:1949:AWKa

[Ima49] Isao Imai. Applications of the W.K.B. method to the flow of a compressible fluid, I. Journal of Mathematics and Physics (MIT), 28(1–4):

173–182, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281173.

Ingram:1924:NC

[Ing24] W. H. Ingram. Note on the curl. Journal of Mathematics and Physics (MIT), 3(3):186–187, April 1924. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192433186.

Ingraham:1934:RSM

[Ing34] M. H. Ingraham. On the rational solutions of the matrix equation P (X) = A. Journal of Mathematics and Physics (MIT), 13(1–4):46– 50, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193413146.

Ikenberry:1952:CEH

[IR52] E. Ikenberry and W. A. Rutledge. Convergence of expansions in the Hermite polynomials H_{n}(hw). Journal of Mathematics and Physics (MIT), 31(1–4):180–183, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1952311180.

Isaacs:1947:NDU

[Isa47] Rufus Isaacs. Numerical determination by use of special computational devices of an integral operator in the theory of compressible fluids: II. Determination of the coefficients of the integral operator by interpolatory means. Journal of Mathematics and Physics (MIT), 26(1–4): 165–181, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261165.

Infeld:1947:SSB

[ISC47] L. Infeld, V. G. Smith, and W. Z. Chien. On some series of Bessel functions. Journal of Mathematics and Physics (MIT), 26(1–4):22– 28, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194726122.

J:1930:RPG

[J30] F. W. Sohon S. J. Rotation and perversion groups in Euclidean space of four dimensions. Journal of Mathematics and Physics (MIT), 9(1–4): 194–260, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193091194.

Jaeger:1948:RIB

[Jae48] J. C. Jaeger. Repeated integrals of Bessel functions and the theory of transients in filter circuits. Journal of Mathematics and Physics (MIT), 27(1–4):210–219, April 1948. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1948271210.

Jaeger:1955:NVT

[Jae55] J. C. Jaeger. Numerical values for the temperature in radial heat flow. Journal of Mathematics and Physics (MIT), 34(1–4):316–321, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341316.

Janos:1962:EEQ

[Jan62] W. A. Janos. On error estimation in quantum statistics perturbation theory. Journal of Mathematics and Physics (MIT), 41(1–4):246– 253, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411246.

Jenssen:1960:AID

[Jen60] Osmund Jenssen. Asymptotic integration of the differential equation for a special case of symmetrically loaded toroidal shells. Journal of Mathematics and Physics (MIT), 39(1–4):1–17, April 1960. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19603911.

Jessen:1935:SAP

[Jes35] B. Jessen. Some analytical problems relating to probability. Journal of Mathematics and Physics (MIT), 14(1–4):24–27, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193514124.

Jones:1958:AEM

[JK58] Douglas S. Jones and Morris Kline. Asymptotic expansion of multiple integrals and the method of stationary phase. Journal of Mathematics and Physics (MIT), 37(1–4):1–28, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19583711.

Joffe:1927:MBD

[JKS27] A. Joffé, T. Kurchatoff, and K. Sinelnikoff. The mechanism of breakdown of dielectrics. Journal of Mathematics and Physics (MIT), 6(1–4):

133–144, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761133.

Johnson:1930:ERT

[Joh30] E. A. Johnson. The effect of reverberation time in the reproduction of sounds. Journal of Mathematics and Physics (MIT), 9(1–4):1–10, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1930911.

Johnson:1963:BLT

[Joh63] M. W. Johnson, Jr. A boundary layer theory for unsymmetric deformations of circular cylindrical elastic shells. Journal of Mathematics and Physics (MIT), 42(1–4):167–188, April 1963. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1963421167.

Jones:1956:MCT

[Jon56] E. E. Jones. The magnetostatic characteristics of two non-magnetic elliptic cylinders. Journal of Mathematics and Physics (MIT), 35(1–4): 266–277, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351266.

Jones:1964:CSIa

[Jon64a] D. S. Jones. On a certain singular integral equation I. Journal of Mathematics and Physics (MIT), 43(1–4):27–33, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196443127.

Jones:1964:CSIb

[Jon64b] D. S. Jones. On a certain singular integral equation II. Journal of Mathematics and Physics (MIT), 43(1–4):263–273, April 1964. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431263.

Jones:1968:EII

[Jon68] A. L. Jones. An extension of an inequality involving modified Bessel functions. Journal of Mathematics and Physics (MIT), 47(1–4):220– 221, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471220.

Johnson:1955:IDS

[JR55] Millard W. Johnson and Eric Reissner. On inextensional deformations of shallow elastic shells. Journal of Mathematics and Physics (MIT), 34(1–4):335–350, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341335.

Johnson:1958:FTT

[JR58] M. W. Johnson and E. Reissner. On the foundations of the theory of thin elastic shells. Journal of Mathematics and Physics (MIT), 37(1–4): 371–392, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371371.

Kagno:1935:NHC

[Kag35] I. N. Kagno. A note on the Heawood color formula. Journal of Mathematics and Physics (MIT), 14(1–4):228–231, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1935141228.

Kagno:1936:TSH

[Kag36] I. N. Kagno. The triangulation of surfaces and the Heawood color formula. Journal of Mathematics and Physics (MIT), 15(1–4):179– 186, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151179.

Kagno:1937:MGS

[Kag37] I. N. Kagno. The mapping of graphs on surfaces. Journal of Mathematics and Physics (MIT), 16(1–4):46–75, April 1937. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193716146.

Kagno:1938:PSS

[Kag38] I. N. Kagno. Perfect subdivision of surfaces. Journal of Mathematics and Physics (MIT), 17(1–4):76–111, April 1938. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193817176.

Kagno:1940:C

[Kag40] I. N. Kagno. Correction. Journal of Mathematics and Physics (MIT), 19(1–4):158, April 1940. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1940191158.

Kagno:1941:CNS

[Kag41] I. N. Kagno. On a certain non-separating graph on an orientable surface. Journal of Mathematics and Physics (MIT), 20(1–4):370–387, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201370.

Kales:1935:CS

[Kal35a] Morris L. Kales. On the closure of the set {e^{inθ}}. Journal of Mathematics and Physics (MIT), 14(1–4):187–188, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1935141187.

Kales:1935:SNL

[Kal35b] Morris L. Kales. A simple non-linear integral equation. Journal of Mathematics and Physics (MIT), 14(1–4):189–190, April 1935. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1935141189.

Kane:1962:PRW

[Kan62] Julius Kane. The propagation of Rayleigh waves past a fluid-loaded boundary. Journal of Mathematics and Physics (MIT), 41(1–4):179– 190, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411179.

Kane:1964:NQS

[Kan64] J. Kane. Note on the quadrature of some exponential transforms of cylindrical functions. Journal of Mathematics and Physics (MIT), 43(1–4):140–143, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431140.

Kanwal:1965:DDI

[Kan65] R. P. Kanwal. Dynamic displacements in an infinite elastic space and matched asymptotic expansions. Journal of Mathematics and Physics (MIT), 44(1–4):275–283, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1965441275.

Kanwal:1968:EDD

[Kan68] R. P. Kanwal. Errata: Dynamical displacements in an infinite elastic space and matched asymptotic expansions. Journal of Mathematics and Physics (MIT), 47(1–4):222, April 1968. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1968471222.

Kaplan:1934:VRP

[Kap34] N. Kaplan. V_{m} in R_{m+1} possessing an M-tuply orthogonal net along the lines of curvature. Journal of Mathematics and Physics (MIT), 13(1–4):426–432, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1934131426.

Kaplan:1936:CTD

[Kap36] Nathan Kaplan. On certain three-dimensional manifolds in Euclidean sixspace (V_{3} in R_{6}). Journal of Mathematics and Physics (MIT), 15(1–4):219–248, April 1936. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1936151219.

Kaplan:1946:ATC

[Kap46] E. L. Kaplan. Auxiliary table of complete elliptic integrals. Journal of Mathematics and Physics (MIT), 25(1–4):26–36, April 1946. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194625126.

Kaplan:1948:ATI

[Kap48] E. L. Kaplan. Auxiliary table for the incomplete elliptic integrals. Journal of Mathematics and Physics (MIT), 27(1–4):11–36, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194827111.

Kaplan:1950:MEI

[Kap50] E. L. Kaplan. Multiple elliptic integrals. Journal of Mathematics and Physics (MIT), 29(1–4):69–75, April 1950. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195029169.

Kaplan:1952:NIN

[Kap52] E. L. Kaplan. Numerical integration near a singularity. Journal of Mathematics and Physics (MIT), 31(1–4):1–28, April 1952. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19523111.

Karlin:1943:CSE

[Kar43] Meyer Karlin. Characterization of the system of extremals of variation problem of higher order in the plane. Journal of Mathematics and Physics (MIT), 22(1–4):158–180, April 1943. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1943221158.

Karlin:1949:NEC

[Kar49] Meyer Karlin. Note on the expansion of confluent hypergeometric functions in terms of Bessel functions of integral order. Journal of Mathematics and Physics (MIT), 28(1–4):43–44, April 1949. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194928143.

Kaye:1955:TFE

[Kay55] Joseph Kaye. A table of the first eleven repeated integrals of the error function. Journal of Mathematics and Physics (MIT), 34(1–4):119– 125, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341119.

Kozin:1963:CBL

[KB63] F. Kozin and J. L. Bogdanoff. A comment on “The Behavior of Linear Systems with Random Parametric Excitation”. Journal of Mathematics and Physics (MIT), 42(1–4):336–337, April 1963. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1963421336. See [CD62].

Kurtz:1962:CAA

[KC62] Edward F. Kurtz and Stephen H. Crandall. Computer-aided analysis of hydrodynamic stability. Journal of Mathematics and Physics (MIT), 41(1–4):264–279, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411264.

Kopal:1951:NFR

[KCK51] Zdeněk Kopal, Pierre Carrus, and Katherine E. Kavanagh. A new formula for repeated mechanical quadratures. Journal of Mathematics and Physics (MIT), 30(1–4):44–48, April 1951. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195130144.

Kasner:1948:GAT

[KD48] Edward Kasner and John De Cicco. Generalization of Appell’s transformation. Journal of Mathematics and Physics (MIT), 27(1–4):262– 269, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271262.

Keilson:1962:NSM

[Kei62a] Julian Keilson. Non-stationary Markov walks on the lattice. Journal of Mathematics and Physics (MIT), 41(1–4):205–211, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1962411205.

Keilson:1962:UGF

[Kei62b] Julian Keilson. The use of Green’s functions in the study of bounded random walks with application to queuing theory. Journal of Mathematics and Physics (MIT), 41(1–4):42–52, April 1962. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196241142.

Keith:1963:GRA

[Kei63] James Clark Keith. Gravitational radiation and aberrated centripetal force reactions in G. D. Birkhoff relativity theory, Part I: Part One. Retarded gravitational interaction forces. Journal of Mathematics and Physics (MIT), 42(1–4):248–262, April 1963. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1963421248.

Kelman:1967:STC

[Kel67] Robert B. Kelman. The steady temperature in cylinders with mixed Dirichlet and radiation conditions on the lateral wall. Journal of Mathematics and Physics (MIT), 46(1–4):333–342, April 1967. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1967461333.

Keyes:1922:CVG

[Key22] Frederick G. Keyes. The constant volume gas thermometer. Journal of Mathematics and Physics (MIT), 1(2):89–102, March 1922. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19221289.

Kuerti:1956:ODS

[KF56] G. Kuerti and K. Faymon. One-dimensional shock-motion of a gas with γ = 3, enclosed between fixed walls. Journal of Mathematics and Physics (MIT), 35(1–4):384–393, April 1956. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1956351384.

King:1934:ICP

[Kin34] G. W. King. The indeterminate and composite product of matrices. Journal of Mathematics and Physics (MIT), 13(1–4):433–454, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1934131433.

King:1965:ANL

[Kin65] J. P. King. An application of a non-linear transform to infinite products. Journal of Mathematics and Physics (MIT), 44(1–4):408–409, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441408.

Kirkham:1957:GFZ

[Kir57] Don Kirkham. Graphs and formulas for zeros of cross product Bessel functions. Journal of Mathematics and Physics (MIT), 36(1–4):371– 377, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361371.

Kline:1948:SBE

[Kli48] Morris Kline. Some Bessel equations and their applications to guide and cavity theory. Journal of Mathematics and Physics (MIT), 27(1–4): 37–48, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194827137.

Kotik:1962:CTA

[KM62] J. Kotik and V. Mangulis. Comparison of two approximate dipole distributions for a lenticular cylinder in a semi-infinite fluid. Journal of Mathematics and Physics (MIT), 41(1–4):280–290, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1962411280.

Knobel:1926:PEH

[KN26] Max Knobel and Ralph B. Norton. The penetration of electrolytic hydrogen through iron. Journal of Mathematics and Physics (MIT), 5(1–4):75–83, January 1926. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19265175.

Knobel:1926:TO

[Kno26] Max Knobel. The theory of overvoltage. Journal of Mathematics and Physics (MIT), 5(1–4):65–74, January 1926. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19265165.

Kogbetliantz:1935:MAS

[Kog35] Ervand Kogbetliantz. Sur les moyennes arithmétiques des séries — noyaux des développements en séries d’Hermite et de Laguerre et sur celles de ces séries — noyaux derivées terme a terme. (French) [On the arithmetic means of series — kernels of developments in the series of Hermite and of Laguerre and on those of these series — kernels derived term by term]. Journal of Mathematics and Physics (MIT), 14(1–4): 37–99, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193514137.

Kopal:1948:TCH

[Kop48] Zdeněk Kopal. A table of the coefficients of the Hermite quadrature formula. Journal of Mathematics and Physics (MIT), 27(1–4):259– 261, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271259.

Kotorynski:1968:SCP

[Kot68] W. P. Kotorynski. Self-convection in plane flow due to a source singularity of heat. Journal of Mathematics and Physics (MIT), 47(1–4): 170–184, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471170.

Kozin:1959:QD

[Koz59] Frank Kozin. On the quadratic detector — I. Journal of Mathematics and Physics (MIT), 38(1–4):119–125, April 1959. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1959381119.

Kozin:1963:ASS

[Koz63] Frank Kozin. On almost sure stability of linear systems with random coefficients. Journal of Mathematics and Physics (MIT), 42(1–4):59– 67, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196342159.

Katz:1953:SNL

[KP53] S. Katz and A. M. Peiser. A system of non-linear partial differential equations arising in heat transfer. Journal of Mathematics and Physics (MIT), 32(1–4):256–268, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321256.

Kuo:1967:STS

[KP67] Nan-Hung Kuo and Martin A. Plonus. A systematic technique in the solution of diffraction by a right-angled dielectric wedge. Journal of Mathematics and Physics (MIT), 46(1–4):394–407, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1967461394.

Kruse:1951:I

[KR51] U. E. Kruse and N. F. Ramsey. The integral. Journal of Mathematics and Physics (MIT), 30(1–4):40–43, April 1951. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195130140.

Knowles:1956:DES

[KR56] James K. Knowles and Eric Reissner. A derivation of the equations of shell theory for general orthogonal coordinates. Journal of Mathematics and Physics (MIT), 35(1–4):351–358, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1956351351.

Knowles:1958:NSS

[KR58] James K. Knowles and Eric Reissner. Note on stress-strain relations for thin elastic shells. Journal of Mathematics and Physics (MIT), 37(1–4): 269–282, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371269.

Kelly:1960:RVV

[KR60] Edward J. Kelly and William L. Root. A representation of vectorvalued random processes. Journal of Mathematics and Physics (MIT), 39(1–4):211–216, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391211.

Kramer:1959:GST

[Kra59] H. P. Kramer. A generalized sampling theorem. Journal of Mathematics and Physics (MIT), 38(1–4):68–72, April 1959. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195938168.

Kretzmer:1950:AAC

[Kre50] Ernest R. Kretzmer. An application of auto-correlation analysis. Journal of Mathematics and Physics (MIT), 29(1–4):179–190, April 1950.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291179.

Kron:1934:NRD

[Kro34] Gabriel Kron. Non-Riemannian dynamics of rotating electrical machinery. Journal of Mathematics and Physics (MIT), 13(1–4):103–194, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1934131103.

Kroner:1963:DNC

[Krö63] Ekkehart Kröner. Dislocation: a new concept in the continuum theory of plasticity. Journal of Mathematics and Physics (MIT), 42(1–4): 27–37, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196342127.

Kornhauser:1952:VTA

[KS52] E. T. Kornhauser and I. Stakgold. A variational theorem for ∇^{2}u + λu = 0 and its application. Journal of Mathematics and Physics (MIT), 31(1–4):45–54, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195231145.

Keyes:1922:ESM

[KSJ22] Frederick G. Keyes, Leighton B. Smith, and David B. Joubert. The equation of state for methane gas phase. Journal of Mathematics and Physics (MIT), 1(4):191–210, August 1922. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192214191.

Keyes:1922:TPM

[KTS22] Frederick G. Keyes, Robert S. Taylor, and Leighton B. Smith. The thermodynamic properties of methane. Journal of Mathematics and Physics (MIT), 1(4):211–242, August 1922. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192214211.

Keyes:1922:EAT

[KTY22] Frederick G. Keyes, Bailey Townshend, and Louis H. Young. The establishing of the absolute temperature scale below the melting point of ice. Journal of Mathematics and Physics (MIT), 1(4):243–312, August 1922. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192214243.

Ku:1926:NME

[Ku26] Y. H. Ku. Note on a method of evaluating the complex roots of a quartic equation. Journal of Mathematics and Physics (MIT), 5(1–4):126–128, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192651126.

Kuerti:1951:BLC

[Kue51] G. Kuerti. Boundary layer in convergent flow between spiral walls. Journal of Mathematics and Physics (MIT), 30(1–4):106–115, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1951301106.

Kuehl:1960:CIL

[Kue60] Hans H. Kuehl. Current on an infinitely long cylindrical antenna. Journal of Mathematics and Physics (MIT), 39(1–4):121–125, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1960391121.

Kuehl:1964:KTW

[Kue64] Hans H. Kuehl. The kinetic theory of waves in a warm plasma excited by a current source. Journal of Mathematics and Physics (MIT), 43(1–4):218–226, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431218.

Kuehl:1966:EWP

[Kue66] H. H. Kuehl. Excitation of a warm plasma by an electric dipole. Journal of Mathematics and Physics (MIT), 45(1–4):318–321, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451318.

Kuhn:1957:NPT

[Kuh57] H. W. Kuhn. A note on Prager’s transportation problem. Journal of Mathematics and Physics (MIT), 36(1–4):107–111, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1957361107.

Kuo:1943:FCV

[Kuo43] Y. H. Kuo. The flow of a compressible viscous fluid through a straight pipe. Journal of Mathematics and Physics (MIT), 22(1–4):13–30, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194322113.

Kuo:1953:FIV

[Kuo53] Y. H. Kuo. On the flow of an incompressible viscous fluid past a flat plate at moderate Reynolds numbers. Journal of Mathematics and Physics (MIT), 32(1–4):83–101, April 1953. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195332183.

Kuscer:1955:MPA

[Kuš55] I. Kuščer. Milne’s problem for anisotropic scattering. Journal of Mathematics and Physics (MIT), 34(1–4):256–266, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341256.

Kuscer:1958:DRL

[Kuš58] I. Kuščer. Diffuse reflection of light from a semiinfinite scattering medium. Journal of Mathematics and Physics (MIT), 37(1–4):52– 57, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195837152.

Knowles:1960:BEP

[KW60] James K. Knowles and Neng-Ming Wang. On the bending of an elastic plate containing a crack. Journal of Mathematics and Physics (MIT), 39(1–4):223–236, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391223.

Lowan:1943:TI

[LA43] Arnold N. Lowan and Milton Abramowitz. Tables of the integrals of and. Journal of Mathematics and Physics (MIT), 22(1–4):2–12, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19432212.

Lam:1968:SEC

[Lam68] John Lam. Scattering by an expanding conducting sphere. Journal of Mathematics and Physics (MIT), 47(1–4):199–212, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1968471199.

Lanczos:1938:TIE

[Lan38] C Lanczos. Trigonometric interpolation of empirical and analytical functions. Journal of Mathematics and Physics (MIT), 17(1–4):123– 199, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1938171123.

Langenhop:1951:NLE

[Lan51] C. E. Langenhop. Note on Levinson’s existence theorem for forced periodic solutions of a second order differential equation. Journal of Mathematics and Physics (MIT), 30(1–4):36–39, April 1951. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195130136.

Langenhop:1959:NAP

[Lan59] C. E. Langenhop. Note on almost periodic solutions of non-linear differential equations. Journal of Mathematics and Physics (MIT), 38(1–4): 126–129, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381126.

Lardner:1966:SDS

[Lar66] Thomas J. Lardner. On the stress distribution in a shallow logarithmic shell of revolution. Journal of Mathematics and Physics (MIT), 45(1– 4):23–34, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196645123.

Latta:1963:SDE

[Lat63] G. E. Latta. Some differential equations of the Mathieu type, and related integral equations. Journal of Mathematics and Physics (MIT), 42(1–4):139–146, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421139.

Lawrence:1934:TLM

[Law34] Lawrence Lawrence. The theory of linear matrix transformations with applications to the theory of linear matrix equations. Journal of Mathematics and Physics (MIT), 13(1–4):399–420, April 1934. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1934131399.

Lowan:1943:TRF

[LBA43] Arnold N. Lowan, G. Blanch, and M. Abramowitz. Table of and related functions. Journal of Mathematics and Physics (MIT), 22(1–4):51– 57, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194322151.

Levi-Civita:1934:GST

[LC34] Tullio Levi-Civita. A general survey of the theory of adiabatic invariants. Journal of Mathematics and Physics (MIT), 13(1–4):18–40, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193413118.

LaRosa:1954:GTW

[LC54] R. LaRosa and Herbert J. Carlin. A general theory of wideband matching with dissipative 4 poles. Journal of Mathematics and Physics (MIT), 33(1–4):331–345, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1954331331.

Lee:1932:SEN

[Lee32] Y. W. Lee. Synthesis of electric networks by means of the Fourier transforms of Laguerre’s functions. Journal of Mathematics and Physics (MIT), 11(1–4):83–113, April 1932. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193211183.

Lee:1967:IIP

[Lee67] P. A. Lee. On integrals involving parabolic cylinder functions. Journal of Mathematics and Physics (MIT), 46(1–4):215–219, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1967461215.

Lellouche:1967:APA

[Lel67] Gerald S. Lellouche. An application of Popov’s asymptotic stability criterion. Journal of Mathematics and Physics (MIT), 46(1–4):225– 236, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461225.

Lemaitre:1925:NSU

[Lem25] G. Lemaitre. Note on De Sitter’s universe. Journal of Mathematics and Physics (MIT), 4(1–4):188–192, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192541188.

Leutert:1951:CUA

[Leu51] Werner Leutert. On the convergence of unstable approximate solutions of the heat equation to the exact solution. Journal of Mathematics and Physics (MIT), 30(1–4):245–251, April 1951. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1951301245.

Levinson:1935:FTS

[Lev35] Norman Levinson. The Fourier transform solution of ordinary and partial differential equations. Journal of Mathematics and Physics (MIT), 14(1–4):195–227, April 1935. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1935141195.

Levinson:1937:TRN

[Lev37] Norman Levinson. A theorem relating non-vanishing and analytic functions. Journal of Mathematics and Physics (MIT), 16(1–4):185–190, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1937161185.

Levinson:1940:HTZ

[Lev40] Norman Levinson. On Hardy’s theorem on the zeros of the zeta function. Journal of Mathematics and Physics (MIT), 19(1–4):159–160, April 1940. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1940191159.

Levinson:1941:IIP

[Lev41] Norman Levinson. An integral inequality of the Phragmen Lindelof type. Journal of Mathematics and Physics (MIT), 20(1–4):89–98, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194120189.

Levinson:1943:NLD

[Lev43a] Norman Levinson. On a non-linear differential equation of the second order. Journal of Mathematics and Physics (MIT), 22(1–4):181– 187, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1943221181.

Levinson:1943:EPS

[Lev43b] Norman Levinson. On the existence of periodic solutions for second order differential equations with a forcing term. Journal of Mathematics and Physics (MIT), 22(1–4):41–48, April 1943. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194322141.

Levinson:1946:WRM

[Lev46] Norman Levinson. The Wiener (root mean square) error criterion in filter design and prediction. Journal of Mathematics and Physics (MIT), 25(1–4):261–278, April 1946. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1946251261.

Levinson:1947:HEW

[Lev47] Norman Levinson. A heuristic exposition of Wiener’s mathematical theory of prediction and filtering. Journal of Mathematics and Physics (MIT), 26(1–4):110–119, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1947261110.

Levinson:1949:ODE

[Lev49] Norman Levinson. An ordinary differential equation with an interval of stability, a separation point, and an interval of instability. Journal of Mathematics and Physics (MIT), 28(1–4):215–222, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281215.

Lewin:1963:EVB

[Lew63] Myrna Lewin. The effect of vertical barriers on progressing waves. Journal of Mathematics and Physics (MIT), 42(1–4):287–300, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1963421287.

Lomen:1967:FBP

[LF67] D. O. Lomen and L. L. Fontenot. Fluid behavior in parabolic containers undergoing vertical excitation. Journal of Mathematics and Physics (MIT), 46(1–4):43–53, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746143.

Lyon:1926:MSP

[LH26] W. V. Lyon and F. L. Hitchcock. A method of solving power networks by means of conjugate vectors. Journal of Mathematics and Physics (MIT), 5(1–4):122–125, January 1926. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192651122.

Lowan:1942:F

[LH42] Arnold N. Lowan and William Horenstein. On the function H(m, a, x) = exp(−ιx)F (m + 1 − ιa, 2m + 2; 2ιx). Journal of Mathematics and Physics (MIT), 21(1–4):264–283, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1942211264.

Lowan:1943:STF

[LH43] A. N. Lowan and A. Hillman. A short table of the first five zeros of the transcendental equation J_{0}(x)Y_{0}(kt) − J_{0}(kx)Y_{0}(x) = 0. Journal of Mathematics and Physics (MIT), 22(1–4):208–209, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1943221208.

Lifshitz:1942:FAO

[Lif42a] Jaime Lifshitz. On the Fourier analysis of orbits in the equatorial plane of a magnetic dipole. Journal of Mathematics and Physics (MIT), 21(1–4):94–116, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194221194.

Lifshitz:1942:SPP

[Lif42b] Jaime Lifshitz. On the stability of principal periodic orbits in the theory of primary cosmic rays. Journal of Mathematics and Physics (MIT), 21(1–4):284–292, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1942211284.

Lindsay:1924:AMAb

[Lin24] Robert Bruce Lindsay. On the atomic models of the alkali metals. Journal of Mathematics and Physics (MIT), 3(4):191–236, May 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192434191.

Lin:1941:MSA

[Lin41] Shih-Nge Lin. A method of successive approximations of evaluating the real and complex roots of cubic and higher-order equations. Journal of Mathematics and Physics (MIT), 20(1–4):231–242, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1941201231.

Lin:1947:NSC

[Lin47a] Shih-Nge Lin. Numerical solution of complex roots of quartic equations. Journal of Mathematics and Physics (MIT), 26(1–4):279–283, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261279.

Ling:1947:SNP

[Lin47b] Chih-Bing Ling. On the stresses in a notched plate under tension. Journal of Mathematics and Physics (MIT), 26(1–4):284–289, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1947261284.

Lin:1949:SFT

[Lin49] C. C. Lin. On the subsonic flow through circular and straight lattices of airfoils. Journal of Mathematics and Physics (MIT), 28(1–4):117– 130, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281117.

Ling:1952:TVI

[Lin52] Chih-Bing Ling. Tables of values of the integrals. Journal of Mathematics and Physics (MIT), 31(1–4):58–62, April 1952. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195231158.

Lin:1954:PTB

[Lin54] C. C. Lin. On a perturbation theory based on the method of characteristics. Journal of Mathematics and Physics (MIT), 33(1–4):117– 134, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331117.

Lin:1957:NGP

[Lin57] C. C. Lin. Note on Garabedian’s paper “Numerical Construction of Detached Shock Waves”. Journal of Mathematics and Physics (MIT), 36(1–4):206–209, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361206. See [Gar57].

Lipka:1922:GMC

[Lip22] Joseph Lipka. On the geometry of motion in curved N-space. Journal of Mathematics and Physics (MIT), 1(1):21–41, January 1922. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19221121.

Lipka:1923:CP

[Lip23a] Joseph Lipka. On conformal parallelism. Journal of Mathematics and Physics (MIT), 2(1–4):175–194, October 1923. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192321175.

Lipka:1923:HCE

[Lip23b] Joseph Lipka. On Hamilton’s canonical equations and infinitesimal contact transformations. Journal of Mathematics and Physics (MIT), 2(1–4):31–46, October 1923. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19232131.

Lipka:1923:IDS

[Lip23c] Joseph Lipka. On irreversible dynamical systems. Journal of Mathematics and Physics (MIT), 2(1–4):73–87, October 1923. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19232173.

Lipka:1924:RCR

[Lip24] Joseph Lipka. On Ricci’s coefficients of rotation. Journal of Mathematics and Physics (MIT), 3(1):7–23, January 1924. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1924317.

Littauer:1929:NTT

[Lit29] S. B. Littauer. A new Tauberian theorem with application to the summability of Fourier series and integrals. Journal of Mathematics and Physics (MIT), 8(1–4):216–234, April 1929. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192981216.

Lowan:1943:TFC

[LL43] Arnold N. Lowan and Jack Laderman. Table of Fourier coefficients. Journal of Mathematics and Physics (MIT), 22(1–4):136–147, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1943221136.

Leonard:1965:RKC

[LM65] A. Leonard and T. W. Mullikin. The resolvent kernel for a class of integral operators with difference kernels on a finite interval. Journal of Mathematics and Physics (MIT), 44(1–4):327–340, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441327.

Leutert:1951:CAS

[LO51] Werner Leutert and George G. O’Brien. On the convergence of approximate solutions of the wave equation to the exact solution. Journal of Mathematics and Physics (MIT), 30(1–4):252–256, April 1951.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1951301252.

Lotkin:1952:NIP

[Lot52] Mark Lotkin. A new integrating procedure of high accuracy. Journal of Mathematics and Physics (MIT), 31(1–4):29–34, April 1952. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195231129.

Lotkin:1953:NIP

[Lot53] M. Lotkin. A new integration procedure. Journal of Mathematics and Physics (MIT), 32(1–4):171–179, April 1953. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1953321171.

Lotkin:1954:SIN

[Lot54] Mark Lotkin. The solution by iteration of nonlinear integral equations. Journal of Mathematics and Physics (MIT), 33(1–4):346–355, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331346.

Lotkin:1956:NSL

[Lot56] Mark Lotkin. Note on the sensitivity of least squares solutions. Journal of Mathematics and Physics (MIT), 35(1–4):309–311, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1956351309.

Lotkin:1958:NIH

[Lot58] Mark Lotkin. The numerical integration of heat conduction equations. Journal of Mathematics and Physics (MIT), 37(1–4):178–187, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371178.

Loud:1949:LRE

[Lou49] Warren S. Loud. On the long-run error in the numerical solution of certain differential equations. Journal of Mathematics and Physics (MIT), 28(1–4):45–49, April 1949. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194928145.

Loud:1955:PSD

[Lou55] W. S. Loud. On periodic solutions of Duffing’s equation with damping. Journal of Mathematics and Physics (MIT), 34(1–4):173–178, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341173.

Loud:1961:LIM

[Lou61] W. S. Loud. The location of the invariant manifold for a perturbed autonomous system. Journal of Mathematics and Physics (MIT), 40(1–4): 87–102, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196140187.

Louchard:1965:HPB

[Lou65] G. Louchard. Hitting probabilities for Brownian motion in three dimensions. Journal of Mathematics and Physics (MIT), 44(1–4):177– 181, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441177.

Lowan:1945:PHC

[Low45] Arnold N. Lowan. On the problem of heat conduction in thin plates. Journal of Mathematics and Physics (MIT), 24(1–4):22–29, April 1945. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194524122.

Lee:1963:DEI

[LP63] K. S. H. Lee and C. H. Papas. Doppler effects in inhomogeneous anisotropic ionized gases. Journal of Mathematics and Physics (MIT), 42(1–4):189–199, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421189.

Lin:1948:FBC

[LR48] C. C. Lin and S. I. Rubinov. On the flow behind curved shocks. Journal of Mathematics and Physics (MIT), 27(1–4):105–129, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1948271105.

Lee:1960:EPJ

[LR60] Kiuck Lee and Lee G. Radosevich. Evaluation of ∂J_{ν}(z)/∂ν. Journal of Mathematics and Physics (MIT), 39(1–4):293–299, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1960391293.

Lin:1948:TDN

[LRT48] C. C. Lin, E. Reissner, and H. S. Tsien. On two-dimensional non-steady motion of a slender body in a compressible fluid. Journal of Mathematics and Physics (MIT), 27(1–4):220–231, April 1948. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1948271220.

Lowan:1943:TCN

[LS43] A. N. Lowan and Herbert Salzer. Table of coefficients in numerical integration formulae. Journal of Mathematics and Physics (MIT), 22(1–4): 49–50, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194322149.

Lowan:1944:CIW

[LS44] Arnold N. Lowan and Herbert E. Salzer. Coefficients for interpolation within a square grid in the complex plane. Journal of Mathematics and Physics (MIT), 23(1–4):156–166, April 1944. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1944231156.

Lowan:1945:TCN

[LS45] Arnold N. Lowan and Herbert E. Salzer. Table of coefficients for numerical integration without differences. Journal of Mathematics and Physics (MIT), 24(1–4):1–21, April 1945. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19452411.

Luke:1951:RAE

[LU51] Yudell L. Luke and Dolores Ufford. On the roots of algebraic equations. Journal of Mathematics and Physics (MIT), 30(1–4):94–101, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195130194.

Ludford:1951:BIP

[Lud51] G. S. S. Ludford. The behavior at infinity of the potential function of a two dimensional subsonic compressible flow. Journal of Mathematics and Physics (MIT), 30(1–4):117–130, April 1951. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1951301117.

Luke:1950:SNI

[Luk50] Yudell L. Luke. Some notes on integrals involving Bessel functions. Journal of Mathematics and Physics (MIT), 29(1–4):27–30, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195029127.

Luke:1952:ABF

[Luk52a] Yudell L. Luke. An associated Bessel function. Journal of Mathematics and Physics (MIT), 31(1–4):131–138, April 1952. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1952311131.

Luke:1952:CFI

[Luk52b] Yudell L. Luke. Coefficients to facilitate interpolation and integration of linear sums of exponential functions. Journal of Mathematics and Physics (MIT), 31(1–4):267–275, April 1952. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1952311267.

Luke:1955:SFE

[Luk55] Yudell L. Luke. Simple formulas for the evaluation of some higher transcendental functions. Journal of Mathematics and Physics (MIT), 34(1–4):298–307, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341298.

Luke:1958:PTT

[Luk58] Yudell L. Luke. The Padé table and the τ-method. Journal of Mathematics and Physics (MIT), 37(1–4):110–127, April 1958. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1958371110.

Luke:1959:ERT

[Luk59] Yudell L. Luke. On economic representations of transcendental functions. Journal of Mathematics and Physics (MIT), 38(1–4):279–294, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381279.

Lynch:1966:EBN

[Lyn66] Robert E. Lynch. Error bound on a numerical solution of a quasi-linear wave equation. Journal of Mathematics and Physics (MIT), 45(1–4): 370–382, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451370.

Lyon:1924:NME

[Lyo24] W. V. Lyon. Note on a method of evaluating the complex roots of a quartic equation. Journal of Mathematics and Physics (MIT), 3(3):

188–190, April 1924. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192433188.

Lipow:1955:RE

[LZ55] M. Lipow and S. A. Zwick. On the roots of the equation:Y_{1}(mx)[xJ_{1}(x)− BJ_{0}(x)] − J_{1}(mx)[xY_{1}(x) − BY_{0}(x)] = 0. Journal of Mathematics and Physics (MIT), 34(1–4):308–315, April 1955. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1955341308.

Moravcsik:1965:LSA

[MA65] Michael J. Moravcsik and Henry D. I. Abarbanel. Least square analysis and significance tests for a set of functions of discrete abscissae. Journal of Mathematics and Physics (MIT), 44(1–4):255–265, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441255.

Maccoll:1941:TME

[Mac41] L. A. Maccoll. Trajectories of monoenergetic electrons, in an arbitrary static electromagnetic field, in the neighborhood of a given trajectory. Journal of Mathematics and Physics (MIT), 20(1–4):355–369, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201355.

MacDonald:1949:PCH

[Mac49] A. D. MacDonald. Properties of the confluent hypergeometric function. Journal of Mathematics and Physics (MIT), 28(1–4):183–191, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281183.

Magnusson:1942:NMS

[Mag42] Philip C. Magnusson. A numerical method of solving integral equations in two independent variables. Journal of Mathematics and Physics (MIT), 21(1–4):250–263, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1942211250.

Mahler:1927:SAA

[Mah27] Kurt Mahler. The spectrum of an array and its application to the study of the translation properties of a simple class of arithmetical functions: Part Two. On the translation properties of a simple class of arithmetical functions. Journal of Mathematics and Physics (MIT), 6(1–4):158– 163, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761158.

Malin:1936:EFG

[Mal36] Henry Malin. On entire functions with gaps. Journal of Mathematics and Physics (MIT), 15(1–4):90–96, April 1936. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193615190.

Manneback:1922:IES

[Man22] Charles Manneback. An integral equation for skin effect in parallel conductors. Journal of Mathematics and Physics (MIT), 1(3):123– 146, April 1922. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192213123.

Manwell:1955:FPC

[Man55] A. R. Manwell. A family of plane compressible flows past a certain semi-infinite body. Journal of Mathematics and Physics (MIT), 34(1–4):113–118, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341113.

Martin:1950:PRP

[Mar50a] M. H. Martin. Plane rotational Prandtl–Meyer flows. Journal of Mathematics and Physics (MIT), 29(1–4):76–89, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195029176.

Martin:1950:SPR

[Mar50b] M. H. Martin. Steady, plane, rotational Prandtl–Meyer flow of a polytropic gas. Journal of Mathematics and Physics (MIT), 29(1–4): 263–281, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291263.

Marx:1953:RRP

[Mar53] Imanuel Marx. Recurrence relations for prolate spheroidal wave functions. Journal of Mathematics and Physics (MIT), 32(1–4):269–275, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321269.

Marguerre:1956:VSP

[Mar56] K. Marguerre. Vibration and stability problems of beams treated by matrices. Journal of Mathematics and Physics (MIT), 35(1–4):28–43, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195635128.

Martin:1957:FTR

[Mar57] A. I. Martin. On a formula for the torsional rigidity of thinsymmetrical-sections. Journal of Mathematics and Physics (MIT), 36(1–4):20–25, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195736120.

Martin:1963:UHF

[Mar63a] M. H. Martin. On the uniqueness of harmonic functions under boundary conditions. Journal of Mathematics and Physics (MIT), 42(1–4): 1–13, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19634211.

Marx:1963:RCN

[Mar63b] Imanuel Marx. Remark concerning a non-linear sequence-to-sequence transform. Journal of Mathematics and Physics (MIT), 42(1–4):334–335, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1963421334. See remark [Tuc66].

Marcinkowski:1967:PWS

[Mar67] Chester J. Marcinkowski. Plane wave scattering by a linearly varying surface impedance. Journal of Mathematics and Physics (MIT), 46(1–4):311–324, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461311.

Mayer:1939:CD

[May39] Walther Mayer. Charaktersysteme und Dualitätstheoreme. (German) [Character systems and duality theorems]. Journal of Mathematics and Physics (MIT), 18(1–4):1–27, April 1939. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19391811.

Meyers:1950:BIF

[MB50] Leroy F. Meyers and Arthur Bard. Best interpolation formulas. Journal of Mathematics and Physics (MIT), 29(1–4):198–206, April 1950.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291198.

Mann:1956:IAD

[MBC56] W. Robert Mann, C. L. Bradshaw, and J. Grady Cox. Improved approximations to differential equations by difference equations. Journal of Mathematics and Physics (MIT), 35(1–4):408–415, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1956351408.

McLachlan:1946:MFT

[McL46] N. W. McLachlan. Mathieu functions and their classification. Journal of Mathematics and Physics (MIT), 25(1–4):209–240, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1946251209. See corrections

[McL47a].

McLachlan:1947:CPM

[McL47a] N. W. McLachlan. Corrections to paper “Mathieu Functions and Their Classification”. Journal of Mathematics and Physics (MIT), 26(1–4):78, April 1947. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194726178. See [McL46].

McLachlan:1947:MFF

[McL47b] N. W. McLachlan. Mathieu functions of fractional order. Journal of Mathematics and Physics (MIT), 26(1–4):29–41, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194726129.

McMillan:1939:NTN

[McM39] Brockway McMillan. A note on transcendental numbers. Journal of Mathematics and Physics (MIT), 18(1–4):28–33, April 1939. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193918128.

Meligy:1963:EEI

[ME63] A. S. Meligy and E. M. El Gazzy. Expansion of the exponential integral in terms of Bessel functions. Journal of Mathematics and Physics (MIT), 42(1–4):157–161, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421157.

Meligy:1964:EGE

[ME64] A. S. Meligy and E. M. El Gazzy. Expansion of the generalized exponential integral in terms of Bessel functions. Journal of Mathematics and Physics (MIT), 43(1–4):60–64, April 1964. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196443160.

Mei:1966:NGW

[Mei66] C. C. Mei. Nonlinear gravity waves in a thin sheet of viscous fluid. Journal of Mathematics and Physics (MIT), 45(1–4):266–288, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451266.

Mersman:1950:NFS

[Mer50] W. A. Mersman. A new form of solution of Hermite’s equation. Journal of Mathematics and Physics (MIT), 29(1–4):191–197, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1950291191.

Mezger:1958:EIA

[Mez58] Fritz W. Mezger. Evaluation of integrals arising in particle attenuation problems. Journal of Mathematics and Physics (MIT), 37(1–4):79– 88, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195837179.

Moore:1931:GAP

[MF31a] C. L. E. Moore and P. Franklin. The geometry of algebraic Pfaffians. Journal of Mathematics and Physics (MIT), 10(1–4):19–49, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193110119.

Moore:1931:PPF

[MF31b] C. L. E. Moore and P. Franklin. Pfaffians in parametric form. Journal of Mathematics and Physics (MIT), 10(1–4):95–105, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193110195.

Moore:1932:DP

[MF32] C. L. E. Moore and P. Franklin. Dual Pfaffians. Journal of Mathematics and Physics (MIT), 11(1–4):12–26, April 1932. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193211112.

Martin:1954:ETP

[MG54] John C. Martin and Nathan Gerber. The effect of thickness on pitching airfoils at supersonic speeds. Journal of Mathematics and Physics (MIT), 33(1–4):46–56, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195433146.

Murphy:1964:WCFa

[MG64a] E. L. Murphy and R. H. Good, Jr. WKB connection formulas. Journal of Mathematics and Physics (MIT), 43(1–4):251–254, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431251.

Murphy:1964:WCFb

[MG64b] E. L. Murphy and R. H. Good, Jr. WKB connection formulas. Journal of Mathematics and Physics (MIT), 43(1–4):261–262, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1964431261.

Miles:1949:CIE

[Mil49] John W. Miles. On certain integral equations in diffraction theory. Journal of Mathematics and Physics (MIT), 28(1–4):223–226, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281223.

Miles:1954:LEN

[Mil54] John W. Miles. Linearization of the equations of non-steady flow in a compressible fluid. Journal of Mathematics and Physics (MIT), 33(1–4):135–143, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1954331135.

Miranker:1963:SPE

[Mir63] W. L. Miranker. Singular perturbation eigenvalues by a method of undetermined coefficients. Journal of Mathematics and Physics (MIT), 42(1–4):47–58, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196342147.

Mireles:1966:IPE

[Mir66] R. Mireles. The inverse problem of electromagnetic scattering theory. i. a uniqueness theorem for cylinders. Journal of Mathematics and Physics (MIT), 45(1–4):179–187, April 1966. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1966451179.

Marshall:1959:NTB

[ML59a] F. J. Marshall and H. F. Ludloff. A note on Timoshenko beam theory. Journal of Mathematics and Physics (MIT), 38(1–4):175–180, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381175.

Morduchow:1959:CMA

[ML59b] Morris Morduchow and Lionel Levin. Comparison of the method of averages with the method of least squares: Fitting a parabola. Journal of Mathematics and Physics (MIT), 38(1–4):181–192, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1959381181.

Miller:1953:MBE

[MM53] K. S. Miller and F. J. Murray. A mathematical basis for an error analysis of differential analyzers. Journal of Mathematics and Physics (MIT), 32(1–4):136–163, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321136.

Montroll:1946:NBF

[Mon46] Elliott W. Montroll. A note on Bessel functions of purely imaginary argument. Journal of Mathematics and Physics (MIT), 25(1–4):37– 48, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194625137.

Moore:1922:H

[Moo22a] C. L. E. Moore. Hyperquaternions. Journal of Mathematics and Physics (MIT), 1(2):63–77, March 1922. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19221263.

Moore:1922:NNP

[Moo22b] C. L. E. Moore. Note on the normal planes to a surface in a space of four dimensions. Journal of Mathematics and Physics (MIT), 1(3): 147–156, April 1922. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192213147.

Moore:1923:NVD

[Moo23] C. L. E. Moore. Note on the vanishing of the determinant of the second fundamental form of a hypersurface. Journal of Mathematics and Physics (MIT), 2(1–4):125–130, October 1923. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192321125.

Moore:1924:NDS

[Moo24] C. L. E. Moore. Note on developable surfaces in hyperspace. Journal of Mathematics and Physics (MIT), 3(1):1–6, January 1924. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1924311.

Moore:1925:MVT

[Moo25] C. L. E. Moore. Minimal varieties of two and three dimensions whose element of arc is a perfect square. Journal of Mathematics and Physics (MIT), 4(1–4):167–178, January 1925. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192541167.

Moore:1926:GGR

[Moo26] C. L. E. Moore. Grassmannian geometry in Riemannian space. Journal of Mathematics and Physics (MIT), 5(1–4):191–200, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192651191.

Moore:1927:NSN

[Moo27] C. L. E. Moore. Note on surfaces in a non-Riemannian space. Journal of Mathematics and Physics (MIT), 6(1–4):39–49, April 1927. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19276139.

Moore:1928:NCR

[Moo28a] C. L. E. Moore. Note on congruences in Riemannian space. Journal of Mathematics and Physics (MIT), 7(1–4):185–192, April 1928. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192871185.

Moore:1928:PG

[Moo28b] C L. E. Moore. Parallel geodesics. Journal of Mathematics and Physics (MIT), 7(1–4):81–85, April 1928. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19287181.

Moore:1929:GCC

[Moo29a] C. L. E. Moore. Geometry of a complex of curves in a Riemannian space of n dimensions. Journal of Mathematics and Physics (MIT), 8(1–4):247–265, April 1929. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192981247.

Moore:1929:HWC

[Moo29b] C. L. E. Moore. Hypersurfaces for which the Codazzi equations take the simplest form. Journal of Mathematics and Physics (MIT), 8(1–4): 141–147, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192981141.

Moore:1929:PDS

[Moo29c] C. L. E. Moore. Principal directions in a subspace V_{m} immersed in a Riemannian space V_{n}. Journal of Mathematics and Physics (MIT), 8(1–4):65–69, April 1929. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19298165.

Moore:1930:DPG

[Moo30] C. L. E. Moore. Differential projective geometry of a system of Pfaffian differential equations. Journal of Mathematics and Physics (MIT), 9(1–4):320–332, April 1930. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193091324.

Moon:1937:TPF

[Moo37] Parry Moon. A table of Planck’s function from 3500 to 8000^{o}K. Journal of Mathematics and Physics (MIT), 16(1–4):133–157, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1937161133.

Moon:1940:TFR

[Moo40] Parry Moon. A table of Fresnel reflections. Journal of Mathematics and Physics (MIT), 19(1–4):1–33, April 1940. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19401911.

Morawetz:1954:ASS

[Mor54] Cathleen S. Morawetz. Asymptotic solutions of the stability equations of a compressible fluid. Journal of Mathematics and Physics (MIT), 33(1–4):1–26, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19543311.

Morton:1957:GAV

[Mor57] K. W. Morton. A generalisation of the antithetic variate technique for evaluating integrals. Journal of Mathematics and Physics (MIT), 36(1–4):289–293, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361289.

Morrison:1965:DFE

[Mor65a] J. A. Morrison. Dual formulation for the eigenfunctions corresponding to the bandpass kernel, in the case of degeneracy. Journal of Mathematics and Physics (MIT), 44(1–4):313–326, April 1965. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1965441313.

Morrison:1965:EFF

[Mor65b] J. A. Morrison. Eigenfunctions of the finite Fourier transform operator over a hyperellipsoidal region. Journal of Mathematics and Physics (MIT), 44(1–4):245–254, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1965441245.

Moser:1958:RPP

[Mos58] J. Moser. Remarks on the preceding paper of Louis N. Howard. Journal of Mathematics and Physics (MIT), 37(1–4):299–304, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1958371299.

Mow:1967:ASE

[Mow67] V. C. Mow. An asymptotic solution for an elastica-viscous liquid between rotating spheres. Journal of Mathematics and Physics (MIT), 46(1–4):267–280, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461267.

Mathematical:1946:TSF

[MP46] The Mathematical and Tables Project. Table of the Struve functions L_{v}(x) and H_{v}(x). Journal of Mathematics and Physics (MIT), 25(1–4): 252–259, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1946251252.

Moon:1945:VDA

[MS45] Parry Moon and Domina Eberle Spencer. Visual dark adaptation: a mathematical formulation. Journal of Mathematics and Physics (MIT), 24(1–4):65–105, April 1945. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194524165.

Moon:1946:AEP

[MS46] Parry Moon and Domina Eberle Spencer. Analytic expressions in photometry and colorimetry. Journal of Mathematics and Physics (MIT), 25(1–4):111–190, April 1946. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1946251111.

Meyers:1950:BAI

[MS50] Leroy F. Meyers and Arthur Sard. Best approximate integration formulas. Journal of Mathematics and Physics (MIT), 29(1–4):118–123, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291118.

Melvin:1957:ECP

[MS57] M. A. Melvin and N. V. V. J. Swamy. Evaluation of certain physically interesting integrals and hypergeometric sums. Journal of Mathematics and Physics (MIT), 36(1–4):157–163, April 1957. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1957361157.

Moon:1961:SSB

[MS61] Parry Moon and Domina Eberle Spencer. Series solutions of Bôcher equations. Journal of Mathematics and Physics (MIT), 40(1–4):116– 124, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401116.

McClellan:1966:SNH

[MS66] R. P. McClellan and G. W. Stroke. Solution of the non-homogeneous Helmholtz equation for optical gratings with perfectly conducting boundaries. Journal of Mathematics and Physics (MIT), 45(1–4):383– 390, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451383.

Muckenhoupt:1929:APF

[Muc29] Carl F. Muckenhoupt. Almost periodic functions and vibrating systems. Journal of Mathematics and Physics (MIT), 8(1–4):163–199, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192981163.

Muller:1955:CII

[Mul55] G. M. Muller. On certain infinite integrals involving Bessel functions. Journal of Mathematics and Physics (MIT), 34(1–4):179–181, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341179.

Muller:1956:DOC

[Mul56] Mervin E. Muller. On discrete operators connected with the Dirichlet problem. Journal of Mathematics and Physics (MIT), 35(1–4):89– 113, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195635189.

Mulligan:1967:RBS

[Mul67] Bernard Mulligan. A relationship between solutions of Liouville’s equation. Journal of Mathematics and Physics (MIT), 46(1–4):211–214, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461211.

Munakata:1952:VES

[Mun52] K. Munakata. On the vibration and elastic stability of a rectangular plate clamped at its four edges. Journal of Mathematics and Physics (MIT), 31(1–4):69–74, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195231169.

Murray:1933:CLW

[Mur33] F. H. Murray. The capacity of a long wire. Journal of Mathematics and Physics (MIT), 12(1–4):306–310, April 1933. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1933121306.

Murray:1943:SLO

[Mur43] F. J. Murray. The solution of linear operator equations. Journal of Mathematics and Physics (MIT), 22(1–4):148–157, April 1943. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1943221148.

Murray:1967:SMD

[Mur67] J. D. Murray. A simple method for determining asymptotic forms of Navier–Stokes solutions for a class of large Reynolds number flows. Journal of Mathematics and Physics (MIT), 46(1–4):1–20, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19674611.

Murray:1968:SPC

[Mur68] J. D. Murray. Singular perturbations of a class of nonlinear hyperbolic and parabolic equations. Journal of Mathematics and Physics (MIT), 47(1–4):111–133, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471111.

Motzkin:1952:ALE

[MW52] T. S. Motzkin and W. Wasow. On the approximation of linear elliptic differential equations by difference equations with positive coefficients. Journal of Mathematics and Physics (MIT), 31(1–4):253– 259, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311253.

Nagamatsu:1951:CCF

[Nag51] H. T. Nagamatsu. Circular cylinder and flat plate airfoil in a flow field with parabolic velocity distribution. Journal of Mathematics and Physics (MIT), 30(1–4):131–139, April 1951. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1951301131.

Nagaraja:1964:NCI

[Nag64] K. S. Nagaraja. A note on certain integrals. Journal of Mathematics and Physics (MIT), 43(1–4):55–59, April 1964. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196443155.

Nagaraja:1965:CV

[Nag65] K. S. Nagaraja. Concerning the value of. Journal of Mathematics and Physics (MIT), 44(1–4):182–188, April 1965. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1965441182.

Nagelberg:1967:ATM

[Nag67] Elliott R. Nagelberg. Approximate theory of microwave interaction with bounded gyrotropie media. Journal of Mathematics and Physics (MIT), 46(1–4):54–62, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746154.

Navot:1961:EEM

[Nav61] Israel Navot. An extension of the Euler–Maclaurin summation formula to functions with a branch singularity. Journal of Mathematics and Physics (MIT), 40(1–4):271–276, April 1961. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1961401271.

Navot:1962:FEE

[Nav62] Israel Navot. A further extension of the Euler–Maclaurin summation formula. Journal of Mathematics and Physics (MIT), 41(1–4):155– 163, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411155.

Nayfeh:1965:PMT

[Nay65a] Ali Hasan Nayfeh. A perturbation method for treating nonlinear oscillation problems. Journal of Mathematics and Physics (MIT), 44(1–4): 368–374, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441368.

Naylor:1965:EPC

[Nay65b] D. Naylor. An eigenvalue problem in cylindrical harmonics. Journal of Mathematics and Physics (MIT), 44(1–4):391–402, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441391.

Naylor:1966:SEP

[Nay66] D. Naylor. On some eigenvalue problems involving modified Bessel functions. Journal of Mathematics and Physics (MIT), 45(1–4):95– 103, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196645195.

Nayfeh:1967:ASE

[Nay67] Ali Hasan Nayfeh. Asymptotic solutions of an eigenvalue problem with two turning points — heat transfer in a tube. Journal of Mathematics and Physics (MIT), 46(1–4):349–354, April 1967. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1967461349.

Newman:1960:NMS

[New60] D. J. Newman. Numerical method for solution of an elliptic Cauchy problem. Journal of Mathematics and Physics (MIT), 39(1–4):72–75, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196039172.

Nielsen:1947:ALS

[NG47] K. L. Nielsen and L. Goldstein. An algorithm for least squares. Journal of Mathematics and Physics (MIT), 26(1–4):120–132, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1947261120.

Ng:1967:RFC

[Ng67] E. Wai-Kwok Ng. Recursive formulae for the computation of certain integrals of Bessel functions. Journal of Mathematics and Physics (MIT), 46(1–4):223–224, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461223.

Lin:1943:MFR

[nL43] Shin nge Lin. A method for finding roots of algebraic equations. Journal of Mathematics and Physics (MIT), 22(1–4):60–77, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194322160.

Noble:1958:CDI

[Nob58] B. Noble. Certain dual integral equations. Journal of Mathematics and Physics (MIT), 37(1–4):128–136, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1958371128.

Nohel:1959:CND

[Noh59] John A. Nohel. A class of nonlinear delay differential equations. Journal of Mathematics and Physics (MIT), 38(1–4):295–311, April 1959.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1959381295.

Nemenyi:1948:SGP

[NP48] P. Neményi and R. Prim. Some geometric properties of plane gas flow. Journal of Mathematics and Physics (MIT), 27(1–4):130–135, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271130.

Olmstead:1966:ESOb

[OB66] W. E. Olmstead and W. J. Byrne, Jr. An exact solution for Oseen flow past a half plane and a vertical force singularity. Journal of Mathematics and Physics (MIT), 45(1–4):402–407, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451402.

Oberhettinger:1955:ASF

[Obe55] F. Oberhettinger. On asymptotic series for functions occurring in the theory of diffraction of waves by wedges. Journal of Mathematics and Physics (MIT), 34(1–4):245–255, April 1955. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1955341245.

Oberhettinger:1958:DBF

[Obe58] F. Oberhettinger. On the derivative of Bessel functions with respect to the order. Journal of Mathematics and Physics (MIT), 37(1–4): 75–78, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195837175.

Olmstead:1966:NOF

[OH66] W. E. Olmstead and D. L. Hector. On the nonuniqueness of Oseen flow past a half plane. Journal of Mathematics and Physics (MIT), 45(1–4): 408–417, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451408.

OBrien:1950:SNS

[OHK50] George G. O’Brien, Morton A. Hyman, and Sidney Kaplan. A study of the numerical solution of partial differential equations. Journal of Mathematics and Physics (MIT), 29(1–4):223–251, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291223.

See remarks [Hym51].

Okubo:1952:SDA

[Ōku52] H. Ōkubo. The stress distribution in an aeolotropic circular disk compressed diametrically. Journal of Mathematics and Physics (MIT), 31(1–4):75–83, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195231175.

Olcer:1967:GHF

[Ölç67] N. Y. Ölçer. On a general heat flow problem. Journal of Mathematics and Physics (MIT), 46(1–4):99–106, April 1967. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196746199.

Olmstead:1965:NOF

[Olm65] W. E. Olmstead. A note on Oseen flow in the plane. Journal of Mathematics and Physics (MIT), 44(1–4):240–244, April 1965. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1965441240.

Olmstead:1966:ESOa

[Olm66] W. E. Olmstead. An exact solution for Oseen flow past a half plane and a horizontal force singularity. Journal of Mathematics and Physics (MIT), 45(1–4):156–161, April 1966. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1966451156.

OMathuna:1963:RSD

[O’M63] Diarmuid O’Mathuna. Rotationally symmetric deformations in helicoidal shells. Journal of Mathematics and Physics (MIT), 42(1–4): 85–111, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196342185.

Oono:1950:SFT

[Oon50] Yosibo Oono. Synthesis of a finite 2n-terminal network by a group of networks each of which contains only one Ohmic resistance. Journal of Mathematics and Physics (MIT), 29(1–4):13–26, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195029113.

Opatowski:1941:MEP

[Opa41] I. Opatowski. On the motion of an electric particle. Journal of Mathematics and Physics (MIT), 20(1–4):418–424, April 1941. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1941201418.

Olsson:1940:PBE

[OR40] R. Gran Olsson and Eric Reissner. A problem of buckling of elastic plates of variable thickness. Journal of Mathematics and Physics (MIT), 19(1–4):131–139, April 1940. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1940191131.

Ortenblad:1930:MTP

[Ort30] Alberto Ortenblad. Mathematical theory of the process of consolidation of mud deposits. Journal of Mathematics and Physics (MIT), 9(1–4): 73–149, April 1930. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19309173.

Osborn:1966:AEN

[Osb66] John E. Osborn. Approximation of the eigenvalues of non self-adjoint operators. Journal of Mathematics and Physics (MIT), 45(1–4):391– 401, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451391.

Paley:1933:OM

[Pal33] R. E. A. C. Paley. On orthogonal matrices. Journal of Mathematics and Physics (MIT), 12(1–4):311–320, April 1933. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1933121311.

Papas:1949:RTS

[Pap49] Charles H. Papas. Radiation from a transverse slot in an infinite cylinder. Journal of Mathematics and Physics (MIT), 28(1–4):227–236, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281227.

Papas:1954:ASC

[Pap54] Charles H. Papas. An application of Sommerfeld’s complex order wave functions to antenna theory. Journal of Mathematics and Physics (MIT), 33(1–4):269–275, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1954331269.

Papoulis:1958:SNL

[Pap58] A. Papoulis. Strongly non-linear oscillations. Journal of Mathematics and Physics (MIT), 37(1–4):147–156, April 1958. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1958371147.

Parke:1949:OA

[Par49] Nathan Grieb Parke III. Optical algebra. Journal of Mathematics and Physics (MIT), 28(1–4):131–139, April 1949. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1949281131.

Park:1954:APB

[Par54] David Park. Asymptotic properties of Bessel functions and the radiation from a synchrotron. Journal of Mathematics and Physics (MIT), 33(1–4):179–184, April 1954. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1954331179.

Paria:1957:ACP

[Par57] Gunadhar Paria. Axisymmetric consolidation for a porous elastic material containing a fluid. Journal of Mathematics and Physics (MIT), 36(1–4):338–346, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361338.

Paul:1959:WPA

[Pau59] David I. Paul. Wave propagation in acoustics using the saddle point method. Journal of Mathematics and Physics (MIT), 38(1–4):1–15, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19593811.

Paul:1961:SES

[Pau61] David I. Paul. Some exact solutions to the two-dimensional spin wave Hamiltonian. Journal of Mathematics and Physics (MIT), 40(1–4): 249–252, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401249.

Payne:1953:STF

[Pay53] W. T. Payne. Spinor theory of four-terminal networks. Journal of Mathematics and Physics (MIT), 32(1–4):19–33, April 1953. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195332119.

Payne:1959:RFS

[Pay59] L. E. Payne. Representation formulas for solutions of a class of partial differential equations. Journal of Mathematics and Physics (MIT), 38(1–4):145–149, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1959381145.

Payne:1960:NIP

[Pay60] L. E. Payne. A note on inequalities for plate eigenvalues. Journal of Mathematics and Physics (MIT), 39(1–4):155–159, April 1960. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1960391155.

Payton:1964:BSC

[Pay64] R. G. Payton. Bond stress in cylindrical shells subjected to an end velocity step. Journal of Mathematics and Physics (MIT), 43(1–4): 169–190, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431169.

Pridmore-Brown:1964:RFM

[PB64] D. C. Pridmore-Brown. The radiation field of a magnetic dipole antenna in a conical sheath. Journal of Mathematics and Physics (MIT), 43(1–4):199–217, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431199.

Pridmore-Brown:1968:WHS

[PB68] D. C. Pridmore-Brown. A Wiener–Hopf solution of a radiation problem in conical geometry. Journal of Mathematics and Physics (MIT), 47(1– 4):79–94, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196847179.

Podolsky:1955:DGM

[PD55] Boris Podolsky and Harry Denman. A derivation of generalized macroscopic electrodynamic equations. Journal of Mathematics and Physics (MIT), 34(1–4):198–207, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1955341198.

Pease:1937:RBS

[Pea37] E. M. J. Pease. On the relation between simple distillation and true boiling point curves of complex mixtures the components of which obey Raoult’s Law. Journal of Mathematics and Physics (MIT), 16(1–4): 202–214, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1937161202.

Pearson:1960:IRS

[Pea60] Carl E. Pearson. Integral representation for solutions of Bessel’s difference equation. Journal of Mathematics and Physics (MIT), 39(1–4): 287–292, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391287.

Pearson:1968:DEB

[Pea68a] Carl E. Pearson. On a differential equation of boundary layer type. Journal of Mathematics and Physics (MIT), 47(1–4):134–154, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471134.

Pearson:1968:NLO

[Pea68b] Carl E. Pearson. On non-linear ordinary differential equations of boundary layer type. Journal of Mathematics and Physics (MIT), 47(1–4):351–358, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471351.

Pennell:1928:GOA

[Pen28] W. O. Pennell. A general operational analysis. Journal of Mathematics and Physics (MIT), 7(1–4):24–38, April 1928. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19287124.

Peter:1957:PDC

[Pet57] Martin Peter. A property of dielectric constants of dielectrics in thermal equilibrium. Journal of Mathematics and Physics (MIT), 36(1–4):347– 350, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361347.

Petryshyn:1965:TVM

[Pet65] W. V. Petryshyn. On two variants of a method for the solution of linear equations with unbounded operators and their applications. Journal of Mathematics and Physics (MIT), 44(1–4):297–312, April 1965.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441297.

Phillips:1922:ESA

[Phi22a] H. B. Phillips. The equation of state, with applications to viscosity. Journal of Mathematics and Physics (MIT), 1(1):42–53, January 1922. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19221142.

Phillips:1922:NET

[Phi22b] H. B. Phillips. Note on Einstein’s theory of gravitation. Journal of Mathematics and Physics (MIT), 1(3):177–190, April 1922. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192213177.

Phillips:1933:FLB

[Phi33] H. B. Phillips. Faraday’s Law as a basis of electromagnetic theory. Journal of Mathematics and Physics (MIT), 12(1–4):259–273, April 1933. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1933121259.

Phillips:1934:ESD

[Phi34a] H. B. Phillips. Effect of surface discontinuity on the distribution of potential. Journal of Mathematics and Physics (MIT), 13(1–4):261– 267, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1934131261.

Phillips:1934:SF

[Phi34b] H. B. Phillips. Stress functions. Journal of Mathematics and Physics (MIT), 13(1–4):421–425, April 1934. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1934131421.

Phillips:1948:CDP

[Phi48] Aris Phillips. Calculation of the displacements in plastic torsion. Journal of Mathematics and Physics (MIT), 27(1–4):270–273, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1948271270.

Pinney:1946:LFM

[Pin46] Edmund Pinney. Laguerre functions in the mathematical foundations of the electromagnetic theory of the paraboloidal reflector. Journal of Mathematics and Physics (MIT), 25(1–4):49–79, April 1946. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194625149.

Pinney:1947:EFP

[Pin47] Edmund Pinney. Electromagnetic fields in a paraboloidal reflector. Journal of Mathematics and Physics (MIT), 26(1–4):42–55, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194726142.

Pinney:1951:TUW

[Pin51] Edmund Pinney. A theorem of use in wave theory. Journal of Mathematics and Physics (MIT), 30(1–4):1–10, April 1951. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19513011.

Pinney:1957:SDB

[Pin57] Edmund Pinney. Some discontinuous Bessel integrals. Journal of Mathematics and Physics (MIT), 36(1–4):362–370, April 1957. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1957361362.

Pitt:1937:TAC

[Pit37] H. R. Pitt. A theorem on absolutely convergent trigonometrical series. Journal of Mathematics and Physics (MIT), 16(1–4):191–195, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1937161191.

Pan:1956:CVF

[PK56] L. J. Pan and Y. H. Kuo. Compressible viscous flow past a wedge moving at hypersonic speeds. Journal of Mathematics and Physics (MIT), 35(1–4):179–193, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351179.

Polya:1952:RFP

[Pól52] G. Pólya. Remarks on the foregoing paper. Journal of Mathematics and Physics (MIT), 31(1–4):55–57, April 1952. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195231155.

Pomraning:1968:RCF

[Pom68] G. C. Pomraning. Reciprocal and canonical forms of variational problems involving linear operators. Journal of Mathematics and Physics (MIT), 47(1–4):155–169, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471155.

Poritsky:1932:FDT

[Por32] Hillil Poritsky. The field due to two equally charged parallel conducting cylinders. Journal of Mathematics and Physics (MIT), 11(1–4):213– 217, April 1932. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1932111213.

Poritsky:1942:VFF

[Por42] H Poritsky. Vortices in fluid flow. Journal of Mathematics and Physics (MIT), 21(1–4):218–239, April 1942. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1942211218.

Poritsky:1960:PCC

[Por60] H. Poritsky. Potential of a charged cylinder between two parallel grounded planes. Journal of Mathematics and Physics (MIT), 39(1–4): 35–48, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196039135.

Payne:1956:RCE

[PPW56] L. E. Payne, G. Pólya, and H. F. Weinberger. On the ratio of consecutive eigenvalues. Journal of Mathematics and Physics (MIT), 35(1–4): 289–298, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351289.

Prager:1957:GHT

[Pra57] William Prager. A generalization of Hitchcock’s transportation problem. Journal of Mathematics and Physics (MIT), 36(1–4):99–106, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195736199.

Prim:1949:UFG

[Pri49] R. C. Prim. On the uniqueness of flows with given streamlines. Journal of Mathematics and Physics (MIT), 28(1–4):50–53, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194928150.

Project:1944:T

[Pro44] The Mathematical Tables Project. Table of. Journal of Mathematics and Physics (MIT), 23(1–4):45–60, April 1944. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194423145.

Parzen:1956:AGS

[PS56] Emanuel Parzen and Norman Shiren. Analysis of a general system for the detection of amplitude-modulated noise. Journal of Mathematics and Physics (MIT), 35(1–4):278–288, April 1956. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1956351278.

Purcell:1953:VMS

[Pur53] Everett W. Purcell. The vector method of solving simultaneous linear equations. Journal of Mathematics and Physics (MIT), 32(1–4):180– 183, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321180.

Price:1966:AOM

[PVW66] Harvey S. Price, Richard S. Varga, and Joseph E. Warren. Application of oscillation matrices to diffusion-convection equations. Journal of Mathematics and Physics (MIT), 45(1–4):301–311, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451301.

Phillips:1923:NDP

[PW23] H. B. Phillips and N. Wiener. Nets and the Dirichlet problem. Journal of Mathematics and Physics (MIT), 2(1–4):105–124, October 1923. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192321105.

Payne:1954:NBH

[PW54] L. E. Payne and H. F. Weinberger. New bounds in harmonic and biharmonic problems. Journal of Mathematics and Physics (MIT), 33(1–4): 291–307, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331291.

Payne:1960:FKI

[PW60] L. E. Payne and H. F. Weinberger. A Faber–Krahn inequality for wedge-like membranes. Journal of Mathematics and Physics (MIT), 39(1–4):182–188, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391182.

Protter:1965:CCC

[PW65] M. H. Protter and H. F. Weinberger. On the capacity of composite conductors. Journal of Mathematics and Physics (MIT), 44(1–4):375– 383, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441375.

Radlow:1959:DLT

[Rad59] James Radlow. On the double Laplace transforms of some Green’s functions. Journal of Mathematics and Physics (MIT), 38(1–4):203– 205, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381203.

Rainville:1953:HCP

[Rai53] Earl D. Rainville. A heat conduction problem and the product of two error functions. Journal of Mathematics and Physics (MIT), 32(1–4): 43–47, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195332143.

Ralston:1956:PBH

[Ral56] Anthony Ralston. On the problem of buckling of a hyperbolic paraboloidal shell loaded by its own weight. Journal of Mathematics and Physics (MIT), 35(1–4):53–59, April 1956. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195635153.

Rand:1950:PMF

[Ran50] Robert C. Rand. Prandtl–Meyer flow behind a curved shock wave. Journal of Mathematics and Physics (MIT), 29(1–4):124–132, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1950291124.

Ranger:1964:CBV

[Ran64] K. B. Ranger. On a class of boundary value problems for the biharmonic equation and the Stokes repeated operator equation. Journal of Mathematics and Physics (MIT), 43(1–4):75–82, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196443175.

Ranger:1965:SPG

[Ran65] K. B. Ranger. Some properties of generalized axially symmetric potentials. Journal of Mathematics and Physics (MIT), 44(1–4):341–352, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441341.

Ranger:1966:PST

[Ran66] K. B. Ranger. The perturbation of some two dimensional viscous flows in a finite region. Journal of Mathematics and Physics (MIT), 45(1–4): 418–424, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451418.

Ranger:1967:IVF

[Ran67] K. B. Ranger. Incompressible viscous flow in a two dimensional channel under a transverse magnetic field. Journal of Mathematics and Physics (MIT), 46(1–4):21–27, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746121.

Ranger:1968:MFV

[Ran68] K. B. Ranger. Multipole flows of a viscous fluid exterior and interior to a circular cylinder. Journal of Mathematics and Physics (MIT), 47(1–4):456–460, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471456.

Rutledge:1930:TME

[RC30] George Rutledge and Prescott Crout. Tables and methods of extending tables for interpolation without differences. Journal of Mathematics and Physics (MIT), 9(1–4):166–180, April 1930. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm193091166.

Rutledge:1935:RVP

[RD35] George Rutledge and R. D. Douglass. The range of de la Vallée Poussin summation. Journal of Mathematics and Physics (MIT), 14(1–4):191– 194, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1935141191.

Redheffer:1949:DCA

[Red49a] R. M. Redheffer. Design of a circuit to approximate a prescribed amplitude and phase. Journal of Mathematics and Physics (MIT), 28(1–4):140–147, April 1949. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1949281140.

Redheffer:1949:RBN

[Red49b] R. M. Redheffer. Remarks on the basis of network theory. Journal of Mathematics and Physics (MIT), 28(1–4):237–258, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281237.

Redheffer:1960:CLF

[Red60] R. M. Redheffer. On a certain linear fractional transformation. Journal of Mathematics and Physics (MIT), 39(1–4):269–286, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1960391269.

Redheffer:1962:RTL

[Red62] Raymond Redheffer. On the relation of transmission-line theory to scattering and transfer. Journal of Mathematics and Physics (MIT), 41(1– 4):1–41, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19624111.

Redheffer:1964:ERP

[Red64] R. M. Redheffer. Elementary remarks on problems of mixed type. Journal of Mathematics and Physics (MIT), 43(1–4):1–14, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19644311.

Redheffer:1968:SGH

[Red68] Ray Redheffer. On Serrin’s generalization of Harnack’s inequality. Journal of Mathematics and Physics (MIT), 47(1–4):432–436, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1968471432.

Reissner:1937:RTB

[Rei37] M. Eric Reissner. Remark on the theory of bending of plates of variable thickness. Journal of Mathematics and Physics (MIT), 16(1–4):43– 45, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193716143.

Reissner:1941:CSI

[Rei41] Eric Reissner. On a class of singular integral equations. Journal of Mathematics and Physics (MIT), 20(1–4):219–223, April 1941. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1941201219.

Reissner:1944:NTS

[Rei44a] Eric Reissner. Note on the theorem of the symmetry of the stress tensor. Journal of Mathematics and Physics (MIT), 23(1–4):192–194, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1944231192.

Reissner:1944:TBE

[Rei44b] Eric Reissner. On the theory of bending of elastic plates. Journal of Mathematics and Physics (MIT), 23(1–4):184–191, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1944231184.

Reissner:1946:SSDa

[Rei46a] Eric Reissner. Stresses and small displacements of shallow spherical shells. I. Journal of Mathematics and Physics (MIT), 25(1–4):80–85, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194625180.

Reissner:1946:SSDb

[Rei46b] Eric Reissner. Stresses and small displacements of shallow spherical shells. II. Journal of Mathematics and Physics (MIT), 25(1–4):279– 300, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1946251279.

See corrections [Rei48a].

Reissner:1947:NMT

[Rei47] Eric Reissner. Note on the membrane theory of shell of revolution. Journal of Mathematics and Physics (MIT), 26(1–4):290–293, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261290.

Reissner:1948:CPS

[Rei48a] Eric Reissner. Corrections to the paper “Stresses and Small Displacements of Shallow Spherical Shells II”. Journal of Mathematics and Physics (MIT), 27(1–4):240, April 1948. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1948271240. See [Rei46b].

Reissner:1948:NMC

[Rei48b] Eric Reissner. Note on the method of complementary energy. Journal of Mathematics and Physics (MIT), 27(1–4):159–160, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1948271159.

Reissner:1950:VTE

[Rei50] Eric Reissner. On a variational theorem in elasticity. Journal of Mathematics and Physics (MIT), 29(1–4):90–95, April 1950. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195029190.

Reich:1951:DI

[Rei51] Edgar Reich. On the definition of information. Journal of Mathematics and Physics (MIT), 30(1–4):156–161, April 1951. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1951301156.

Reissner:1952:NUT

[Rei52a] Eric Reissner. On non-uniform torsion of cylindrical rods. Journal of Mathematics and Physics (MIT), 31(1–4):214–221, April 1952. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1952311214.

Reissner:1952:SSR

[Rei52b] Eric Reissner. Stress strain relations in the theory of thin elastic shells. Journal of Mathematics and Physics (MIT), 31(1–4):109–119, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311109.

Reissner:1953:VTF

[Rei53] Eric Reissner. On a variational theorem for finite elastic deformations. Journal of Mathematics and Physics (MIT), 32(1–4):129–135, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321129.

Reid:1955:DSS

[Rei55] Walter P. Reid. Distribution of sizes of spheres in a solid from a study of slices of the solid. Journal of Mathematics and Physics (MIT), 34(1–4): 95–102, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195534195.

Reissner:1959:DSD

[Rei59] Eric Reissner. On the determination of stresses and displacements for unsymmetrical deformations of shallow spherical shells. Journal of Mathematics and Physics (MIT), 38(1–4):16–35, April 1959. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195938116.

Reissner:1961:NFI

[Rei61] E. Reissner. Note on finite inextensional deformations of shallow elastic shells. Journal of Mathematics and Physics (MIT), 40(1–4):253– 259, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401253.

Reissner:1963:PBP

[Rei63a] Eric Reissner. On pure bending of pressurized toroidal membranes. Journal of Mathematics and Physics (MIT), 42(1–4):38–46, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196342138.

Reissner:1963:DTT

[Rei63b] Eric Reissner. On the derivation of the theory of thin elastic shells. Journal of Mathematics and Physics (MIT), 42(1–4):263–277, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1963421263.

Reynolds:1951:DPM

[Rey51] R. R. Reynolds. The Dirichlet problem for multiply connected domains. Journal of Mathematics and Physics (MIT), 30(1–4):11–22, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195130111.

Reyn:1967:SHR

[Rey67] J. W. Reyn. Solutions in the hyperbolic region of an equation, which approximates Chaplygin’s equation near the vacuum line. Journal of Mathematics and Physics (MIT), 46(1–4):28–42, April 1967. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196746128.

Reza:1954:CBC

[Rez54] F. M. Reza. Conversion of a brune cycle with an ideal transformer into a cycle without an ideal transformer. Journal of Mathematics and Physics (MIT), 33(1–4):194–198, April 1954. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1954331194.

Rhodes:1965:ESD

[Rho65a] Donald R. Rhodes. Errata: On some double orthogonality properties of the spheroidal and Mathieu functions. Journal of Mathematics and Physics (MIT), 44(1–4):411, April 1965. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1965441411. See [Rho65b].

Rhodes:1965:SDO

[Rho65b] Donald R. Rhodes. On some double orthogonality properties of the spheroidal and Mathieu functions. Journal of Mathematics and Physics (MIT), 44(1–4):52–65, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm196544152. See errata [Rho65a].

Rice:1922:CTP

[Ric22a] Lepine Hall Rice. A certain type of product and the combinatory analysis involved in its expansion. Journal of Mathematics and Physics (MIT), 1(2):86–88, March 1922. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19221286.

Rice:1922:EST

[Ric22b] Lepine Hall Rice. On the expression of the sum of any two determinants as a determinant of more dimensions. Journal of Mathematics and Physics (MIT), 1(3):160–166, April 1922. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192213160.

Rice:1924:CGD

[Ric24] Lepine Hall Rice. A contribution to the generalization of a determinantal theorem of Frobenius. Journal of Mathematics and Physics (MIT), 3(2):118–126, March 1924. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192432118.

Rice:1925:DEM

[Ric25a] Lepine Hall Rice. Determinantal expression of multiple cross and dot products of polyadics in three dimensions. Journal of Mathematics and Physics (MIT), 4(1–4):130–132, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192541130.

Rice:1925:FMO

[Ric25b] Lepine Hall Rice. File multiplication of ordered determinants. Journal of Mathematics and Physics (MIT), 4(1–4):200–204, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192541200.

Rice:1925:TED

[Ric25c] Lepine Hall Rice. A Taylor’s expansion of a determinant. Journal of Mathematics and Physics (MIT), 4(1–4):62–63, January 1925. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19254162.

Rice:1926:AID

[Ric26a] Lappine Hall Rice. Adjoint and inverse determinants and matrices. Journal of Mathematics and Physics (MIT), 5(1–4):55–64, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19265155.

Rice:1926:CSP

[Ric26b] Lepine Hall Rice. Compounds of Scott product determinants. Journal of Mathematics and Physics (MIT), 5(1–4):238–250, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192651238.

Rice:1927:CCP

[Ric27] Lepine Hall Rice. Compounds of Cayley products of determinants of higher class. Journal of Mathematics and Physics (MIT), 6(1–4):33– 38, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19276133.

Rice:1928:CRG

[Ric28] Lepine Hall Rice. Couche ranks in the general matrix. Journal of Mathematics and Physics (MIT), 7(1–4):93–96, April 1928. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19287193.

Rice:1929:DD

[Ric29] Lepine Hall Rice. Decomposition of determinants. Journal of Mathematics and Physics (MIT), 8(1–4):56–64, April 1929. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19298156.

Rice:1930:IHD

[Ric30] Lepine Hall Rice. Introduction to higher determinants. Journal of Mathematics and Physics (MIT), 9(1–4):47–71, April 1930. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19309147.

Rice:1932:RMV

[Ric32] Lepine Hall Rice. The rank of a matrix, the value of a determinant, and the solution of a system of linear equations. Journal of Mathematics and Physics (MIT), 11(1–4):146–149, April 1932. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1932111146.

Robbins:1936:TDC

[Rob36] Charles K. Robbins. Three-dimensional contact transformations in solved form. Journal of Mathematics and Physics (MIT), 15(1–4):268– 282, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151268.

Roberts:1948:NLT

[Rob48] Richard Calvin Roberts. Note on the lift of a triangular wing at supersonic speeds. Journal of Mathematics and Physics (MIT), 27(1–4): 49–55, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194827149.

Roberts:1957:NSE

[Rob57a] Leonard Roberts. On the numerical solution of the equations for spherical waves of finite amplitude, II. Journal of Mathematics and Physics (MIT), 36(1–4):329–337, April 1957. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1957361329.

Robinson:1957:WPH

[Rob57b] A. Robinson. Wave propagation in a heterogeneous elastic medium. Journal of Mathematics and Physics (MIT), 36(1–4):210–222, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361210.

Roettinger:1948:NUA

[Roe48] Ida Roettinger. Note on the use of almost periodic functions in the solution of certain boundary value problems. Journal of Mathematics and Physics (MIT), 27(1–4):232–239, April 1948. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1948271232.

Rosen:1930:NGL

[Ros30] N. Rosen. Note on the general Lorentz transformation. Journal of Mathematics and Physics (MIT), 9(1–4):181–187, April 1930. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193091181.

Rose:1958:EIT

[Ros58] M. E. Rose. The electrostatic interaction of two arbitrary charge distributions. Journal of Mathematics and Physics (MIT), 37(1–4):215– 222, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371215.

Rosenstock:1964:EMF

[Ros64] Herbert B. Rosenstock. Euler-MacLaurin formula in three dimensions. Journal of Mathematics and Physics (MIT), 43(1–4):342–346, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431342.

Rosen:1966:SCN

[Ros66a] Gerald Rosen. Solutions of a certain nonlinear wave equation. Journal of Mathematics and Physics (MIT), 45(1–4):235–265, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451235.

Ross:1966:TSA

[Ros66b] Edward W. Ross, Jr. Transition solutions for axisymmetric shell vibrations. Journal of Mathematics and Physics (MIT), 45(1–4):335– 355, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451335.

Rouse:1967:ASN

[Rou67] Carl A. Rouse. Accurate and stable numerical solutions to the Sehrödinger equation. Journal of Mathematics and Physics (MIT), 46(1–4):63–73, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196746163.

Riordan:1942:NTT

[RS42] John Riordan and C. E. Shannon. The number of two-terminal seriesparallel networks. Journal of Mathematics and Physics (MIT), 21(1–4): 83–93, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194221183.

Roth:1956:VMS

[RS56] J. P. Roth and D. S. Scott. A vector method for solving linear equations and inverting matrices. Journal of Mathematics and Physics (MIT), 35(1–4):312–317, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351312.

Reissner:1957:BIC

[RS57] Eric Reissner and M. B. Sledd. Bounds on influence coefficients for circular cylindrical shells. Journal of Mathematics and Physics (MIT), 36(1–4):1–19, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19573611.

Reissner:1966:ASB

[RS66a] E. Reissner and J. G. Simmonds. Asymptotic solutions of boundary value problems for elastic semi-infinite circular cylindrical shells. Journal of Mathematics and Physics (MIT), 45(1–4):1–22, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19664511.

Rosencrans:1966:SOO

[RS66b] S. I. Rosencrans and D. H. Sattinger. On the spectrum of an operator occurring in the theory of hydrodynamic stability. Journal of Mathematics and Physics (MIT), 45(1–4):289–300, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451289.

Reissner:1946:NSS

[RT46] E. Reissner and G. B. Thomas. Note on the shear stresses in a bent cantilever beam of rectangular cross section. Journal of Mathematics and Physics (MIT), 25(1–4):241–243, April 1946. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1946251241.

Rubin:1964:APS

[Rub64] Hanan Rubin. Anecdote on power series expansions of Mathieu functions. Journal of Mathematics and Physics (MIT), 43(1–4):339–341, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431339.

Russell:1933:MEF

[Rus33a] J. B. Russell. A method of evaluating the Fourier integral. Journal of Mathematics and Physics (MIT), 12(1–4):274–290, April 1933. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1933121274.

Russell:1933:THF

[Rus33b] J. B. Russell. A table of Hermite functions. Journal of Mathematics and Physics (MIT), 12(1–4):291–297, April 1933. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1933121291.

Rutledge:1922:EDC

[Rut22] George Rutledge. Explicit determination of Cotes’ coefficients for polynomial area. Journal of Mathematics and Physics (MIT), 1(2):78–85, March 1922. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19221278.

Rutledge:1923:PDP

[Rut23] George Rutledge. The polynomial determined by 2n + 1 points. Journal of Mathematics and Physics (MIT), 2(1–4):47–62, October 1923. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19232147.

Rutledge:1927:CIC

[Rut27] George Rutledge. Convergent interpolation coefficients with convergent sum. Journal of Mathematics and Physics (MIT), 6(1–4):240– 249, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192761240.

Rutledge:1929:FTL

[Rut29a] George Rutledge. Fundamental table for Lagrangean coefficients. Journal of Mathematics and Physics (MIT), 8(1–4):1–12, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1929811.

Rutledge:1929:LVL

[Rut29b] George Rutledge. Limiting values of Lagrangean coefficients. Journal of Mathematics and Physics (MIT), 8(1–4):13–17, April 1929. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19298113.

Rutledge:1930:NSC

[Rut30] George Rutledge. A necessary and sufficient condition for convergence of a sequence of Lagrange polynomials. Journal of Mathematics and Physics (MIT), 9(1–4):261–277, April 1930. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm193091261.

Rutledge:1932:IMV

[Rut32] George Rutledge. The inverse matrix for De la Vallée–Poussin summation. Journal of Mathematics and Physics (MIT), 11(1–4):73–82, April 1932. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193211173.

Robacker:1956:NSN

[RW56] John T. Robacker and Albert D. Wheelon. A note on the solution of a non-linear transonic flow equation. Journal of Mathematics and Physics (MIT), 35(1–4):321–322, April 1956. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1956351321.

Reissner:1968:AET

[RW68] E. Reissner and F. Y. M. Wan. On axial extension and torsion of helicoidal shells. Journal of Mathematics and Physics (MIT), 47(1–4): 1–31, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19684711.

Saibel:1928:ATT

[Sai28] Edward Saibel. Analytic and topological transformations of closed surfaces. Journal of Mathematics and Physics (MIT), 7(1–4):97–108, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19287197.

Salem:1942:SMF

[Sal42] R Salem. On singular monotonic functions of the Cantor type. Journal of Mathematics and Physics (MIT), 21(1–4):69–82, April 1942.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194221169.

Salzer:1943:CND

[Sal43a] Herbert E. Salzer. Coefficients for numerical differentiation with central differences. Journal of Mathematics and Physics (MIT), 22(1–4):115– 135, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1943221115.

Salzeri:1943:TCI

[Sal43b] Herbert E. Salzeri. Table of coefficients for inverse interpolation with central differences. Journal of Mathematics and Physics (MIT), 22(1–4):210–224, April 1943. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1943221210.

Salzer:1944:TCD

[Sal44a] Herbert E. Salzer. Table of coefficients for differences in terms of the derivatives. Journal of Mathematics and Physics (MIT), 23(1–4):210– 212, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1944231210.

Salzer:1944:TCI

[Sal44b] Herbert E. Salzer. Table of coefficients for inverse interpolation with advancing differences. Journal of Mathematics and Physics (MIT), 23(1–4):75–102, April 1944. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194423175.

Salzer:1945:FDI

[Sal45a] Herbert E. Salzer. Formulas for direct and inverse interpolation of a complex function tabulated along equidistant circular arcs. Journal of Mathematics and Physics (MIT), 24(1–4):141–143, April 1945. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1945241141.

Salzer:1945:IIE

[Sal45b] Herbert E. Salzer. Inverse interpolation for eight-, nine-, ten-, and eleven-point direct interpolation. Journal of Mathematics and Physics (MIT), 24(1–4):106–108, April 1945. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1945241106.

Salzer:1945:TCD

[Sal45c] Herbert E. Salzer. Table of coefficients for double quadrature without difference, for integrating second order differential equations. Journal of Mathematics and Physics (MIT), 24(1–4):135–140, April 1945. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1945241135.

Salzer:1946:CFU

[Sal46a] Herbert E. Salzer. Coefficients for facilitating the use of the Gaussian quadrature formula. Journal of Mathematics and Physics (MIT), 25(1–4):244–246, April 1946. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1946251244.

Salzer:1946:NCN

[Sal46b] Herbert E. Salzer. Note on coefficients for numerical integration with differences. Journal of Mathematics and Physics (MIT), 25(1–4):86– 88, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194625186.

Salzer:1947:AFD

[Sal47a] Herbert E. Salzer. Alternative formulas for direct interpolation of a complex function tabulated along equidistant circular arcs. Journal of Mathematics and Physics (MIT), 26(1–4):56–61, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194726156.

Salzer:1947:TFU

[Sal47b] Herbert E. Salzer. Tables for facilitating the use of Chebyshev’s quadrature formula. Journal of Mathematics and Physics (MIT), 26(1–4): 191–194, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261191.

Salzer:1947:TCI

[Sal47c] Herbert E. Salzer. Tables of coefficients for interpolating in functions of the variables. Journal of Mathematics and Physics (MIT), 26(1–4): 294–305, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1947261294.

Salzer:1948:CCQ

[Sal48a] Herbert E. Salzer. Coefficients for complex quartic, quintic, and sextic interpolation with a square grid. Journal of Mathematics and Physics (MIT), 27(1–4):136–156, April 1948. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1948271136.

Salzer:1948:CFT

[Sal48b] Herbert E. Salzer. Coefficients for facilitating trigonometric interpolation. Journal of Mathematics and Physics (MIT), 27(1–4):274–278, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1948271274.

Salzer:1949:CRI

[Sal49a] Herbert E. Salzer. Coefficients for repeated integration with central differences. Journal of Mathematics and Physics (MIT), 28(1–4):54– 61, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194928154.

Salzer:1949:FCC

[Sal49b] Herbert E. Salzer. Formulas for complex Cartesian interpolation of higher degree. Journal of Mathematics and Physics (MIT), 28(1–4): 200–204, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281200.

Salzer:1950:CPC

[Sal50a] Herbert E. Salzer. Coefficients for polar complex interpolation. Journal of Mathematics and Physics (MIT), 29(1–4):96–104, April 1950. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195029196.

Salzer:1950:FNI

[Sal50b] Herbert E. Salzer. Formulas for numerical integration of first and second order differential equations in the complex plane. Journal of Mathematics and Physics (MIT), 29(1–4):207–216, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1950291207.

Salzer:1952:FND

[Sal52] Herbert E. Salzer. Formulas for numerical differentiation in the complex plane. Journal of Mathematics and Physics (MIT), 31(1–4):155– 169, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311155.

Saltzer:1953:ABM

[Sal53a] Charles Saltzer. An abridged block method for the solution of the Dirichlet problem for the Laplace difference equation. Journal of Mathematics and Physics (MIT), 32(1–4):63–67, April 1953. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195332163.

Salzer:1953:RTO

[Sal53b] Herbert E. Salzer. Radix table for obtaining hyperbolic and inverse hyperbolic functions to many places. Journal of Mathematics and Physics (MIT), 32(1–4):197–202, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321197.

Salzer:1954:SMS

[Sal54] Hebert E. Salzer. A simple method for summing certain slowly convergent series. Journal of Mathematics and Physics (MIT), 33(1–4): 356–359, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331356.

Salzer:1955:EWQ

[Sal55a] Herbert E. Salzer. Equally weighted quadrature formulas over semiinfinite and infinite intervals. Journal of Mathematics and Physics (MIT), 34(1–4):54–63, April 1955. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195534154.

Salzer:1955:OQF

[Sal55b] Herbert E. Salzer. Osculatory quadrature formulas. Journal of Mathematics and Physics (MIT), 34(1–4):103–112, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341103.

Salzer:1956:CCO

[Sal56] Herbert E. Salzer. Coefficients for complex osculatory interpolation over a Cartesian grid. Journal of Mathematics and Physics (MIT), 35(1–4):152–163, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351152.

Salzer:1957:FCF

[Sal57] Herbert E. Salzer. Formulas for calculating Fourier coefficients. Journal of Mathematics and Physics (MIT), 36(1–4):96–98, April 1957.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195736196.

Salzer:1958:TNC

[Sal58] Herbert E. Salzer. Tables for the numerical calculation of inverse Laplace transforms. Journal of Mathematics and Physics (MIT), 37(1–4):89–109, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195837189.

Salzer:1960:FCC

[Sal60a] Herbert E. Salzer. Formulae for complex Cartesian hyperosculatory interpolation. Journal of Mathematics and Physics (MIT), 39(1–4): 300–307, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1960391300.

Salzer:1960:NFT

[Sal60b] Herbert E. Salzer. New formulas for trigonometric interpolation. Journal of Mathematics and Physics (MIT), 39(1–4):83–96, April 1960. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196039183.

Salzer:1961:AFT

[Sal61] Herbert E. Salzer. Additional formulas and tables for orthogonal polynomials originating from inversion integrals. Journal of Mathematics and Physics (MIT), 40(1–4):72–86, April 1961. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196140172.

Salzer:1962:MQB

[Sal62] Herbert E. Salzer. Multiple quadrature with backward differences. Journal of Mathematics and Physics (MIT), 41(1–4):219–227, April 1962. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1962411219.

Salzer:1963:NIE

[Sal63a] Herbert E. Salzer. Numerical integration employing “Overdifferentiation” — some simple experiments. Journal of Mathematics and Physics (MIT), 42(1–4):307–322, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421307.

Salzer:1963:TCO

[Sal63b] Herbert E. Salzer. Tables of coefficients for obtaining central differences from the derivatives. Journal of Mathematics and Physics (MIT), 42(1–4):162–165, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1963421162.

Salzer:1964:OTI

[Sal64] Herbert E. Salzer. Obreschkoff-type integration and interpolation formulas for periodic functions. Journal of Mathematics and Physics (MIT), 43(1–4):241–250, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1964431241.

Samuelsos:1945:CIP

[Sam45] Paul A. Samuelsos. A convergent iterative process. Journal of Mathematics and Physics (MIT), 24(1–4):131–134, April 1945. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1945241131.

Samuelson:1949:ICC

[Sam49] P. A. Samuelson. Iterative computation of complex roots. Journal of Mathematics and Physics (MIT), 28(1–4):259–267, April 1949. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281259.

Samuelson:1952:RCS

[Sam52] Paul A. Samuelson. Rapidly converging solutions to integral equations. Journal of Mathematics and Physics (MIT), 31(1–4):276–286, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311276.

Samoloff:1954:CSC

[Sam54] J. Samoloff. The convergence of the solutions of a class of iterative difference equations. Journal of Mathematics and Physics (MIT), 33(1–4): 105–110, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331105.

Sard:1956:AP

[Sar56] Arthur Sard. Approximation and projection. Journal of Mathematics and Physics (MIT), 35(1–4):127–144, April 1956. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1956351127.

Sattinger:1966:EIE

[Sat66] D. H. Sattinger. The eigenvalues of an integral equation in anistotropic neutron transport theory. Journal of Mathematics and Physics (MIT), 45(1–4):188–196, April 1966. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1966451188.

Savin:1937:SLH

[Sav37] S. A. Savin. Solution of Laplace’s harmonic equation with three independent variables. Journal of Mathematics and Physics (MIT), 16(1–4):196–201, April 1937. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1937161196.

Savin:1938:AFS

[Sav38] S A. Savin. Abridged formulas for the solution of differential equations of the theory of elasticity by means of integral polynomials. Journal of Mathematics and Physics (MIT), 17(1–4):201–207, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1938171201.

Savin:1939:FRC

[Sav39] S. A. Savin. Flexure of a rectangular covering slab elastically fixed by its four sides. Journal of Mathematics and Physics (MIT), 18(1–4): 239–255, April 1939. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1939181239.

Savin:1941:MIA

[Sav41a] S. A. Savin. Method of integral algebraical functions in the theory of elasticity of three dimensions. Journal of Mathematics and Physics (MIT), 20(1–4):1–17, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19412011.

Savin:1941:SSE

[Sav41b] S. A. Savin. On some solutions of the equations of internal equilibrium of the theory of elasticity. Journal of Mathematics and Physics (MIT), 20(1–4):23–29, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194120123.

Savin:1941:SVC

[Sav41c] S. A. Savin. Saint Venant’s conditions of compatibility in the method of integral algebraical functions. Journal of Mathematics and Physics (MIT), 20(1–4):18–22, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm194120118.

Shohat:1942:SAT

[SB42] J. A. Shohat and A. V. Bushkovitch. On some applications of the Tchebysheff inequality for definite integrals. Journal of Mathematics and Physics (MIT), 21(1–4):211–217, April 1942. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1942211211.

Schouten:1931:DEGa

[Sch31a] J. A. Schouten. Dirac equations in general relativity: 1. Four dimensional theory. Journal of Mathematics and Physics (MIT), 10(1–4): 239–271, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1931101239.

Schouten:1931:DEGb

[Sch31b] J. A. Schouten. Dirac equations in general relativity: 2. Five dimensional theory. Journal of Mathematics and Physics (MIT), 10(1–4): 272–283, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1931101272.

Schwartz:1941:GCR

[Sch41] Abraham Schwartz. The Gauss–Codazzi–Ricci equations in Riemannian manifolds. Journal of Mathematics and Physics (MIT), 20(1–4): 30–79, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194120130.

Scheffe:1942:LDE

[Sch42] Henry Scheffé. Linear differential equations with two-term recurrence formulas. Journal of Mathematics and Physics (MIT), 21(1–4):240– 249, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1942211240.

Schmidt:1955:TLX

[Sch55a] Paul W. Schmidt. Tables of ∫_{0}^{x} J_{0}(t) dt for large x. Journal of Mathematics and Physics (MIT), 34(1–4):169–172, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341169.

Schweizer:1955:SGL

[Sch55b] Berthold Schweizer. A symmetric generalization of the Lagrange interpolation formula. Journal of Mathematics and Physics (MIT), 34(1–4): 157–159, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1955341157.

Schweizer:1957:AEO

[Sch57] Berthold Schweizer. On approximate eigenvalues obtained by the method of least squares. Journal of Mathematics and Physics (MIT), 36(1–4):284–288, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361284.

Schoenfeld:1959:AE

[Sch59] Lowell Schoenfeld. The asymptotic expansion of. Journal of Mathematics and Physics (MIT), 38(1–4):193–202, April 1959. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1959381193.

Schweikert:1966:ICU

[Sch66] Daniel G. Schweikert. An interpolation curve using a spline in tension. Journal of Mathematics and Physics (MIT), 45(1–4):312–317, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451312.

Siegel:1952:ZPP

[SCKH52] K M. Siegel, J. W. Crispin, R. E. Kleinman, and H. E. Hunter. The zeros of P ′_{nι}(x_{0}) of non-integral degree. Journal of Mathematics and Physics (MIT), 31(1–4):170–179, April 1952. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1952311170.

Siegel:1953:NZ

[SCKH53] K. M. Siegel, J. W. Crispin, R. E. Kleinman, and H. E. Hunter. Note on the zeros of. Journal of Mathematics and Physics (MIT), 32(1–4): 193–196, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321193.

Scozzafava:1966:CZE

[Sco66] R. Scozzafava. On the complex zeros of an exponential polynomial occurring in the study of the stability of a vacuum diode. Journal of Mathematics and Physics (MIT), 45(1–4):104–108, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451104.

Samuels:1959:SLS

[SE59] J. Clifton Samuels and A. Cemal Eringen. On stochastic linear systems. Journal of Mathematics and Physics (MIT), 38(1–4):83–103, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195938183.

Sears:1949:LPT

[Sea49] W. R. Sears. The linear-perturbation theory for rotational flow. Journal of Mathematics and Physics (MIT), 28(1–4):268–271, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281268.

Segel:1960:UVA

[Seg60] Lee A. Segel. A uniformly-valid asymptotic expansion of the solution to an unsteady boundary-layer problem. Journal of Mathematics and Physics (MIT), 39(1–4):189–197, April 1960. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1960391189.

Serrin:1954:CTS

[Ser54] J. B. Serrin. Comparison theorems for subsonic flows. Journal of Mathematics and Physics (MIT), 33(1–4):27–45, April 1954. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195433127.

Seth:1952:FEP

[Set52] B. R. Seth. Finite elastic-plastic torsion. Journal of Mathematics and Physics (MIT), 31(1–4):84–90, April 1952. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195231184.

Shannon:1941:MTD

[Sha41a] Claude E. Shannon. Mathematical theory of the differential analyzer. Journal of Mathematics and Physics (MIT), 20(1–4):337–354, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1941201337.

Sharp:1941:CME

[Sha41b] Henry S. Sharp. A comparison of methods for evaluating the complex roots of quartic equations. Journal of Mathematics and Physics (MIT), 20(1–4):243–258, April 1941. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1941201243.

Shannon:1949:TCL

[Sha49] Claude E. Shannon. A theorem on coloring the lines of a network. Journal of Mathematics and Physics (MIT), 28(1–4):148–152, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281148.

Shanks:1955:NLT

[Sha55] Daniel Shanks. Non-linear transformations of divergent and slowly convergent sequences. Journal of Mathematics and Physics (MIT), 34(1– 4):1–42, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19553411.

Sharp:1968:NR

[Sha68] R. T. Sharp. Notes on O(4) representations. Journal of Mathematics and Physics (MIT), 47(1–4):359–365, April 1968. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1968471359.

Shen:1948:HFS

[She48] S. F. Shen. Hypersonic flow over a slender cone. Journal of Mathematics and Physics (MIT), 27(1–4):56–66, April 1948. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194827156.

Shen:1952:NDS

[She52] S. F. Shen. Note on the distribution of singular directions of a two-dimensional shock. Journal of Mathematics and Physics (MIT), 31(1–4):102–108, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1952311102.

Shield:1954:SVF

[Shi54] R. T. Shield. Stress and velocity fields in soil mechanics. Journal of Mathematics and Physics (MIT), 33(1–4):144–156, April 1954. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1954331144.

Shinbrot:1957:IEO

[Shi57] Marvin Shinbrot. On the integral equation occurring in optimization theory with nonstationary inputs. Journal of Mathematics and Physics (MIT), 36(1–4):121–129, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361121.

Siegel:1955:BLF

[Sie55] Keeve M. Siegel. Bounds of the Legendre function. Journal of Mathematics and Physics (MIT), 34(1–4):43–49, April 1955. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195534143.

Simons:1956:PSS

[Sim56] Roger M. Simons. A power series solution of the nonlinear equations for axi-symmetrical bending of shallow spherical shells. Journal of Mathematics and Physics (MIT), 35(1–4):164–176, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1956351164.

Simmonds:1966:ICS

[Sim66] James G. Simmonds. Influence coefficients for semi-infinite and infinite circular cylindrical elastic shells. Journal of Mathematics and Physics (MIT), 45(1–4):127–149, April 1966. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1966451127.

Sternberg:1952:ATS

[SK52] R. L. Sternberg and H. Kauffman. Applications of the theory of systems of differential equations to multiple non-uniform transmission lines. Journal of Mathematics and Physics (MIT), 31(1–4):244–252, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311244.

Sternberg:1953:GST

[SK53] R. L. Sternberg and H. Kaufman. A general solution of the twofrequency modulation product problem. I. Journal of Mathematics and Physics (MIT), 32(1–4):233–242, April 1953. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1953321233.

Stulen:1956:MSI

[SL56] F. B. Stulen and F. G. Lehman. A method of solving inhomogeneous linear simultaneous equations. Journal of Mathematics and Physics (MIT), 35(1–4):123–126, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351123.

Sleatoh:1960:VSP

[Sle60] F. B. Sleatoh. A variational solution to the problem of scalar scattering by a prolate spheroid. Journal of Mathematics and Physics (MIT), 39(1–4):105–120, April 1960. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1960391105.

Slepian:1965:SAE

[Sle65] David Slepian. Some asymptotic expansions for prolate spheroidal wave functions. Journal of Mathematics and Physics (MIT), 44(1–4):99– 140, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196544199.

Smith:1943:AE

[Smi43] V. G. Smith. An asymptotic expansion of J_{i_{0}(x)} = ∫_{x}^{∞} J_{0}(t)/t dt. Journal of Mathematics and Physics (MIT), 22(1–4):58–59, April 1943. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194322158.

Smith:1957:TIP

[Smi57] R. C. T. Smith. Tension of an infinite plate cut along a circular arc. Journal of Mathematics and Physics (MIT), 36(1–4):223–233, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361223.

Smith:1968:SIH

[Smi68] J. H. Smith. Solution of the inhomogeneous Helmholtz equation with Lebedev transforms. Journal of Mathematics and Physics (MIT), 47(1–4):442–450, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471442.

Riley:1962:AER

[sNR62] s N. Riley. Asymptotic expansions in radial jets. Journal of Mathematics and Physics (MIT), 41(1–4):132–146, April 1962. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1962411132.

Soh:1933:TGE

[Soh33] Hsin P. Soh. Theory of gravitation and electromagnetism. Journal of Mathematics and Physics (MIT), 12(1–4):298–305, April 1933. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1933121298.

Soni:1965:IMB

[Son65] R. P. Soni. On an inequality for modified Bessel functions. Journal of Mathematics and Physics (MIT), 44(1–4):406–407, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441406.

Spencer:1941:TR

[Spe41a] D. C. Spencer. On a theorem of rengels. Journal of Mathematics and Physics (MIT), 20(1–4):118–123, April 1941. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1941201118.

Spencer:1941:DAT

[Spe41b] D. C. Spencer. On distortion in analytic transformations. Journal of Mathematics and Physics (MIT), 20(1–4):124–126, April 1941. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1941201124.

Spencer:1942:LPT

[Spe42a] D. C. Spencer. The lattice points of tetrahedra. Journal of Mathematics and Physics (MIT), 21(1–4):189–197, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1942211189.

Spencer:1942:NMO

[Spe42b] D. C. Spencer. Note on mean one-valent functions. Journal of Mathematics and Physics (MIT), 21(1–4):178–188, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1942211178.

Spencer:1942:SRC

[Spe42c] D C. Spencer. Some remarks concerning the coefficients of Schlicht functions. Journal of Mathematics and Physics (MIT), 21(1–4):63– 68, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194221163.

Spencer:1944:GFA

[Spe44a] Domina Eberle Spencer. Geometric figures in affine space. Journal of Mathematics and Physics (MIT), 23(1–4):1–23, April 1944. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19442311.

Spencer:1944:TRF

[Spe44b] Domina Eberle Spencer. The tensor representation of the figures of Study’s “Geometrie der Dynamen”. Journal of Mathematics and Physics (MIT), 23(1–4):103–115, April 1944. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1944231103.

Spiegel:1957:DDF

[Spi57] Murrary R. Spiegel. The Dirac delta function and the summation of certain Bessel series. Journal of Mathematics and Physics (MIT), 36(1–4): 378–380, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361378.

Spillers:1964:SFP

[Spi64] William R. Spillers. A shrink fit problem. Journal of Mathematics and Physics (MIT), 43(1–4):65–71, April 1964. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196443165.

Sponsler:1955:TFI

[Spo55] George C. Sponsler. Two formulae involving generalized Legendre functions of integral and nonintegral index. Journal of Mathematics and Physics (MIT), 34(1–4):50–53, April 1955. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195534150.

Steinberg:1952:STA

[SR52] R. Steinberg and R. M. Redheffer. Simultaneous trigonometric approximation. Journal of Mathematics and Physics (MIT), 31(1–4):

260–266, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311260.

Shirtliffe:1961:COA

[SS61] C. J. Shirtliffe and D. G. Stephenson. A computer oriented adaption of Salzer’s method for inverting Laplace transforms. Journal of Mathematics and Physics (MIT), 40(1–4):135–141, April 1961. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1961401135.

Schweizer:1963:ICH

[SS63] B. Schweizer and A. Sklar. Inequalities for the confluent hypergeometric function. Journal of Mathematics and Physics (MIT), 42(1–4):329– 330, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1963421329.

Schroeder:1967:GVT

[SS67] J. Schroeder and A. N. Sherbourne. A general variational theorem in classical plasticity. Journal of Mathematics and Physics (MIT), 46(1–4):175–187, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461175.

Schroeder:1968:BUY

[SS68a] J. Schroeder and A. N. Sherbourne. Bounds for unsymmetrical yield point loads of thin shells. Journal of Mathematics and Physics (MIT), 47(1–4):249–261, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471249.

Schroeder:1968:GTT

[SS68b] J. Schroeder and A. N. Sherbourne. A general theorem for thin shells in classical plasticity. Journal of Mathematics and Physics (MIT), 47(1–4):95–108, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196847195.

Tsien:1946:ODF

[sT46] Hsue shen Tsien. One-dimensional flows of a gas characterized by van der Waal’s equation of state. Journal of Mathematics and Physics (MIT), 25(1–4):301–324, April 1946. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1946251301. See correction [Tsi47].

Tsien:1947:FCN

[sT47] Hsue shen Tsien. Flow conditions near the intersection of a shock wave with solid boundary. Journal of Mathematics and Physics (MIT), 26(1– 4):69–75, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm194726169.

Staniland:1936:AAT

[Sta36a] A. E. Staniland. Analytic affine transformations in Euclidean space of 2n dimensions. Journal of Mathematics and Physics (MIT), 15(1–4): 283–290, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1936151283.

Staniland:1936:NAR

[Sta36b] A. E. Staniland. A note on analytic rotations in Euclidean space of four dimensions. Journal of Mathematics and Physics (MIT), 15(1–4): 65–72, April 1936. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm193615165.

Sternberg:1954:GSTa

[Ste54a] Robert L. Sternberg. A general solution of the two-frequency modulation product problem II. Tables of the functions A_{mn}(h, k). Journal of Mathematics and Physics (MIT), 33(1–4):68–79, April 1954. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm195433168.

Sternberg:1954:GSTb

[Ste54b] Robert L. Sternberg. A general solution of the two-frequency modulation product problem III. Rectifiers and limiters. Journal of Mathematics and Physics (MIT), 33(1–4):199–205, April 1954. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1954331199.

Sternberg:1955:SAE

[Ste55] Robert L. Sternberg. Successive approximation and expansion methods in the numerical design of microwave dielectric lenses. Journal of Mathematics and Physics (MIT), 34(1–4):209–235, April 1955. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1955341209.

Stewartson:1957:AET

[Ste57] K. Stewartson. On asymptotic expansions in the theory of boundary layers. Journal of Mathematics and Physics (MIT), 36(1–4):173– 191, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1957361173.

See note [Cra59].

Stern:1968:TGM

[Ste68] Melvin E. Stern. Towards a generalization of “Minimum free-energy” applicable to statistically steady turbulent motion. Journal of Mathematics and Physics (MIT), 47(1–4):416–428, April 1968. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1968471416.

Stickforth:1965:DCC

[Sti65] J. Stickforth. On the derivation of the conditions of compatibility from Castigliano’s principle by means of three-dimensional stress functions. Journal of Mathematics and Physics (MIT), 44(1–4):214–226, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441214.

Stone:1948:SFP

[Sto48] A. H. Stone. On supersonic flow past a slightly yawing cone. Journal of Mathematics and Physics (MIT), 27(1–4):67–81, April 1948. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194827167.

Stone:1951:SFP

[Sto51] A. H. Stone. On supersonic flow past a slightly yawing cone II. Journal of Mathematics and Physics (MIT), 30(1–4):200–213, April 1951. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1951301200. See corrections

[Sto52].

Stone:1952:CPS

[Sto52] A. H. Stone. Corrections to the paper, “On Supersonic Flow Past a Slightly Yawing Cone II”. Journal of Mathematics and Physics (MIT), 31(1–4):300, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1952311300a. See [Sto51].

Struik:1928:SPD

[Str28a] D. J. Struik. On sets of principal directions in a Riemannian manifold of four dimensions. Journal of Mathematics and Physics (MIT), 7(1–4): 193–197, April 1928. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192871193.

Struik:1928:RTQ

[Str28b] D. J. Struik. A relativistic theory of Quanta. Journal of Mathematics and Physics (MIT), 7(1–4):1–23, April 1928. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1928711.

Struik:1930:DPD

[Str30] D. J. Struik. On the derivation of Poisson’s distribution law. Journal of Mathematics and Physics (MIT), 9(1–4):151–161, April 1930. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm193091151.

Struik:1932:CM

[Str32] D. J. Struik. C. L. E. Moore. Journal of Mathematics and Physics (MIT), 11(1–4):1–11, April 1932. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19321111.

Struik:1935:FPT

[Str35] D. J. Struik. Five papers on the theory of probability. Journal of Mathematics and Physics (MIT), 14(1–4):1–3, April 1935. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19351411.

Strang:1959:OCC

[Str59] W. Gilbert Strang. On the order of convergence of the Crank–Nicolson procedure. Journal of Mathematics and Physics (MIT), 38(1–4):141– 144, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381141.

Strang:1962:TPD

[Str62a] Gilbert Strang. Trigonometric polynomials and difference methods of maximum accuracy. Journal of Mathematics and Physics (MIT), 41(1–4):147–154, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411147.

Stroh:1962:SSP

[Str62b] A. N. Stroh. Steady state problems in anisotropic elasticity. Journal of Mathematics and Physics (MIT), 41(1–4):77–103, April 1962. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196241177.

Streifer:1965:ZBF

[Str65] William Streifer. On the zeros of Bessel functions as a function of order. Journal of Mathematics and Physics (MIT), 44(1–4):403–405, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441403.

Streifer:1968:UAE

[Str68] William Streifer. Uniform asymptotic expansions for prolate spheroidal wave functions. Journal of Mathematics and Physics (MIT), 47(1–4): 407–415, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471407.

Su:1959:EAT

[Su59] Hsuan-Loh Su. Electric analog for theory of adjustment and regression. Journal of Mathematics and Physics (MIT), 38(1–4):312–326, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381312.

Sutton:1937:IAD

[Sut37] Charles S. Sutton. An investigation of the average distribution of twin prime numbers. Journal of Mathematics and Physics (MIT), 16(1–4): 1–42, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19371611.

Smith:1966:SPT

[SW66] T. E. Smith and W. E. Warren. Some problems in two-dimensional electrostriction. Journal of Mathematics and Physics (MIT), 45(1–4): 45–51, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196645145. See

corrigenda [SW68].

Smith:1968:CSP

[SW68] Thomas E. Smith and William E. Warren. Corrigenda to “Some Problems in Two-Dimensional Electrostriction”. Journal of Mathematics and Physics (MIT), 47(1–4):109–110, April 1968. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1968471109. See [SW66].

Synge:1936:SQV

[Syn36] J. L. Synge. The stability of quadratic velocity-distributions for an inviscid liquid flowing between parallel planes. Journal of Mathematics and Physics (MIT), 15(1–4):205–210, April 1936. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1936151205.

Synge:1956:EWA

[Syn56] J. L. Synge. Elastic waves in anisotropic media. Journal of Mathematics and Physics (MIT), 35(1–4):323–334, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1956351323.

Szasz:1949:SSC

[Szá49] Otto Szász. Summation of slowly convergent series. Journal of Mathematics and Physics (MIT), 28(1–4):272–279, April 1949. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1949281272.

Trent:1963:GAL

[TA63] H. M. Trent and L. Auslander. Generalized automorphisms of labelled graphs and graphical invariants. Journal of Mathematics and Physics (MIT), 42(1–4):78–84, April 1963. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196342178.

Tao:1959:TS

[Tao59] L. N. Tao. On toroidal shells. Journal of Mathematics and Physics (MIT), 38(1–4):130–134, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1959381130.

Taylor:1923:FSR

[Tay23] J. S. Taylor. A four-space representation of complex plane analytics. Journal of Mathematics and Physics (MIT), 2(1–4):1–30, October 1923. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1923211.

Taylor:1924:TTF

[Tay24] J. S. Taylor. The theory of testimony: First paper: a single witness. Journal of Mathematics and Physics (MIT), 3(3):162–173, April 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192433162.

Taylor:1939:CSA

[Tay39] W. C. Taylor. A complete set of asymptotic formulas for the Whittaker function and the LaGuerre polynomials. Journal of Mathematics and Physics (MIT), 18(1–4):34–49, April 1939. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm193918134.

Tea:1945:GMN

[Tea45] Peter L. Tea. A graphical method for the numerical solution of Fredholm’s integral equation of the second kind. Journal of Mathematics and Physics (MIT), 24(1–4):109–125, April 1945. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1945241109.

Tellegen:1953:SNC

[Tel53] B. D. H. Tellegen. Synthesis of 2n by networks containing the minimum number of elements. Journal of Mathematics and Physics (MIT), 32(1– 4):1–18, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19533211.

Terzaghi:1929:SBD

[Ter29] Charles Terzaghi. Settlement of buildings due to progressive consolidation of individual strata. Journal of Mathematics and Physics (MIT), 8(1–4):266–292, April 1929. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192981266.

Tomotika:1950:TFC

[TH50] S. Tomotika and Z. Hasimoto. On the transonic flow of a compressible fluid through an axially symmetrical nozzle. Journal of Mathematics and Physics (MIT), 29(1–4):105–117, April 1950. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1950291105.

Thomas:1940:CFR

[Tho40] T. Y. Thomas. The characterization of flat Riemann spaces as a boundary value problem. Journal of Mathematics and Physics (MIT), 19(1–4):247–259, April 1940. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1940191247.

Thomas:1944:SMS

[Tho44] T. Y. Thomas. Surfaces of maximum shearing stress. Journal of Mathematics and Physics (MIT), 23(1–4):167–172, April 1944. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1944231167.

Thomas:1946:TED

[Tho46a] T. Y. Thomas. On the transformation of the equations of dynamics. Journal of Mathematics and Physics (MIT), 25(1–4):191–208, April 1946. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1946251191.

Thomas:1946:RDS

[Tho46b] T. Y. Thomas. Reducible dynamical systems. Journal of Mathematics and Physics (MIT), 25(1–4):89–91, April 1946. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm194625189.

Thomas:1947:CSW

[Tho47] T. Y. Thomas. On curved shock waves. Journal of Mathematics and Physics (MIT), 26(1–4):62–68, April 1947. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm194726162.

Thomas:1948:CCA

[Tho48] T. Y. Thomas. Calculation of curvatures of attached shock waves. Journal of Mathematics and Physics (MIT), 27(1–4):279–297, April 1948. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1948271279.

Thomas:1949:CRS

[Tho49a] T. Y. Thomas. The consistency relations for shock waves. Journal of Mathematics and Physics (MIT), 28(1–4):62–90, April 1949. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194928162.

Thomas:1949:DSS

[Tho49b] T. Y. Thomas. The distribution of singular shock directions. Journal of Mathematics and Physics (MIT), 28(1–4):153–172, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1949281153.

Thomas:1949:CSP

[Tho49c] T. Y. Thomas. On conditions for steady plane flow with shock waves. Journal of Mathematics and Physics (MIT), 28(1–4):91–98, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194928191.

Thomas:1956:DES

[Tho56] T. Y. Thomas. Deformation energy and the stress-strain relations for isotropic materials. Journal of Mathematics and Physics (MIT), 35(1–4):335–350, April 1956. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1956351335.

Tiem:1968:CFG

[Tie68] Dang Hong Tiem. Canonical forms of gravitational field equations in empty space. Journal of Mathematics and Physics (MIT), 47(1–4): 461–464, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471461.

Ting:1953:DRW

[Tin53] Lu Ting. Diffraction and reflection of weak shocks by structures. Journal of Mathematics and Physics (MIT), 32(1–4):102–116, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1953321102.

Ting:1959:MTP

[Tin59] Lu Ting. On the mixing of two parallel streams. Journal of Mathematics and Physics (MIT), 38(1–4):153–165, April 1959. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1959381153.

Ting:1965:ICB

[Tin65] Lu Ting. On the initial conditions for boundary layer equations. Journal of Mathematics and Physics (MIT), 44(1–4):353–367, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1965441353.

Taylor:1940:TCC

[TM40] Donald W. Taylor and Wilfred Merchant. A theory of clay consolidation accounting for secondary compression. Journal of Mathematics and Physics (MIT), 19(1–4):167–185, April 1940. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1940191167.

Todd:1933:CP

[Tod33] J. A. Todd. A combinatorial problem. Journal of Mathematics and Physics (MIT), 12(1–4):321–333, April 1933. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1933121321.

Torda:1952:BLC

[Tor52] T. Paul Torda. Boundary layer control by continuous surface suction or injection. Journal of Mathematics and Physics (MIT), 31(1–4):206– 213, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1952311206.

Torda:1953:BLC

[Tor53] T. P. Torda. Boundary layer control by distributed surface suction or injection. Bi-parametric general solution. Journal of Mathematics and Physics (MIT), 32(1–4):312–314, April 1953. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1953321312.

Ting:1967:NMT

[TR67] Lu Ting and Charles J. Ruger. A note on the mixing of two parallel subsonic streams. Journal of Mathematics and Physics (MIT), 46(1–4): 343–348, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1967461343.

Trifan:1956:STP

[Tri56] D. Trifan. Stress theory of plastic flow. Journal of Mathematics and Physics (MIT), 35(1–4):44–52, April 1956. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195635144.

Tsang:1954:END

[Tsa54] N. F. Tsang. On electrical network determinants. Journal of Mathematics and Physics (MIT), 33(1–4):185–193, April 1954. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1954331185.

Tsien:1946:SLH

[Tsi46] Hsue-Shen Tsien. Similarity laws of hypersonic flows. Journal of Mathematics and Physics (MIT), 25(1–4):247–251, April 1946. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1946251247.

Tsien:1947:CPO

[Tsi47] Hsue-Shen Tsien. Corrections on the paper “One-Dimensional Flows of a Gas Characterized by van der Waal’s Equation of State”. Journal of Mathematics and Physics (MIT), 26(1–4):76–77, April 1947. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194726176. See [sT46].

Tsu:1961:EIN

[Tsu61] Raphael Tsu. The evaluation of incomplete normalization integrals and derivatives with respect to the order of associated Legendre polynomials. Journal of Mathematics and Physics (MIT), 40(1–4):232– 243, April 1961. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1961401232.

Tucker:1966:RCP

[Tuc66] Richard R. Tucker. Remark concerning a paper by Imanuel Marx. Journal of Mathematics and Physics (MIT), 45(1–4):233–234, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451233. See [Mar63b].

Turner:1949:ATO

[Tur49] M. J. Turner. Aerodynamic theory of oscillating sweptback wings. Journal of Mathematics and Physics (MIT), 28(1–4):280–293, April 1949. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1949281280.

Trilling:1953:TFP

[TW53] L. Trilling and K. Walker, Jr. On the transonic flow past a finite wedge. Journal of Mathematics and Physics (MIT), 32(1–4):72–79, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195332172.

Wu:1956:FST

[tW56a] T. Yao tsu Wu. A free streamline theory for two-dimensional fully cavitated hydrofoils. Journal of Mathematics and Physics (MIT), 35(1–4): 236–265, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351236.

Wu:1956:SPU

[tW56b] T. Yao tsu Wu. Small perturbations in the unsteady flow of a compressible viscous and heat-conducting fluid. Journal of Mathematics and Physics (MIT), 35(1–4):13–27, April 1956. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm195635113.

Tomotika:1956:FVL

[TY56] S. Tomotika and H. Yosinobu. The flow of a viscous liquid past a flat plate at small Reynolds numbers. Journal of Mathematics and Physics (MIT), 35(1–4):1–12, April 1956. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19563511.

Tomotika:1957:CHC

[TY57] S. Tomotika and H. Yosinobu. On the convection of heat from cylinders immersed in a low-speed stream of incompressible fluid. Journal of Mathematics and Physics (MIT), 36(1–4):112–120, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1957361112.

Uhlenbeck:1935:PPC

[Uhl35] G. E. Uhlenbeck. The probability of position in a canonical ensemble. Journal of Mathematics and Physics (MIT), 14(1–4):10–16, April 1935. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm193514110.

Vallarta:1924:NQN

[Val24a] M. S. Vallarta. Note on the quantization of non-conditioned-periodic systems. Journal of Mathematics and Physics (MIT), 3(2):108–117, March 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192432108.

Vallarta:1924:NDS

[Val24b] M. Sandoval Vallarta. Notes on dynamical systems non-integrable by separation of variables and on the existence of “Unmechanical” orbits in the atom. Journal of Mathematics and Physics (MIT), 3(3):174– 181, April 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192433174.

Vallarta:1925:STF

[Val25] M. Sandoval Vallarta. Sommerfeld’s theory of fine structure from the standpoint of general relativity. Journal of Mathematics and Physics (MIT), 4(1–4):65–83, January 1925. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm19254165.

Vallarta:1926:TCX

[Val26] M. Sandoval Vallarta. Theory of the continuous X-ray spectrum. Journal of Mathematics and Physics (MIT), 5(1–4):1–7, January 1926.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1926511.

Vallarta:1927:CVM

[Val27] M. S. Vallarta. On the conditions of validity of macromechanics. Journal of Mathematics and Physics (MIT), 6(1–4):209–222, April 1927. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192761209.

Vallarta:1929:NSI

[Val29] M. S. Vallarta. Note on the statistical interpretation of Maxwell’s equations. Journal of Mathematics and Physics (MIT), 8(1–4):155– 161, April 1929. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192981155.

Valanis:1965:PAW

[Val65] K. C. Valanis. Propagation and attenuation of waves in linear viscoelastic solids. Journal of Mathematics and Physics (MIT), 44(1–4): 227–239, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441227.

Valanis:1966:TLV

[Val66] K. C. Valanis. Thermodynamics of large viscoelastic deformations. Journal of Mathematics and Physics (MIT), 45(1–4):197–212, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451197.

Valanis:1967:EFM

[Val67] K. C. Valanis. Entropy, fading memory and Onsager’s relations. Journal of Mathematics and Physics (MIT), 46(1–4):164–174, April 1967. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1967461164.

Valanis:1968:VPT

[Val68] K. C. Valanis. The visoelastic potential and its thermodynamic foundations. Journal of Mathematics and Physics (MIT), 47(1–4):262–275, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471262.

VanNostrand:1954:OHF

[Van54] Robert G. Van Nostrand. The orthogonality of the hyperboloid functions. Journal of Mathematics and Physics (MIT), 33(1–4):276–282, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331276.

VanDyke:1958:PRS

[Van58] M. D. Van Dyke. The paraboloid of revolution in subsonic flow. Journal of Mathematics and Physics (MIT), 37(1–4):38–51, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195837138.

Varga:1961:HOS

[Var61] Richard S. Varga. On higher order stable implicit methods for solving parabolic partial differential equations. Journal of Mathematics and Physics (MIT), 40(1–4):220–231, April 1961. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1961401220.

vonkarman:1947:SLT

[vk47] Theodore von karman. The similarity law of transonic flow. Journal of Mathematics and Physics (MIT), 26(1–4):182–190, April 1947. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1947261182.

Wadsworth:1932:STL

[Wad32] G. P. Wadsworth. Systems of three linear partial differential equations of the third order. Journal of Mathematics and Physics (MIT), 11(1–4): 255–272, April 1932. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1932111255.

Wadsworth:1935:GLH

[Wad35] G. P. Wadsworth. The geometry of linear homogeneous partial differential equations of the first order. Journal of Mathematics and Physics (MIT), 14(1–4):279–290, April 1935. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1935141279.

Wagner:1951:NEF

[Wag51a] Carl Wagner. On the numerical evaluation of Fredholm integral equations with the aid of the Liouville–Neumann series. Journal of Mathematics and Physics (MIT), 30(1–4):232–234, April 1951. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1951301232. See erratum [Wag52].

Wagner:1951:SFI

[Wag51b] Carl Wagner. On the solution of Fredholm integral equations of second kind by iteration. Journal of Mathematics and Physics (MIT), 30(1–4):23–30, April 1951. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm195130123.

Wagner:1952:ENE

[Wag52] Carl Wagner. Erratum: On the Numerical Evaluation of Fredholm Integral Equations with the Aid of the Liouville-Neumann Series. Journal of Mathematics and Physics (MIT), 31(1–4):300, April 1952. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1952311300b. See [Wag51a].

Wagner:1953:NSV

[Wag53] Carl Wagner. On the numerical solution of Volterra integral equations. Journal of Mathematics and Physics (MIT), 32(1–4):289–301, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321289.

Walker:1938:NDC

[Wal38] Laurence R. Walker. A note on the determinant of certain matrices. Journal of Mathematics and Physics (MIT), 17(1–4):1–4, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19381711.

Waligorski:1959:TG

[Wal59] S. Waligórski. On a theorem of E. N. Gilbert. Journal of Mathematics and Physics (MIT), 38(1–4):206–207, April 1959. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1959381206.

Wang:1967:FPC

[Wan67] Chang-Yi Wang. The flow past a circular cylinder which is started impulsively from rest. Journal of Mathematics and Physics (MIT), 46(1–4):195–202, April 1967. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1967461195.

Wan:1968:TVT

[Wan68a] Frederic Y. M. Wan. Two variational theorems for thin shells. Journal of Mathematics and Physics (MIT), 47(1–4):429–431, April 1968.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1968471429.

Wang:1968:NDI

[Wan68b] Chang-Yi Wang. A note on the drag of an impulsively started circular cylinder. Journal of Mathematics and Physics (MIT), 47(1–4):451– 455, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471451.

Wang:1968:NAS

[Wan68c] Kuo Chang Wang. A new approach to ‘not-so-slender’ wing theory. Journal of Mathematics and Physics (MIT), 47(1–4):391–406, April 1968. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1968471391.

Wang:1968:EPT

[Wan68d] Neng-Ming Wang. Effects of plate thickness on the bending of an elastic plate containing a crack. Journal of Mathematics and Physics (MIT), 47(1–4):371–390, April 1968. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1968471371.

Warga:1952:CIP

[War52] J. Warga. On a class of iterative procedures for solving normal systems of ordinary differential equations. Journal of Mathematics and Physics (MIT), 31(1–4):223–243, April 1952. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/

sapm1952311223. See note [War53].

Warga:1953:NPC

[War53] J. Warga. A note on the paper “On a Class of Iterative Procedures for Solving Normal Systems of Ordinary Differential Equations”. Journal of Mathematics and Physics (MIT), 32(1–4):315, April 1953. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1953321315. See [War52].

Wasow:1944:ASB

[Was44] Wolfgang Wasow. On the asymptotic solution of boundary value problems for ordinary differential equations containing a parameter. Journal of Mathematics and Physics (MIT), 23(1–4):173–183, April 1944. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1944231173.

Washizu:1953:BSB

[Was53] K. Washizu. Bounds for solutions of boundary value problems in elasticity. Journal of Mathematics and Physics (MIT), 32(1–4):117–128, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1953321117.

Washizu:1957:NCC

[Was57] K. Washizu. A note on the conditions of compatibility. Journal of Mathematics and Physics (MIT), 36(1–4):306–312, April 1957. CO-DEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1957361306.

Wasow:1959:TPP

[Was59] Wolfgang Wasow. A turning point problem for a system of two linear differential equations. Journal of Mathematics and Physics (MIT), 38(1–4):257–278, April 1959. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1959381257.

Washizu:1964:SCN

[Was64] K. Washizu. Some considerations on a naturally curved and twisted slender beam. Journal of Mathematics and Physics (MIT), 43(1–4): 111–116, April 1964. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1964431111.

Waterman:1963:RLF

[Wat63] P. C. Waterman. Roots of Legendre functions of variable index. Journal of Mathematics and Physics (MIT), 42(1–4):323–328, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1963421323.

Walltman:1965:CSF

[WB65] Paul Walltman and T. F. Bridgland, Jr. On convergence of solutions of the forced Liénard equation. Journal of Mathematics and Physics (MIT), 44(1–4):284–287, April 1965. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1965441284.

Weaver:1932:UPD

[Wea32] Warren Weaver. Uplift pressure on dams. Journal of Mathematics and Physics (MIT), 11(1–4):114–145, April 1932. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1932111114.

Weeks:1934:SSC

[Wee34a] Dorothy W. Weeks. A study of sixteen coherency matrices. Journal of Mathematics and Physics (MIT), 13(1–4):380–386, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1934131380.

Weeks:1934:TMM

[Wee34b] Dorothy W. Weeks. Three mathematical methods of analyzing polarized light. Journal of Mathematics and Physics (MIT), 13(1–4): 371–379, April 1934. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1934131371.

Weinberger:1953:ULB

[Wei53] H. F. Weinberger. Upper and lower bounds for torsional rigidity. Journal of Mathematics and Physics (MIT), 32(1–4):54–62, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195332154.

Weinitschke:1959:SPS

[Wei59] Hubertus Weinitschke. On the stability problem for shallow spherical shells. Journal of Mathematics and Physics (MIT), 38(1–4):209– 231, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1959381209.

Weinberg:1962:SFD

[Wei62] I. J. Weinberg. Symmetric finite deflections of circular plates subjected to compressive edge forces. Journal of Mathematics and Physics (MIT), 41(1–4):104–115, April 1962. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1962411104.

Weinitschke:1965:ABS

[Wei65] H. J. Weinitschke. On asymmetric buckling of shallow spherical shells. Journal of Mathematics and Physics (MIT), 44(1–4):141–163, April 1965. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1965441141.

Weston:1960:TWF

[Wes60] V. H. Weston. On toroidal wave functions. Journal of Mathematics and Physics (MIT), 39(1–4):64–71, April 1960. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196039164.

Westmann:1964:PSC

[Wes64] R. A. Westmann. Pressurized star crack. Journal of Mathematics and Physics (MIT), 43(1–4):191–198, April 1964. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1964431191.

Westbrook:1968:AIS

[Wes68] D. R. Westbrook. On the accuracy of “Interior” solutions of the reduced wave equation in a plate-like region. Journal of Mathematics and Physics (MIT), 47(1–4):67–78, April 1968. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196847167.

Wheelon:1955:NSF

[Whe55] Albert D. Wheelon. Note on the summation of finite series. Journal of Mathematics and Physics (MIT), 34(1–4):182–186, April 1955. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1955341182.

Whitmore:1938:III

[Whi38] William F. Whitmore. Interreflections inside an infinite cylinder. Journal of Mathematics and Physics (MIT), 17(1–4):218–232, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1938171218.

Wiener:1922:NTI

[Wie22] Norbert Wiener. A new type of integral expansion. Journal of Mathematics and Physics (MIT), 1(3):167–176, April 1922. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192213167.

Wiener:1923:DS

[Wie23] Norbert Wiener. Differential-space. Journal of Mathematics and Physics (MIT), 2(1–4):131–174, October 1923. CODEN JMPHA9.

ISSN 0097-1421. URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm192321131.

Wiener:1924:CNP

[Wie24a] Norbert Wiener. Certain notions in potential theory. Journal of Mathematics and Physics (MIT), 3(1):24–51, January 1924. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19243124.

Wiener:1924:DP

[Wie24b] Norbert Wiener. The Dirichlet problem. Journal of Mathematics and Physics (MIT), 3(3):127–146, April 1924. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192433127.

Wiener:1924:MJL

[Wie24c] Norbert Wiener. In memory of Joseph Lipka. Journal of Mathematics and Physics (MIT), 3(2):63–65, March 1924. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19243263.

Wiener:1924:QVF

[Wie24d] Norbert Wiener. The quadratic variation of a function and its Fourier coefficients. Journal of Mathematics and Physics (MIT), 3(2):72–94, March 1924. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19243272.

Wiener:1925:NPP

[Wie25a] Norbert Wiener. Note on a paper of O. Perron. Journal of Mathematics and Physics (MIT), 4(1–4):21–32, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19254121.

Wiener:1925:NQA

[Wie25b] Norbert Wiener. Note on quasi-analytic functions. Journal of Mathematics and Physics (MIT), 4(1–4):193–199, January 1925. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192541193.

Wiener:1925:SDE

[Wie25c] Norbert Wiener. The solution of a difference equation by trigonometrical integrals. Journal of Mathematics and Physics (MIT), 4(1–4):153– 163, January 1925. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192541153.

Wiener:1926:HAIa

[Wie26a] Norbert Wiener. The harmonic analysis of irregular motion. Journal of Mathematics and Physics (MIT), 5(1–4):99–121, January 1926.

CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19265199.

Wiener:1926:HAIb

[Wie26b] Norbert Wiener. The harmonic analysis of irregular motion (second paper). Journal of Mathematics and Physics (MIT), 5(1–4):158–189, January 1926. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm192651158.

Wiener:1927:SAA

[Wie27] Norbert Wiener. The spectrum of an array and its application to the study of the translation properties of a simple class of arithmetical functions: Part One. The spectrum of an array. Journal of Mathematics and Physics (MIT), 6(1–4):145–157, April 1927. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192761145.

Wiener:1928:CMQ

[Wie28a] Norbert Wiener. Coherency matrices and quantum. Journal of Mathematics and Physics (MIT), 7(1–4):109–125, April 1928. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192871109.

Wiener:1928:NMT

[Wie28b] Norbert Wiener. A new method in Tauberian theorems. Journal of Mathematics and Physics (MIT), 7(1–4):161–184, April 1928. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192871161.

Wiener:1929:HAG

[Wie29a] Norbert Wiener. Harmonic analysis and group theory. Journal of Mathematics and Physics (MIT), 8(1–4):148–154, April 1929. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192981148.

Wiener:1929:HPF

[Wie29b] Norbert Wiener. Hermitian polynomials and Fourier analysis. Journal of Mathematics and Physics (MIT), 8(1–4):70–73, April 1929. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19298170.

Wiener:1931:NDG

[Wie31] Norbert Wiener. A new deduction of the Gaussian distribution. Journal of Mathematics and Physics (MIT), 10(1–4):284–288, April 1931. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1931101284.

Wiener:1935:RF

[Wie35] N. Wiener. Random functions. Journal of Mathematics and Physics (MIT), 14(1–4):17–23, April 1935. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm193514117.

Williams:1942:RSE

[Wil42] Martha Hathaway Williams. Ruled surfaces in Euclidean four space. Journal of Mathematics and Physics (MIT), 21(1–4):144–158, April 1942. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1942211144.

Williams:1957:RRE

[Wil57] W. E. Williams. Reflection and refraction of electromagnetic waves at plane interfaces. Journal of Mathematics and Physics (MIT), 36(1–4): 26–35, April 1957. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195736126.

Williams:1966:EST

[Wil66] M. M. R. Williams. An exact solution of the two group Milne problem by the method of Wiener and Hopf. Journal of Mathematics and Physics (MIT), 45(1–4):64–76, April 1966. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm196645164.

Willke:1967:STS

[Wil67] H. L. Willke, Jr. Stability in time-symmetric flows. Journal of Mathematics and Physics (MIT), 46(1–4):151–163, April 1967. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1967461151.

Winn:1937:CRF

[Win37] C. E. Winn. On certain reductions in the four color problem. Journal of Mathematics and Physics (MIT), 16(1–4):159–171, April 1937. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1937161159.

Wintner:1942:PDA

[Win42] Aurel Wintner. Prime divisors and almost periodicity. Journal of Mathematics and Physics (MIT), 21(1–4):52–56, April 1942. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm194221152.

Wintner:1945:RMS

[Win45] Aurel Wintner. Remarks on the method of stationary phases. Journal of Mathematics and Physics (MIT), 24(1–4):127–130, April 1945. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1945241127.

Wintner:1958:HGE

[Win58] Aurel Wintner. On Heaviside’s generalizations of the exponential function. Journal of Mathematics and Physics (MIT), 37(1–4):143–146, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371143.

Woodruff:1925:NME

[Woo25] L. F. Woodruff. Note on a method of evaluating the complex roots of sixth — and higher — order equations. Journal of Mathematics and Physics (MIT), 4(1–4):164–166, January 1925. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192541164.

Wiener:1938:GIT

[WP38] N. Wiener and H. R. Pitt. A generalization of Ikehara’s theorem. Journal of Mathematics and Physics (MIT), 17(1–4):247–258, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1938171247.

Wells:1945:PCF

[WS45] C. P. Wells and R. D. Spence. The parabolic cylinder functions. Journal of Mathematics and Physics (MIT), 24(1–4):51–64, April 1945. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm194524151.

Wu:1954:THF

[Wu54] Y. T. Wu. A theory for hydrofoils of finite span. Journal of Mathematics and Physics (MIT), 33(1–4):207–248, April 1954. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1954331207.

Wu:1966:MCT

[Wu66] Yau Wu. Markov chain theory of free-molecule flow. Journal of Mathematics and Physics (MIT), 45(1–4):169–178, April 1966. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1966451169.

Walsh:1922:EET

[WW22] J. L. Walsh and N. Wiener. The equivalence of expansions in terms of orthogonal functions. Journal of Mathematics and Physics (MIT), 1(2):103–122, March 1922. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm192212103.

Wiener:1938:SD

[WW38] Norbert Wiener and Aurel Wintner. On singular distributions. Journal of Mathematics and Physics (MIT), 17(1–4):233–246, April 1938. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1938171233.

Walsh:1954:DCS

[WY54] J. L. Walsh and David Young. On the degree of convergence of solutions of difference equations to the solution of the Dirichlet problem. Journal of Mathematics and Physics (MIT), 33(1–4):80–93, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm195433180.

Walsh:1957:LCH

[WY57] J. L. Walsh and David Young. Lipschitz conditions for harmonic and discrete harmonic functions. Journal of Mathematics and Physics (MIT), 36(1–4):138–150, April 1957. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1957361138.

Waterman:1964:NIN

[WYA64] P. C. Waterman, J. M. Yos, and R. J. Abodeely. Numerical integration of non-analytic functions. Journal of Mathematics and Physics (MIT), 43(1–4):45–50, April 1964. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm196443145.

Wynn:1956:NSM

[Wyn56] P. Wynn. A note on Salzer’s method for summing certain slowly convergent series. Journal of Mathematics and Physics (MIT), 35(1–4):

318–320, April 1956. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1956351318.

Yee:1966:IIJ

[Yee66] K. S. Yee. An identity involving the Jacobian elliptic function. Journal of Mathematics and Physics (MIT), 45(1–4):229–230, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1966451229.

Yeh:1963:DCP

[Yeh63] Cavour W. H. Yeh. On the dielectric coated prolate spheroidal antenna. Journal of Mathematics and Physics (MIT), 42(1–4):68–77, April 1963. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm196342168.

Yeh:1966:NII

[Yeh66] C. Yeh. A note on integrals involving parabolic cylinder functions. Journal of Mathematics and Physics (MIT), 45(1–4):231–232, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1966451231.

Yang:1956:RPL

[YL56] Hsun-Tiao Yang and Lester Lees. Rayleigh’s problem at low Mach number according to the kinetic theory of gases. Journal of Mathematics and Physics (MIT), 35(1–4):195–235, April 1956. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1956351195.

Young:1952:EBN

[You52] David Young. An error bound for the numerical quadrature of analytic functions. Journal of Mathematics and Physics (MIT), 31(1–4):42– 44, April 1952. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm195231142.

Young:1953:RMS

[You53] David Young. On Richardson’s method for solving linear systems with positive definite matrices. Journal of Mathematics and Physics (MIT), 32(1–4):243–255, April 1953. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1953321243.

Yuan:1959:TFC

[Yua59] S. W. Yuan. Turbulent flow in channels with porous walls. Journal of Mathematics and Physics (MIT), 38(1–4):166–171, April 1959. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1959381166.

Zadeh:1951:TDH

[Zad51] Lotfi A. Zadeh. Time-dependent Heaviside operators. Journal of Mathematics and Physics (MIT), 30(1–4):73–78, April 1951. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195130173.

Zadeh:1953:CSO

[Zad53] L. A. Zadeh. On a class of stochastic operators. Journal of Mathematics and Physics (MIT), 32(1–4):48–53, April 1953. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm195332148.

Zajac:1964:NOS

[Zaj64] E. E. Zajac. Note on overly-stable difference approximations. Journal of Mathematics and Physics (MIT), 43(1–4):51–54, April 1964. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm196443151.

Zeldin:1922:CCT

[Zel22a] S. D. Zeldin. Commutativity of contact transformations of mechanics. Journal of Mathematics and Physics (MIT), 1(3):157–159, April 1922. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm192213157.

Zeldin:1922:SHA

[Zel22b] S. D. Zeldin. Some hydrodynamic aspects of group theory. Journal of Mathematics and Physics (MIT), 1(1):54–62, January 1922. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm19221154.

Zeldin:1923:CTL

[Zel23a] S. D. Zeldin. Conformal transformations of linear homogeneous difference equations and their invariants. Journal of Mathematics and Physics (MIT), 2(1–4):63–72, October 1923. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm19232163.

Zeldin:1923:QTP

[Zel23b] S. D. Zeldin. On the quadratic ternary partial differential equation admitting Lie-groups of orders four and five. Journal of Mathematics and Physics (MIT), 2(1–4):234–254, October 1923. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm192321234.

Zeldin:1924:TM

[Zel24] S. D. Zeldin. On tautochronous motion. Journal of Mathematics and Physics (MIT), 3(3):147–161, April 1924. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm192433147.

Zeldin:1927:CTI

[Zel27] S. D. Zeldin. Contact transformations in intrinsic geometry. Journal of Mathematics and Physics (MIT), 6(1–4):223–239, April 1927. CODEN

JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm192761223.

Zemanian:1961:TPR

[Zem61] A. H. Zemanian. Transformations of positive-real functions and matrices arising from complex convolution. Journal of Mathematics and Physics (MIT), 40(1–4):260–270, April 1961. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1961401260.

Zwick:1954:DSV

[ZP54] S. A. Zwick and M. S. Plesset. On the dynamics of small vapor bubbles in liquids. Journal of Mathematics and Physics (MIT), 33(1–4):308– 330, April 1954. CODEN JMPHA9. ISSN 0097-1421. URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1954331308.

Zuckerberg:1966:SDP

[Zuc66] Hyam L. Zuckerberg. Solution of the Dirichlet problem in parallel slit domains. Journal of Mathematics and Physics (MIT), 45(1–4): 77–94, April 1966. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196645177.

Zwick:1958:BSPa

[Zwi58a] S. A. Zwick. Behavior of small permanent gas bubbles in a liquid: Part I. Isolated bubbles. Journal of Mathematics and Physics (MIT), 37(1–4):246–268, April 1958. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.1002/
sapm1958371246.

Zwick:1958:BSPb

[Zwi58b] S. A. Zwick. Behavior of small permanent gas bubbles in a liquid: Part II. Bubble clouds. Journal of Mathematics and Physics (MIT), 37(1–4): 339–353, April 1958. CODEN JMPHA9. ISSN 0097-1421. URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1958371339.

Zwick:1958:BSPc

[Zwi58c] S. A. Zwick. Behavior of small permanent gas bubbles in a liquid: Part III. A forced vibration problem. Journal of Mathematics and Physics (MIT), 37(1–4):354–370, April 1958. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1958371354.

Zygmund:1942:TNI

[Zyg42] A Zygmund. Two notes on inequalities. Journal of Mathematics and Physics (MIT), 21(1–4):117–123, April 1942. CODEN JMPHA9. ISSN

0097-1421. URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/sapm1942211117.